### v1

# Count-mean Sketch as an Optimized Framework for Frequency Estimation with Local Differential Privacy

Abstract
This paper identifies that a group of state-of-the-art locally-differentially-private (LDP) algorithms for frequency estimation are equivalent to the private Count-Mean Sketch (CMS) algorithm with different parameters. Therefore, we revisit the private CMS, correct errors in the original CMS paper regarding expectation and variance, modify the CMS implementation to eliminate existing bias, and explore optimized parameters for CMS to achieve optimality in reducing the worst-case mean squared error (MSE), $l_{1}$ loss, and $l_{2}$ loss. Additionally, we prove that pairwise-independent hashing is sufficient for CMS, reducing its communication cost to the logarithm of the cardinality of all possible values (i.e., a dictionary). As a result, the aforementioned optimized CMS is proven theoretically and empirically to be the only algorithm optimized for reducing the worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss when dealing with a very large dictionary. Furthermore, we demonstrate that randomness is necessary to ensure the correctness of CMS, and the communication cost of CMS, though low, is unavoidable despite the randomness being public or private.
1 Introduction
Frequency is a fundamental metric to describe the properties of a dataset, counting the occurrences of specific values. Calculating the frequency of each value is straightforward if the original values are known to the analyst, as one can generate a hash map to record the occurrences of all values in the dataset. This method has linear time and space complexity relative to the dataset size. However, frequency calculation becomes challenging when data is protected, and their values are not directly accessible to the analyst.
One of the most popular approaches to protect data is differential privacy [1], which is considered the de facto standard for extracting statistics from a dataset while protecting individual data privacy. Differential privacy has a local version [2], known as Local Differential Privacy (LDP), which requires enough perturbation on individual data such that an analyst or attacker cannot gain significant knowledge about the individual data values despite their prior knowledge [3]. However, LDP also adds more challenges to calculating statistics from a dataset, including frequency estimation.
A major challenge of LDP-preserving frequency estimation is that traditional mechanisms, such as randomized response [4], result in the variance of the estimated frequency increasing linearly with the number of all possible values in a dataset (referred to as a dictionary), making the estimation infeasible if the dictionary is too large. Several algorithms [5, 6, 7, 8, 9, 10, 11, 12, 13] have been developed to address this problem, which will be discussed in the Previous Work (Section 1.1). Many of these algorithms (e.g., [5, 9, 14, 12]) are found to be equivalent to the Count-Mean Sketch (CMS) algorithm (see Section 4 for more details). CMS hashes the original values through randomly selected hash functions and estimates the original frequencies based on the frequencies of the hashed values. The LDP version of CMS was originally developed by [7], but it had minor issues in calculating the expectation and variance (see Section 3.6) and did not explore the optimized parameters for CMS. Therefore, our work provides a corrected calculation of expectation and variance (Sections 3.1 and 3.2) and explores the optimized parameters of CMS (Section 3.4), which optimally reduce the worst-case mean squared error ($\widehat{MSE}$), $l_{1}$ loss, and $l_{2}$ loss as follows:
$\widehat{MSE}^{*}=\begin{cases}\frac{2(\Delta_{MSE}+e^{\epsilon})}{n(e^{\epsilon}-1)^{2}}&\text{ if }f^{*}\leq\frac{1}{2}\\ \frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}&\text{ if }\frac{1}{2}<f^{*}\leq 1\end{cases}$
$l_{1}^{*}\leq\sqrt{\frac{2d(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}}}\xrightarrow{d\rightarrow\infty}\frac{2de^{\epsilon/2}}{\sqrt{n}(e^{\epsilon}-1)}$
$l_{2}^{*}=\frac{2(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}}\xrightarrow{d\rightarrow\infty}\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$
with
$\Delta_{MSE}=e^{\epsilon/2}\sqrt{[(1-f^{*})e^{\epsilon}+f^{*}][f^{*}e^{\epsilon}+(1-f^{*})]}$
$\Delta_{l}=e^{\epsilon/2}\sqrt{(e^{\epsilon}+d-1)(de^{\epsilon}-e^{\epsilon}+1)}$
where $\epsilon$, $n$, and $d$ are the privacy factor, dataset size, and dictionary size, respectively; $f^{*}$ is defined as $\forall x\in\mathbf{x}:f(x)\leq f^{*}$ where $\mathbf{x}$ are the values of interest; * denotes the optimality. Moreover, our work proves that pairwise-independent hash functions are sufficient for CMS, and Section 3.5 proposes a practical approach to construct a pairwise-independent hashing family given any range of a hash function. The pairwise independence can reduce the communication cost to the logarithm of the dictionary size. As a result, Section 5.1 compares various state-of-the-art algorithms and concludes that our optimized CMS is the only algorithm optimized in the worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss while its communication cost is only logarithmic relative to the dictionary size. Additionally, Section 5.2 proves the necessity of randomness in CMS and provides a counterexample against the existing deterministic algorithm. Section 5.3 shows that the communication cost of CMS is unavoidable, despite the randomness being public or private. Section 6 empirically demonstrates that the optimized CMS outperforms other algorithms.
1.1 Previous Work
RAPPOR [6] is the first algorithm to estimate frequency with accuracy independent of dictionary size. It represents the original value as a one-hot vector and applies LDP perturbation to every bit of the array. However, the communication cost is proportional to the dictionary size, as the one-hot vector is the same size as the dictionary. [15] developed another version of RAPPOR, which hashes an original value to a smaller range than the dictionary size, such that the one-hot encoding results in a much smaller vector. This version uses gradient descent to search for one possible solution for the original frequencies. However, this mapping is equivalent to dimension reduction in linear algebra, so the found solution is not guaranteed to be close to the original frequencies. To tackle the large dictionary, [5] developed an algorithm that transforms an original value to a column of a Hadamard matrix, uniformly samples one element from the column, and perturbs it. The original frequencies can be reconstructed from the transformed values by applying the Hadamard matrix again. The accuracy of this approach is independent of the dictionary size, and the communication cost is the logarithm of the dictionary size. The Apple Privacy team [7] modified the Count-mean Sketch (CMS) algorithm to be LDP. However, they failed to derive the expectation and variance correctly (see Section 3.6) and did not explore the optimized settings of CMS. In addition to CMS, other members of the Count Sketch family have also been adapted to LDP (e.g., [8, 13]). [9] summarized existing approaches and proposed a variant of RAPPOR and a local hashing method, which were claimed to be optimal in terms of MSE. However, [9] assumed the estimated frequencies to be always close to zero, so their optimality did not hold when the original frequency approaches one. At the same time, [10] proposed a Subset Selection algorithm to estimate frequency with optimized accuracy in $l_{1}$ / $l_{2}$ losses, but its communication cost is also proportional to the dictionary size when the privacy requirement is high. [11] later integrated the Hadamard matrix into the subset selection and reduced the communication cost to the logarithm of the dictionary size. [12] further improved this method by applying the Hadamard transform twice. Though its performance can achieve order-optimal in $l_{1}$ / $l_{2}$ losses, it may perform worse than the Hadamard Response in terms of MSE (as discussed in Section 4.3).
2 Background
2.1 Local Differential Privacy (LDP)
Local Differential Privacy (LDP) was initially proposed in [2] to ensure that querying an object does not significantly disclose its actual value.
Definiton 2.1.
A mechanism $M$ is $\epsilon$-LDP if and only if $\forall(v,v^{\prime})\in\mathcal{V}^{2}$, and $\forall u\in\mathcal{U}$,
$Pr(M(v)=u|v)\leq e^{\epsilon}Pr(M(v^{\prime})=u|v^{\prime}).$
While there are other variations of LDP, this paper focuses solely on the most widely-used form mentioned above.
Given the above definition, $P$ is defined as a $|\mathcal{U}|\times|\mathcal{V}|$ matrix, where $P[u,v]$ represents the probability of the LDP mechanism $M$ generating $u$ given the input $v$. Here $u$ and $v$ are represented by an index between $1..|\mathcal{U}|$ and $1..|\mathcal{V}|$, respectively. Suppose there exists another matrix $Q$ of shape $|\mathcal{V}|\times|\mathcal{U}|$, which satisfies $I=QP$ and $I$ is an identity matrix. The decoding process of $M$ is defined to output $Q[:,u]$ given the output $u$ from $M$. The reconstruction process (perturbation and decoding) is denoted as $R$, whose output is an estimator of the input $v$, formulated as:
Property 2.1.
$\begin{gathered}E[R(v)[v]]=1\\ \forall v^{\prime}\neq v:E[R(v)[v^{\prime}]]=0\end{gathered}$
Proof: represent any $u$ from $M(v)$ as a one-hot vector of size $|\mathcal{U}|$, where the $u$-th element is one. A random variable $\mathbf{u}$ represents all the possible outputs from $M(v)$ as one-hot vectors. Subsequently, $R(v)=Q\mathbf{u}$. Therefore, $E[R(v)]=QE[\mathbf{u}]=QP\mathbf{e}_{v}=\mathbf{e}_{v}$, where $\mathbf{e}_{v}$ is a one-hot vector with the $v$-th element being one.
Meanwhile, we require the LDP mechanism $M$ to be stateless, meaning that its output depends only on its input. This is formulated as
$Pr(M(v)=u|v)=Pr(M(v)=u|v,\theta),$
where $\theta$ represents any other information (except the randomness generator used by $M$). Therefore, the outputs of $M$ from two inputs are mutually independent, which is formulated as:
Property 2.2.
If $v$ and $v^{\prime}$ are two inputs into $M$, the reconstruction of them from the outputs of $M$ satisfies $\forall(a,b)\in\mathcal{V}^{2}:Cov(R(v)[a]),R(v^{\prime})[b])=0$.
Most widely-used algorithms (e.g., randomized response [4] and RAPPOR [6]) satisfy this property. Additionally, we observe that these LDP mechanisms also exhibit symmetry in variance, formulated as:
Property 2.3.
When $\forall v,v^{\prime}\in\mathcal{V}^{2}$ and $v\neq v^{\prime}$, $Var(R(v)[v^{\prime}])$ are identical, denoted as $Var(R|\neq)$; $\forall v\in\mathcal{V}$, $Var(R(v)[v])$ are also identical, denoted as $Var(R|=)$.
2.2 Frequency Estimation
Let’s define some symbols: $X^{(i)}$ is the $i$-th object of an arbitrary dataset of size $n$; the possible values of all the objects comprise an dictionary $[d]$, where $d$ is its size, and $\forall i\in[n]:X^{(i)}\in[d]$. The frequency of value $x$ is defined as $f(x)=\frac{\sum_{i\in[n]}\mathbf{1}\{X^{(i)}=x\}}{n}$. The LDP workflow of frequency estimation usually proceeds as follows:
1. Each object $X^{(i)}$ is perturbed by an LDP mechanism, which outputs $Y^{(i)}$.
2. The server collects and aggregates the outputs as $\hat{f}(x)=A_{x}([Y^{(i)}]_{i\in[n]})$, if the server is interested in $f(x)$. The aggregation can be run for every $x\in[d]$.
$\hat{f}(x)$ is an estimator of $f(x)$. Existing literature usually uses $l_{1}$ and $l_{2}$ losses to measure its overall accuracy, which are defined as follows:
$\begin{gathered}l_{1}(\hat{f})=\max_{\forall\mathbf{X}:|\mathbf{X}|=n}E[\sum_{x\in[d]}|\hat{f}(x)-f(x)|]\\ l_{2}(\hat{f})=\max_{\forall\mathbf{X}:|\mathbf{X}|=n}E[\sum_{x\in[d]}(\hat{f}(x)-f(x))^{2}]\end{gathered}$
where $\mathbf{X}=[X^{(i)}]_{i\in[n]}$, and $\forall\mathbf{X}$ assumes the size $n$ is constant. These two losses measure the overall accuracy of the estimation. However, in many cases, an analyst is also interested in the accuracy of the frequency estimation of an individual value. The mean squared error (MSE) is commonly used to measure the accuracy of an estimation, reflecting both the variance and bias of the estimation. Since an analyst is assumed to have no knowledge of the actual dataset, they need to assume the worst case when calculating the MSE, which is defined as
$\widehat{MSE}(\hat{f})=\max_{\forall\mathbf{X}:|\mathbf{X}|=n}\max_{x\in\mathbf{x}}E[(\hat{f}(x)-f(x))^{2}],$ (1)
where $\widehat{MSE}$ denotes the upper bound of MSE, i.e., the worst case; $\mathbf{x}$ denotes the values of interest, which is $[d]$ by default. If the estimator $\hat{f}$ is unbiased, the worse-case MSE is the same as the worst-case variance, which is also positively correlated to the confidence interval (CI). Thus, an unbiased estimator optimized for $\widehat{MSE}$ is also optimized for CI.
3 Revisit Count-mean Sketch
Figure 1: Visualization of Count-mean Sketch.
Count-Mean Sketch (CMS) is a probabilistic data structure for frequency estimation, which (1) uses multiple hash functions to map objects into different counters, and (2) calibrates the bias in each counter to estimate the frequency of the values of interest. Suppose there is a dataset $[X^{(i)}]_{i\in[n]}$, each with a value belonging to a dictionary $[d]$. When setting up the CMS program, a hashing family $\mathcal{H}$ with $k$ hash functions mapping from $[d]$ to $[m]$ is provided. When uniformly sampling a hash function $h$ from $\mathcal{H}$, we expect that $\forall(x_{1},x_{2})\in[d]^{2}:Pr(h(x_{1})=h(x_{2}))\approx\frac{1}{m}$, which will be further discussed in Section 3.1. CMS operates with the following steps (Fig. 1):
1. Each object $X^{(i)}$ is assigned a hash function uniformly sampled from $\mathcal{H}$. The assigned hash function is denoted as $h_{j^{(i)}}$, where $j^{(i)}\in[k]$ is the index of the hash function. The server and the client sync on the hash-function assignment.
2. Hash $X^{(i)}$ as $y^{(i)}=h_{j^{(i)}}(X^{(i)})\in[m]$. Then, $y^{(i)}$ is perturbed by an $\epsilon$-LDP mechanism as $z^{(i)}$.
3. Collect all the responses $(z^{(i)})_{i\in[n]}$, decode them, and acquire the corresponding $\hat{y}^{(i)}$ as an estimator of $y^{(i)}$, i.e., $\hat{y}^{(i)}=R(y^{(i)})$, where $R$ is defined in Section 2.1. Notice that each $\hat{y}^{(i)}$ is a vector of size $m$.
4. The frequency of any value $x\in[d]$ can be estimated as
$\hat{f}(x)=\frac{m}{n(m-1)}\bigl{[}\sum_{i\in[n]}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{]}-\frac{1}{m-1}.$ (2)
The only output that depends on the value of $X^{(i)}$ is $z^{(i)}$, which is perturbed by an $\epsilon$-LDP mechanism. The postprocessing of the outputs from an LDP mechanism does not change its privacy guarantee [16], so we have proved:
Theorem 3.1.
The CMS frequency estimator $\hat{f}(x)$ is $\epsilon$-LDP.
3.1 Expectation and its Bias
This section derives the expectation of $\hat{f}(x)$. Here, we only consider the randomness of assigning hash functions to objects and the LDP perturbation, and regard $\mathcal{H}$ and $[X^{(i)}]_{i\in[n]}$ as constant. That is, $h_{j}(x)$ will not be considered a random variable; $\forall j\in[k],\forall(x,x^{\prime})\in[d]^{2}$, whether $h_{j}(x^{\prime})=h_{j}(x)$ is determined in advance as well, which will be denoted as $c_{j}(x,x^{\prime})$.
The hashing family $\mathcal{H}$ is treated as constant for the following reasons: (1) $\mathcal{H}$ is visible prior to executing CMS, as the server must sync it with all clients in advance to ensure that the clients know which hashing family to sample; and (2) though the hashing family $\mathcal{H}$ can be regenerated before each run of the CMS, we will show later that it is sufficient to have a constant hashing family $\mathcal{H}$ at all times.
Some literature [14] also proposed a deterministic approach to assign hash functions, which will be discussed in Section 5.2. Here, we only focus on the CMS defined above and require the hash-function assignments to be random events. Now, the expectation of the frequency estimation by CMS is derived as:
Theorem 3.2.
The expectation of $\hat{f}(x)$ is
$E[\hat{f}(x)]\\ =\frac{m}{m-1}[f(x)+\sum_{x^{\prime}\in[d]\backslash x}\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}f(x^{\prime})]-\frac{1}{m-1}$ (3)
where $f(x)$ is the frequency of value $x$; $[d]\backslash x$ represents the set of $[d]$ excluding $x$.
Corollary 3.1.
If and only if $\forall x,x^{\prime}:\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}=\frac{1}{m}$,
$E[\hat{f}(x)]=f(x).$
The proof of Theorem 3.2 and Corollary 3.1 is presented in Appendix A. It is notable that $\forall x,x^{\prime}:\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}=\frac{1}{m}$ is equivalent to $\Pr_{h\in\mathcal{H}}(h(x)=h(x^{\prime}))=\frac{1}{m}$, where $h\in\mathcal{H}$ represents uniform sampling, which is also the definition of a pairwise-independent hashing family. Thus, Corollary 3.1 can be rephrased as:
Corollary 3.2.
If and only if $\mathcal{H}$ is pairwise independent, i.e., $\Pr_{h\in\mathcal{H}}(h(x)=h(x^{\prime}))=\frac{1}{m}$, we have $E[\hat{f}(x)]=f(x)$.
The pairwise independence requires the size of $\mathcal{H}$ to be $d^{2}$, so the communication cost of sending one hash function becomes $2\log_{2}(d)$. The creation of such a hashing family is presented in Section 3.5.
3.2 Variance
Assuming that the LDP reconstruction process $R$ in Step 3 of CMS satisfies Properties 2.1 and 2.2, we can derive:
Lemma 3.1.
If $R$ satisfies Properties 2.1 and 2.2, we have
$Var(\hat{f}(x))=\frac{m^{2}}{(m-1)^{2}n}\sum_{x^{\prime}\in[d]}\biggl{[}\biggl{(}\frac{1}{k}\sum_{j\in[k]}\\ Var\bigl{(}R(h_{j}(x^{\prime}))[h_{j}(x)]\bigr{)}\biggr{)}+\bar{c}(x^{\prime},x)-\bar{c}(x^{\prime},x)^{2}\biggr{]}f(x^{\prime}),$ (4)
where $\bar{c}(x^{\prime},x)=\sum_{j\in[k]}\frac{c_{j}(x^{\prime},x)}{k}$.
Combining Lemma 3.1 and Property 2.3, we can derive:
Theorem 3.3.
If $R$ satisfies Properties 2.1 - 2.3, we have
$Var(\hat{f}(x))=\frac{m^{2}}{(m-1)^{2}n}\biggl{[}\\ \sum_{x^{\prime}\in[d]\backslash x}f(x^{\prime})\biggl{(}\bar{c}(x^{\prime},x)Var(R|=)+(1-\bar{c}(x^{\prime},x))\\ Var(R|\neq)+\bar{c}(x^{\prime},x)(1-\bar{c}(x^{\prime},x))\biggr{)}+f(x)Var(R|=)\biggr{]},$ (5)
where $Var(R|=)$ and $Var(R|\neq)$ are defined in Property 2.3.
If the hashing family $\mathcal{H}$ is pairwise independent, we further have:
Corollary 3.3.
If $R$ satisfies Properties 2.1 - 2.3 and $\mathcal{H}$ is pairwise independent (i.e., $\forall x,x^{\prime}:\bar{c}(x^{\prime},x)=\frac{1}{m}$), we have
$Var(\hat{f}(x))\\ =\frac{m}{(m-1)^{2}n}\biggl{[}\bigl{(}1-f(x)\big{)}\biggl{(}Var(R|=)+(m-1)\\ Var(R|\neq)+\frac{m-1}{m}\biggr{)}+mf(x)Var(R|=)\biggr{]}.$ (6)
The lemma, theorem, and corollary of this section are proven in Appendix B.
3.3 Accuracy Analysis
The mean squared error (MSE) measures the accuracy of $\hat{f}(x)$, which is defined as
$MSE(\hat{f}(x))$ $=E[(\hat{f}(x)-f(x))^{2}]$
$=E[\hat{f}(x)-E[\hat{f}(x)]]^{2}+(E[\hat{f}(x)]-f(x))^{2}$
$=Var(\hat{f}(x))+(E[\hat{f}(x)]-f(x))^{2}$ (7)
When $\mathcal{H}$ is pairwise independent, we have $E[\hat{f}(x)]-f(x)=0$, so $MSE(\hat{f}(x))=Var(\hat{f}(x))$. If a hashing family is not pairwise independent but still satisfies $\forall(x,x^{\prime})\in[d]^{2}:\bar{c}(x,x^{\prime})$ is identical, we can create an unbiased frequency estimator of $f(x)$ as
$g(x)=\frac{m^{\prime}}{n(m^{\prime}-1)}\bigl{[}\sum_{i\in[n]}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{]}-\frac{1}{m^{\prime}-1},$ (8)
where $m^{\prime}=\frac{1}{\bar{c}(x,x^{\prime})}$. $g(x)$ has the following properties:
Theorem 3.4.
$E[g(x)]=0$ (9)
and
$MSE(g(x))=\frac{1}{n(1-\bar{c}(x,x^{\prime}))^{2}}\biggl{[}\\ (1-f(x))\biggl{(}\bar{c}(x^{\prime},x)Var(R|=)+(1-\bar{c}(x^{\prime},x))\\ Var(R|\neq)+\bar{c}(x^{\prime},x)(1-\bar{c}(x^{\prime},x))\biggr{)}+f(x)Var(R|=)\biggr{]}.$ (10)
The formal proof is similar to that of the original estimator $\hat{f}(x)$, so the proof is omitted here. The rationale is that the original estimator $\hat{f}(x)$ assumes $\bar{c}(x,x^{\prime})$ being $\frac{1}{m}$, but here we integrate $\bar{c}(x,x^{\prime})$ into the estimation and replace $\frac{1}{m}$ with $\bar{c}(x,x^{\prime})$.
Note that Theorem 10 is not applicable to the original CMS [7] because its $\bar{c}(x,x^{\prime})$ is not constant. We will analyze its accuracy in Section 3.6.
Besides MSE, one may also use the confidence interval (CI) of $\hat{f}(x)$ to measure accuracy. If $\hat{f}(x)$ is unbiased, its CI is the same as its concentration bound. Moreover, $\hat{f}(x)$ is equivalent to the sum of bounded random variables (i.e., $\sum\hat{y}$), so its bounds can be derived from the Chernoff bound. Below is the concentration bound for $\hat{f}(x)$ from the unbiased CMS using randomized response (referred to as CMS+RR):
Theorem 3.5.
CMS+RR with pairwise-independent hashing satisfies $\forall\alpha\in[0,\sqrt{\frac{e^{\epsilon}n}{m-1}}]$,
$Pr\bigl{[}|\hat{f}(x)-f(x)|\geq\alpha\sqrt{Var(\hat{f}(x))}\bigr{]}\\ \leq 2\exp(-\frac{\alpha^{2}}{3}\cdot\frac{m-1}{e^{\epsilon}+m-1}).$
This theorem is proven in Appendix C. Note that CMS+RR will be proven to be optimized in Section 3.4, so this theorem is applicable directly to the optimized CMS.
3.4 Optimized CMS
This section explores the optimized parameters of CMS to reduce the $l_{1}$ loss, $l_{2}$ loss, and the worst-case MSE defined in Section 2.2. It is notable that $l_{2}$ loss is connected to MSE as follows:
$l_{2}(\hat{f})=\max_{\forall\mathbf{X}:|\mathbf{X}|=n}\sum_{x\in[d]}E[(\hat{f}(x)-f(x))^{2}]\\ =\max_{\forall\mathbf{X}:|\mathbf{X}|=n}\sum_{x\in[d]}MSE(\hat{f}(x)).$ (11)
Additionally, given the Cauchy-Schwarz inequality, we have $l_{1}(\hat{f})\leq\sqrt{d\>l_{2}(\hat{f})}$. Thus, both losses are connected to MSE. Based on Eq. (7) and Corollary 3.1, if $\mathcal{H}$ is pairwise independent, MSE is equivalent to $Var(\hat{f}(x))$, which can be calculated using Eq. (6). Assuming that the privacy factor $\epsilon$ and the dictionary $[d]$ are given, the undetermined variables in this equation are only the LDP process $R$ and the range of a hash function $m$.
We start with the LDP process $R$. [9] summarized the state-of-the-art frequency estimation algorithms at that time. When considering different privacy regimes and original frequencies, i.e., $f(x)$, we found that randomized response (RR), symmetric RAPPOR (s-RP), and asymmetric RAPPOR (a-RP) are optimized in some privacy regimes. Though more optimized algorithms exist, their performance is either as optimized as these algorithms in their respective regimes, or they can be converted to CMS+RR (as shown in Section 4). Thus, we only consider RR, s-RP, and a-RP as candidates for the LDP process $R$.
The variance of the these mechanisms are listed below:
$Var(RR(a)[b])=\\ \frac{1}{(e^{\epsilon}-1)^{2}}\begin{cases}e^{\epsilon}(m-1)&\text{ if }a=b\\ e^{\epsilon}+m-2&\text{ if }a\neq b,\end{cases}$ (12)
$Var(aRP(a)[b])=\\ \frac{1}{(e^{\epsilon}-1)^{2}}\begin{cases}(e^{\epsilon}+1)^{2}&\text{ if }a=b\\ 4e^{\epsilon}&\text{ if }a\neq b,\end{cases}$ (13)
and
$\forall a,b:Var(sRP(a)[b])=\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}.$ (14)
Then, we explore the optimized $m$ for different LDP processes and compare their loss functions and the worst-case MSE. We start with the worst-case MSE. Here, we allow an analyst to have prior knowledge of the upper bound of $f(x)$, and we derive:
Theorem 3.6.
If the upper bound of $f(x)$ is known as $\forall x\in\mathbf{x}:f(x)\leq f^{*}$, where $\mathbf{x}$ are the values of interest, the worst-case MSE among $\mathbf{x}$ is optimized as
$\widehat{MSE}(\hat{f})=\begin{cases}\frac{2(\Delta_{MSE}+e^{\epsilon})}{n(e^{\epsilon}-1)^{2}}&\text{ if }f^{*}\leq\frac{1}{2}\\ \frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}&\text{ if }\frac{1}{2}<f^{*}\leq 1\end{cases},$
where
$\Delta_{MSE}=e^{\epsilon/2}\sqrt{[(1-f^{*})e^{\epsilon}+f^{*}][f^{*}e^{\epsilon}+(1-f^{*})]},$
which can be achieved when the LDP process $R$ is randomized response, and $m$ is the closest integer to
$\begin{cases}1+\frac{\Delta_{MSE}}{f^{*}e^{\epsilon}+1-f^{*}}&\text{ if }f^{*}\leq\frac{1}{2}\\ 1+e^{\epsilon/2}&\text{ if }\frac{1}{2}<f^{*}\leq 1.\end{cases}$ (15)
The theorem is proven in Appendix D. Notably, $f^{*}$ is prior knowledge of an analyst and could be larger than $\max_{x}f(x)$, which is unknown to the analyst. Subsequently, we have:
Corollary 3.4.
If the upper bound of $f(x)$ is unknown, i.e., $f^{*}=1$, the worst-case MSE is optimized as
$\widehat{MSE}(\hat{f})=\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}},$
which can be achieved when $R$ is randomized response, and $m$ is the closest integer to $1+e^{\epsilon/2}$
Hereinafter, we assume no prior knowledge of $f(x)$ unless explicitly mentioned, i.e., $f^{*}=1$ by default. On the other hand, if an analyst is interested in some rare values, and the corresponding $f^{*}$ is around zero, we can substitute $f^{*}=0$ into Theorem 15 and obtain $m=1+e^{\epsilon}$, which should be sufficient to approximate the optimized $m$. Its worst-case variance is derived as:
$\widehat{MSE}(\hat{f})=\frac{1}{n}[f^{*}+\frac{4e^{\epsilon}}{(e^{\epsilon}-1)^{2}}].$ (16)
We can see $m=1+e^{\epsilon}$ works well if $f^{*}$ is not a dominant term in Eq. (16).
Next, we derive the optimized $l_{1}$ and $l_{2}$ losses:
Theorem 3.7.
The $l_{1}$ and $l_{2}$ losses are optimized as
$l_{1}^{*}(\hat{f})\leq\sqrt{\frac{2d(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}}},$
and
$l_{2}^{*}(\hat{f})=\frac{2(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}},$
respectively, where
$\Delta_{l}=e^{\epsilon/2}\sqrt{(e^{\epsilon}+d-1)(de^{\epsilon}-e^{\epsilon}+1)}.$
Both optimalities are achieved when the LDP process $R$ is randomized response, and $m$ is the closest integer to $1+\frac{\Delta_{l}}{e^{\epsilon}+d-1}$.
Corollary 3.5.
When the dictionary size $d$ is large enough, the optimized $l_{1}$ and $l_{2}$ losses approach
$\lceil l_{1}\rceil^{*}(\hat{f})\rightarrow\frac{2de^{\epsilon/2}}{\sqrt{n}(e^{\epsilon}-1)},$
and
$l_{2}^{*}(\hat{f})\rightarrow\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}},$
respectively, where $\lceil l_{1}\rceil$ represents the upper bound of the $l_{1}$ loss, and the corresponding $m\rightarrow 1+e^{\epsilon}$.
The theorem and corollary above are proven in Appendix D. It should be noted that the CMS using symmetric or asymmetric RAPPOR as its LDP process cannot serve as the optimized CMS for the following reasons: (1) their performance in reducing worst-case MSE or $l_{1}$ / $l_{2}$ losses is outperformed by CMS+RR in some conditions, and (2) they have very high communication costs. Please refer to Appendix D for more details.
One caveat of the setting to minimize $l_{1}$ / $l_{2}$ losses is that it will lead to a worst-case MSE larger than the optimized value. Assuming $d$ is large enough, and $m=1+e^{\epsilon}$, we have
$\widehat{MSE}(\hat{f})=Var(\hat{f}(x)_{RR}|f(x)=1)=\frac{(e^{\epsilon}+1)^{2}}{n(e^{\epsilon}-1)^{2}},$ (17)
which is always larger than the optimized $\widehat{MSE}$, i.e., $\frac{e^{\epsilon/2}}{n(e^{\epsilon}-1)^{2}}$. What is worse, even when $\epsilon$ is very large, it only converges to $\frac{1}{n}$, not zero. Therefore, before an analyst estimates the frequencies of a dataset, they need to determine whether to focus on (1) the overall accuracy (i.e., $l_{1}$ / $l_{2}$ losses) or (2) the worst-case accuracy of the individual estimations (i.e., $\widehat{MSE}$).
Since optimizing $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses requires different $m$, we will hereinafter refer to the CMS optimized for $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses as MSE-OCMS and $l$-OCMS, respectively. Additionally, the optimized CMS will be compared with the existing algorithms in Section 5.1.
3.5 Unbiased CMS with Imperfect Hashing
An optimized CMS is built upon an unbiased CMS, which requires its hashing to be pairwise independent. However, the construction of a $K$-wise-independent hashing family requires the range of the hashed values (i.e., $[m]$) to comprise a finite field [17]. A finite field must have $p^{l}$ elements [18], where $p$ is a prime number and $l\in\mathbb{N}^{+}$. For instance, if a CMS has $m=6$, there is no way to construct a pairwise independent hashing family for it. In practice, we can still construct an approximately pairwise-independent hashing family for any $m$. Consider a perfect pairwise independent hashing family with $p=2$ and $2^{l}\geq d$:
=$\mathcal{H}=\{h(x)|\forall a_{i}\in[2^{l}]:h(x)=a_{0}+a_{1}x\}$ (18)
where $a_{0}$, $a_{1}$, and $x$ are elements in a finite field $\mathbb{F}(2^{l})$, so the addition and multiplication in Eq. (18) are all field operations. The range of any $h(x)$ is also the field $\mathbb{F}(2^{l})$. Given two random variables $a_{0}$ and $a_{1}$, the size of $\mathcal{H}$ is $2^{2l}=\Theta(d^{2})$. Subsequently, we construct a new hashing family based on $\mathcal{H}$:
=$\mathcal{H}^{\prime}=\{h^{\prime}(x)|h(x)\bmod m\}.$
Since $\forall h\in\mathcal{H}$, $\forall y\in[m]$: $Pr(h(x)=y)=2^{-l}$, and assuming $2^{l}=qm+r$, where $q\in\mathbb{N}$ and $r\in\mathbb{N}$, we will have $r$ hashed values with $Pr(h^{\prime}(x)=y)=\frac{q+1}{2^{l}}$ and the remaining $m-r$ hashed values with $Pr(h^{\prime}(x)=y)=\frac{q}{2^{l}}$. Thus, we have:
Lemma 3.2.
$\forall x_{1},x_{2}\in[d]^{2}$:
$Pr_{h^{\prime}\in\mathcal{H}^{\prime}}(h^{\prime}(x_{1})=h^{\prime}(x_{2}))=r(\frac{q+1}{2^{l}})^{2}+(m-r)(\frac{q}{2^{l}})^{2}\\ =\frac{(2q+1)r+mq^{2}}{(mq+r)^{2}},$ (19)
where $h^{\prime}\in\mathcal{H}^{\prime}$ represents uniform sampling.
Notice that $Pr_{h^{\prime}\in\mathcal{H}^{\prime}}(h^{\prime}(x_{1})=h^{\prime}(x_{2}))=\bar{c}(x_{1},x_{2})$, which is a constant and can be substituted into the $g(x)$ in Theorem 10, where $m^{\prime}=\frac{(mq+r)^{2}}{(2q+1)r+mq^{2}}$. As a result, $g(x)$ is an unbiased estimator of $f(x)$. We can further prove that
Theorem 3.8.
When
$\frac{2mVar(R|=)+m-1}{n(m-1)^{3}}\frac{m}{\tau Var(\hat{f}(x))}\leq(2q+1)^{2},$
we have
$Var(g(x|\mathcal{H}^{\prime}))<(1+\tau)Var(\hat{f}(x)),$
where $g(x|\mathcal{H}^{\prime})$ and $\hat{f}(x)$ represent the $g(x)$ with $\mathcal{H}^{\prime}$ and a standard estimator of $f(x)$ with a pairwise-independent hashing family.
Consider the optimized CMS for the worst-case MSE (Corollary 3.4), we can derive:
Corollary 3.6.
When $2q+1>\sqrt{\frac{1}{\tau}}$ and $m=1+e^{\epsilon/2}$, we have
$MSE(g)<(1+\tau)\widehat{MSE}(\hat{f})$
Corollary 3.7.
When $2q+1>\sqrt{\frac{1}{\tau}}$ and $m=1+e^{\epsilon}$, we have
$l_{2}(g)<(1+\tau)l_{2}^{*}(\hat{f})$
The condition of these two corollaries is independent of $\epsilon$. If $\tau=0.01$, we have $q>4.5$ to satisfy the two corollaries above.
On the other hand, one may prefer using the original frequency estimator $\hat{f}(x)$ in Eq. (2) instead of $g(x)$, while still employing the imperfect hashing. The following theorem will demonstrate how large $q$ is required to make the deviation from $MSE(\hat{f}(x))$ negligible.
Theorem 3.9.
When
$\frac{1}{4(m-1)q(q+1)}\leq\sqrt{\tau Var(\hat{f}(x))}$
and
$\frac{8q^{2}+8q+1}{16q^{2}(q+1)^{2}}\leq\tau,$
we have
$MSE(f(x)|\mathcal{H}^{\prime})<(1+2\tau)MSE(\hat{f}(x)),$
where $MSE(\hat{f}(x))$ represents the MSE of the $f(x)$ with the perfect hashing.
Notably, $Var(\hat{f}(x))=\Theta(1/n)$, so $q=\Theta(n^{1/4})$. That is, the size of $\mathcal{H}^{\prime}$ has to increase with the size of the dataset.
All the theorems and corollaries in this section are proven in Appendix E.
3.6 Correcting the original CMS
In the original CMS paper [7], CMS randomly generates $k$ hash functions, and each hash function $h_{j}(x)$ is generated by sampling $d$ mutually independent random variables, each uniformly distributed among $[m]$. Subsequently, $\forall(x,x^{\prime})\in[d]^{2},c_{j}(x,x^{\prime})$ can be considered as a Bernoulli random variable with a probability of $\frac{1}{m}$ of being one. Since the $k$ hash functions are generated independently, $\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}$ is equivalent to a binomial random variable, whose expectation is $\frac{1}{m}$. However, the probability of the output of a binomial distribution exactly equaling its expectation is very low, so $\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}$ will likely deviate from $\frac{1}{m}$ to some extent (see Slud’s Inequality [19]). Thus, $E[\hat{f}(x)]$ will likely deviate from $f(x)$ to some extent too, i.e., $\hat{f}(x)$ has bias. Notably, this bias is independent of the size of the dataset, so it will still exist even when an infinite amount of data is given for a single run of CMS. This conclusion is also applicable to the non-private Count Sketch families, which are also proven to have bias [20] if they query a dataset only once.
To evaluate the degree of bias, when considering $\forall c_{j}(x,x^{\prime})$ as random variables, we can derive:
Theorem 3.10.
[…]
and
$\mathop{Var}_{\forall c_{j}(x,x^{\prime})}(E[\hat{f}(x)])=\frac{1}{(m-1)k}\sum_{x^{\prime}\in[d]\backslash x}f(x^{\prime})^{2}.$ (21)
The proof is presented in Appendix F.1. The variance indicates that the bias is $O(\frac{1}{\sqrt{(m-1)k}})$. However, the original CMS paper [7] derived $E[\hat{f}(x)]$ to be unbiased. We found that they treated $c_{j}(x,x^{\prime})$ as independent random variables for every pair of $X^{(i)}$ and $X^{(i^{\prime})}$, which is incorrect because all pairs with the same values $x$ and $x^{\prime}$ will always have the same $c_{j}(x,x^{\prime})$.
Eq. (7) demonstrates how the bias of $\hat{f}(x)$ contributes to its MSE. The following theorem demonstrates how to prevent the bias from being a dominant term in $MSE(\hat{f}(x))$:
Theorem 3.11.
If
$\frac{\delta\tau Var(R|\neq)}{n}>\frac{1}{(m-1)k}$
is satisfied, where $\tau$ is a small number, like 0.1, we have
$Pr\biggl{(}(E[\hat{f}(x)]-f(x))^{2}\leq\tau Var(\hat{f}(x))\biggr{)}\geq 1-\delta.$ (22)
The proof is presented in Appendix F.2. This implies that the number of hash functions $k$ should be $\Omega(\frac{n}{m})$ to ensure that the bias of $f(x)$ is negligible. However, this constraint was not mentioned in the original paper.
The original paper also has an issue in its variance calculation, where they required the hashing family $\mathcal{H}$ to be 3-wise-independent because they assumed any $X^{(i_{1})}$ and $X^{(i_{2})}$ were correlated, which was problematic. In contrast, the non-private Count Sketch [20] only requires pairwise independence, which is consistent with our work. Notably, the LDP perturbation only affects the accuracy after the hashing, so it should not alter the pairwise-independence requirement for the hashing family.
4 Connection with Existing Frequency Estimation Algorithms
This section demonstrates that various frequency estimation algorithms proposed in the previous literature are equivalent or closely connected to CMS. First, we define the Transformation-and-Sampling Framework (TnS Framework) as follows:
Definiton 4.1.
(TnS Framework)
1. Transform the original value $x$ to a vector of elements $M[x]$, i.e., a row of a matrix $M$.
2. Uniformly sample an element from $M[x]$, denoted as $M[x][j]$
3. Perturb the sampled element $M[x][j]$ with some LDP mechanism.
The above process is equivalent to (1) uniformly sampling a column from $M$, denoted as $M[:,j]$, (2) mapping the original value $x$ to $M[x,j]$, and (3) perturbing it. The column here functions as a hash function, and all the columns of $M$ comprise a hashing family. Thus, we have proved:
Theorem 4.1.
An algorithm belonging to the TnS Framework is equivalent to CMS .
4.1 Hadamard Encoding
Hadamard Encoding (HE) is an algorithm that (1) creates a Hadamard matrix with a size larger than the dictionary size $d$, (2) select a row from the Hadamard matrix with an index equal to $x+1$, and (3) uniformly samples an element from the selected row, and (4) perturbs the element using randomized response. It was originally proposed by [5] and has been widely used in the literature, e.g., [7, 12].
It is evident that HE belongs to the TnS framework, so it is equivalent to CMS. Additionally, Appendix G proves that decoding process of HE is equivalent to Eq. (2). Moreover, all the rows of a Hadamard matrix, excluding the first element, comprise a 3-wise-independent hashing family [21], mapping to either -1 or 1. Therefore, HE is equivalent to the CMS+RR using pairwise-independent hashing with $m=2$.
4.2 Hadamard Response
Hadamard Response was originally proposed by [11] and later developed in [14]. It has two equivalent versions, and the version in [14] is presented here: (1) publicly and randomly selects a row from a Hadamard matrix with size larger than the dictionary size $d$, (2) asks if the original value is mapped to one in the selected row (1 means yes, 0 means no), and (3) perturbs the answer using randomized response. It is straightforward to see that Step 1 is equivalent to sampling a hash function, and Step 2 is equivalent to hashing the original value. Additionally, its decoding process is equivalent to HE. Therefore, this algorithm is also equivalent to the CMS+RR using pairwise-independent hashing with $m=2$.
4.3 Recursive Hadamard Response
Recursive Hadamard Response (RHR) [12] is an improvement of Hadamard Encoding, increasing accuracy by communicating more bits. It (1) splits the original value $x$ into $x//2^{b-1}$ (floor division) and $x\bmod 2^{b-1}$; (2) creates a Hadamard matrix with a size larger than $d//2^{b-1}$, (3) select a row from the Hadamard matrix with an index equal to $(x//2^{b-1})+1$, and (4) uniformly samples an element from the selected row as $s$, and (5) perturbs $s\times(x\bmod 2^{b-1})$ using randomized response.
Steps 2 - 5 can be fitted in the TnS framework, mapping $x$ to $2^{b}$ possible values. However, the collision probability of two inputs is not uniform. If two inputs $x_{1}$ and $x_{2}$ have the same mod (i.e., $x_{1}\bmod 2^{b-1}=x_{2}\bmod 2^{b-1}$), their collision probability is $1/2$. Otherwise, the probability is zero. In CMS, any pair of inputs have the same collision probability of $1/m$, which is $2^{-b}$ given $2^{b}$ hashed values. Thus, RHR is a skewed version of CMS.
If the dataset is randomly generated, i.e., every pair of objects may have the same mod with a probability of $2^{b-1}$, then the hash collision probability becomes $2^{-b}$. Note that [12] sets $b=\epsilon$ if the dictionary is large enough. This is equivalent to $m=e^{\epsilon}$ in CMS, which is close but not identical to the optimized $m$ used by OCMS, so the accuracy of RHR at most has the same order as OCMS.
However, we can always generate a dataset with all values having the same mod when divided by $2^{b-1}$. This makes all these values have a probability of $1/2$ to collide with one another. Therefore, RHR in this case will perform worse than regular Hadamard Encoding, because they have the same collision probability but the randomized response of RHR considers more possible values, leading to higher variance. In the context of CMS, the MSE of RHR is equivalent to setting $\bar{c}(x,x^{\prime})=\frac{1}{2}$ while setting $R$ as a randomized response with $m>2$ in Eq. (5). Section 6.3 demonstrates that even a real dataset will encounter this problem if it contains values with high frequency.
4.4 Local Hashing
Local Hashing [9] asks each client to (1) publicly and randomly generate a hashing function mapping $[d]$ to $[m]$, (2) hash the original value, and (3) perturb the hashed value using randomized response. It is straightforward to see that randomly generating a hashing function is equivalent to randomly selecting a hash function from the hashing family in CMS. Also, its decoding process is identical to Eq. (2). Therefore, this algorithm is equivalent to CMS. However, [9] did not compute the optimized $\widehat{MSE}$ correctly as they only considered $f(x)=0$, but Section 3.4 demonstrates that both $f(x)=0$ and $f(x)=1$ could be the maximum points of the MSE. Moreover, they did not optimize the communication cost of sending a hash function to the server, which will be discussed in Section 5.1.
4.5 Recursive CMS
When the private CMS was initially proposed in [7], it suggested an algorithm to (1) map an original value to $[m]$ where $m$ is large, and then (2) use Hadamard Encoding to perturb the hashed value. Section 4.1 has demonstrated that Hadamard Encoding is equivalent to CMS, so the CMS with Hadamard Encoding (CMS+HE) forms a recursive CMS. Generally speaking, the perturbation algorithm $R$ in a CMS can also be another CMS, and we will prove that a recursive CMS is equivalent to a regular CMS.
Here, we assume all the CMS involved are unbiased, i.e., the hashing family is pairwise independent. Suppose the first CMS hashes the original values to $[m_{1}]$, and its perturbation algorithm is another CMS hashing its input to $[m_{2}]$. Given the pairwise independence, the collision probability $\bar{c}(x,x^{\prime})$ of two distinct input values is $\frac{m_{1}+m_{2}-1}{m_{1}m_{2}}$. Therefore, the recursive CMS is equivalent to a regular CMS with $m=\frac{m_{1}m_{2}}{m_{1}+m_{2}-1}$.
Returning to the CMS+HE, it is equivalent to the CMS+RR with $m^{\prime}=\frac{2m}{m+1}$, where $m^{\prime}$ and $m$ are the hashing ranges of the equivalent and original CMS, respectively. If $m$ is large enough, it is equivalent to the CMS+RR with $m^{\prime}=2$ (i.e., Hadamard Encoding).
4.6 RAPPOR
The encoding of RAPPOR is equivalent to CMS if each object in RAPPOR belongs to its own cohort and there is only one hash function in each cohort. Assuming there is sufficient data and we can ignore the uncertainty from the LDP mechanism and hash-function assignment, the encoding of RAPPOR and CMS can be considered a linear transformation from $\mathcal{R}^{d}$ (frequencies of original values) to $\mathcal{R}^{km}$ (frequencies of the hashed values). For the unbiased CMS (Corollary 3.1), we have $km=d^{2}m>d$, so CMS can derive the original frequencies without bias. However, RAPPOR utilizes gradient descent to find the original frequencies and does not require $km\geq d$. [15] even chooses small $km$ on purpose to reduce computation. Unfortunately, this is problematic because multiple values in a high-dimension simplex will map to the same value in a low-dimension simplex, and the value RAPPOR finds may not be the original value.
5 Discussion
5.1 Accuracy, Communication Cost, and Optimality
Tables I and II compare the accuracy and communication cost among various frequency estimation algorithms, respectively. We will demonstrate that the optimized CMS is also the optimal algorithm when considering both accuracy and communication cost.
$l_{1}$ $l_{2}$ $\widehat{MSE}$
HE $\frac{d(e^{\epsilon}+1)}{e^{\epsilon}-1}$ $\frac{d(e^{\epsilon}+1)^{2}}{(e^{\epsilon}-1)^{2}}$ $(\frac{e^{\epsilon}+1}{e^{\epsilon}-1})^{2}$
RHR $\Theta(\frac{d}{\sqrt{\min\{\epsilon,\epsilon^{2}\}}})$ $\Theta(\frac{d}{\min\{\epsilon,\epsilon^{2}\}})$ $\Omega((\frac{e^{\epsilon}+1}{e^{\epsilon}-1})^{2})$
OLH $\frac{2de^{\epsilon/2}}{e^{\epsilon}-1}$ $\frac{4de^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ $(\frac{e^{\epsilon}+1}{e^{\epsilon}-1})^{2}$
OCMS $\frac{2de^{\epsilon/2}}{e^{\epsilon}-1}$ $\frac{4de^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ $\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$
CMS+HE $\frac{d(e^{\epsilon}+1)}{e^{\epsilon}-1}$ $\frac{d(e^{\epsilon}+1)^{2}}{(e^{\epsilon}-1)^{2}}$ $(\frac{e^{\epsilon}+1}{e^{\epsilon}-1})^{2}$
SS $\sqrt{\frac{8}{\pi}}\frac{de^{\epsilon/2}}{e^{\epsilon}-1}$ $\frac{4de^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ $(\frac{e^{\epsilon}+1}{e^{\epsilon}-1})^{2}$
a-RP $\frac{2de^{\epsilon/2}}{e^{\epsilon}-1}$ $\frac{4de^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ $(\frac{e^{\epsilon}+1}{e^{\epsilon}-1})^{2}$
s-RP $\frac{de^{\epsilon/4}}{e^{\epsilon/2}-1}$ $\frac{de^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$ $\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$
TABLE I: Accuracy comparison among frequency-estimation algorithms
  • Note: HE, RHR, OLH, OCMS, CMS+HE, SS, a-RP, and s-RP represent Hadamard Encoding, Recursive Hadamard Response, Optimized Local Hashing, Optimized CMS, CMS with Hadamard Encoding, Subset Selection, asymmetric RAPPOR, and symmetric RAPPOR, respectively. Here, OCMS represents MSE-OCMS and $l$-OCMS when comparing $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively. The data source is presented in Section H.
Communication Cost (bits)
HE $\log d$
RHR $\log d+\epsilon$
OLH $d\epsilon$
MSE-OCMS $2\log d+\frac{\epsilon}{2}$
$l$-OCMS $2\log d+\epsilon$
original CMS $\Omega(d\log m)$
SS $\frac{d}{1+e^{\epsilon}}$
RAPPOR $d$
TABLE II: Communication Cost among frequency-estimation algorithms
  • Note: see Table I for the meaning of abbreviations. The communication cost only accounts for the cost between the server and one client. The data source is presented in Section H.
Table I shows that the optimized CMS (OCMS) is optimal in $l_{2}$ loss and the worst-case MSE, even considering the constant term. Its $l_{1}$ loss is order-optimal. Since HE, OLH, and CMS+HE are equivalent to unbiased CMSs without the optimized parameters, it is expected that they are not optimal in at least one dimension. Only OCMS and s-RP are optimal in the worst-case MSE, but Table II shows that s-RP requires $d$ bits to communicate while OCMS only requires $2\log_{2}d+\frac{\epsilon}{2}$ bits. If the communication cost is required to be $O(\log d)$, only RHR has the same order of accuracy as OCMS in $l_{1}$ and $l_{2}$ losses. However, empirical study (Section 6) demonstrates that OCMS outperforms RHR in all datasets, consistent with the theoretical analysis in Section 4.3. Therefore, OCMS is the only algorithm optimized for the worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss while communicating efficiently.
5.2 Necessity of Randomness
[11] proposed an algorithm to convert the Hadamard Response (a special case of CMS) into a deterministic algorithm, arguing that the randomness could be avoided, thereby decreasing the communication cost. However, we will show that one can always create a counterexample causing the frequency estimation to have a constant bias. Assume that we are interested in the frequency of value $x$, and $f(x)$ is set to zero (no $x$ in the dataset). Given any hash function (denoted as $h_{j}$ for convenience), we denote $x_{j}$ as the value colliding with $x$, i.e., $c_{j}(x,x_{j})=1$. Subsequently, we set the value of all objects assigned to the $j$-th hash function as $x_{j}$, and repeat this process for all the hash functions. As a result, we have:
$\sum_{j\in[k]}\sum_{x^{\prime}\in[d]\backslash x}c_{j}(X^{(i)},x)f(x^{\prime})=1.$
Substituting the above formula and $f(x)=0$ into Eq. (3) results in $E[\hat{f}(x)]=1$, which proves:
Theorem 5.1.
If the assignment of hash functions to the objects, i.e., $\forall i:h_{j^{(i)}}$, is constant in CMS, one can always construct a counterexample making
$E[\hat{f}(x)]-f(x)\neq 0.$
The logic behind the theorem is intuitive: CMS utilizes randomness to ensure a stable probability of collision with other values. If one gives up the randomness, an attacker can always generate a dataset to maximize the collision. The reason why [11] believed a deterministic algorithm still worked was that they assumed that each $X^{(i)}$ was sampled from a distribution, so one cannot control the value of $X^{(i)}$. Their algorithm relies on the randomness in distribution sampling instead of the randomness in assigning hash functions.
The next question is how much randomness (in bits) is needed to ensure that CMS will work as expected. Here, we focus only on the randomness in assigning hash functions and assume that the LDP process $R$ always uses its own randomness.
In Section 3.2, the result of $Var(\hat{f}(x))$ is built on $Cov(\hat{y}^{(i_{1})}[h_{j^{(i_{1})}}(x)],\hat{y}^{(i_{2})}[h_{j^{(i_{2})}}(x)])=0$ for any $i_{1}$ and $i_{2}$, which is further derived from the assumption that the hash-function assignments $\mathbf{1}(j^{(i_{1})}=j)$ and $\mathbf{1}(j^{(i_{2})}=j)$ are pairwise independent.
[…]
Theorem 5.2.
$\lceil\max\{4\log_{2}d,2\log_{2}n\}\rceil$ random bits are enough to generate all the hash-function assignments $[j^{(i)}]_{i\in[n]}$ while ensuring their pairwise independence, which is the prerequisite of Eq. (4).
5.3 Public or Private Randomness
Previous literature introduces the concept of public and private randomness as follows:
Public Randomness: the server generates a string of random bytes, which are accessible by the public (i.e., clients).
Private Randomness: the client generates a string of random bytes, which will be used in their algorithm. The client may or may not publish the random bytes.
Previous literature (e.g., [5, 14, 12]) had a strong interest in public randomness, because it was believed that public randomness could save communication costs. However, we argue that public and private randomness make no difference for CMS, and no communication cost can be saved, for the following reasons: the server and a client (i.e., an object in CMS) have to sync on which hash function is assigned to the client, and they need to either sync (1) the identifier of a hash function, or (2) the assignment generator. There are at least $d^{2}$ hash functions in $\mathcal{H}$, so identifying one of them requires $2\log_{2}d$ bits. Meanwhile, Theorem 5.2 shows that at least $4\log_{2}d$ bits are required to represent the assignment generator. Thus, syncing the identifier will be less expensive, and $2\log_{2}d$ bits are needed to sync, regardless of which side generates the assignment. Additionally, $\log_{2}n$ bits are required if the server and a client decide to sync on the assignment generator, because the client needs to send back the object ID (i.e., $i\in[n]$) to identify itself.
When reviewing the existing literature, the work claiming one-bit communication usually missed the communication cost of syncing the public randomness. For instance, [5] proposed a protocol to convert a private-randomness algorithm to a public one. Though the client only sends one bit as a response, it has to sync $b$ bits from the server, where $b$ is the number of bits the private-randomness algorithm would send. Moreover, the server also needs the client/object ID to know where the one bit comes from, which will contribute $\log n$ bits to the communication cost. The same problem also exists in the public-randomness algorithm in [14].
The only scenario where public randomness could save communication costs is when all the clients are co-located, referred to as a client group. The traffic inside a client group is assumed to be free. When utilizing the public randomness, the server will share the hash-function assignment generator with the client group, which costs $\lceil\max\{4\log_{2}d,2\log_{2}n\}\rceil$ bits. Then, the client group shares the generator, calculates their assignment of the hash functions, and replies with their responses in sequence back to the server, which will cost $n\log m\approx n\epsilon$ bits. In this case, the total communication cost between the server and the client group is $\lceil\max\{4\log_{2}d,2\log_{2}n\}\rceil+n\epsilon$. If the client group still uses private randomness, they will return $n(\log d+\epsilon)$ bits, which is larger than the cost from public randomness. However, the client-group assumption is contradictory to the motivation of LDP, which usually assumes the clients are distributed.
Figure 2: Worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss vs. privacy factor $\epsilon$ given the Zipf dataset. See Section 6.1 for details.
Figure 3: Worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss vs. privacy factor $\epsilon$ given the Gaussian dataset. See Section 6.2 for details.
Figure 4: Worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss vs. privacy factor $\epsilon$ given the mini Kosarak dataset. See Section 6.3 for details.
6 Experiment
This section empirically compares the optimized CMS with the algorithms listed below, using both synthetic and real-world datasets:
1. Hadamard Encoding (HE): the first practical frequency estimation algorithm.
2. CMS with HE (CMS+HE): the original CMS algorithm with the same parameter as [7].
3. Recursive Hadamard Response (RHR): the state-of-the-art version of the Hadamard Response family.
Other algorithms listed in Section 5.1 are not included due to their very high communication cost, which is impractical with a very large dictionary.
In the following experiments, we use this estimator to estimate $\widehat{MSE}$:
$\overline{MSE}(\hat{f})=\max_{x\in\mathbf{x}}\frac{1}{t}\sum_{t}(\hat{f}(x)-f(x))^{2},$ (23)
where $\mathbf{x}$ denotes the values of interest and $t$ denotes the experiment rounds. However, $\overline{MSE}(\hat{f})$ is biased because $\frac{1}{t}\sum_{t}(\hat{f}(x)-f(x))^{2}$ of each $x$ is equivalent to a random variable fluctuating around $E[(\hat{f}(x)-f(x))^{2}]$, and the max operator will choose the largest deviated value. Appendix I provides an upper bound of $\overline{MSE}(\hat{f})$. As long as $\overline{MSE}(\hat{f})$ is between the upper bound and the calculated $\widehat{MSE}(\hat{f})$ (both bounds are plotted), we will consider the experimental $\overline{MSE}(\hat{f})$ consistent with the theoretical value from Theorem 15.
Moreover, we will compute $l_{1}$ / $l_{2}$ losses only from a subset of $[d]$, denoted as $\mathbf{x}$. By examining Eq. (48), one could prove that the $l_{1}$ / $l_{2}$ losses from Theorem 3.7 are still valid when replacing $d$ with $|\mathbf{x}|$. Note that the calculated $l_{1}$ loss is an upper bound, so the experimental value is expected to be below it.
All the experiments below will be run 100 times for each privacy factor $\epsilon$ ranging from 1 to 5.
6.1 Verification of Optimality and Equivalence
We use a synthetic dataset sampled from a specific Zipf’s distribution, which has a frequency distribution as
$Zipf(r)\propto\frac{1}{r^{2}},$
where $r$ is the order of a value ranked by its frequency. The dataset values are chosen deliberately, so all have the same mod for the Recursive Hadamard Response (RHR) (see Section 4.3). The dictionary size is set to 100k, and 10,000 objects are sampled from the distribution. The $\widehat{MSE}$, $l_{1}$ loss, and $l_{2}$ loss are evaluated based on the first 100 most frequent values (assuming they are known in advance). MSE-OCMS with $f^{*}=1$ (i.e., no prior knowledge on $f(x)$) and $l$-OCMS are evaluated for $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, while RHR, HE, and CMS+HE are evaluated for all the metrics. Results are plotted in Fig. 4.
MSE-OCMS and $l$-OCMS outperform other algorithms in terms of $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, and their accuracy aligns with the calculated values. HE and CMS+HE have identical performance in all the metrics, supporting our theory that these two algorithms are equivalent. Consistent with our expectation, RHR performs the worst in all the metrics because we deliberately construct a dataset, though using only public information, to ensure every object has a 50% probability of colliding with another in RHR, resulting in poorer performance than HE, as analyzed in Section 4.3.
6.2 Synthetic Dataset with Prior Knowledge
A synthetic database is constructed with 10,000 objects sampled from a normal distribution with a standard deviation of 50. The mean of the distribution is an unknown integer picked from [1000, 9000]. The dictionary size is 10,000, so the values of all the objects are rounded to integers and truncated if they are outside of [0, 10,000]. Through several simulations of this distribution, we are confident that the probability of $\max_{x}f(x)\leq 0.01$ is high. Thus, we set $f^{*}=0$ for MSE-OCMS and verify if its $\widehat{MSE}$ satisfies Eq. (16). Meanwhile, $l$-OCMS is evaluated for $l_{1}$ / $l_{2}$ losses. RHR, HE, and CMS+HE are also evaluated for $\widehat{MSE}$ $l_{1}$ loss, and $l_{2}$ loss. Results are plotted in Fig. 4.
Similar to the above, MSE-OCMS and $l$-OCMS outperform other algorithms in terms of $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, and their accuracy aligns with the calculated values. HE and CMS+HE have identical accuracy in all the metrics, and they perform the worst. The accuracy of RHR lies between OCMS and the other algorithms.
6.3 Real-world Dataset
Kosarak [22] is a dataset containing click-stream data from a Hungarian online news portal. The dataset is shrunk to 1% of its original size, denoted as mini Kosarak, which will be used in this section. It contains over 170,000 entries, all of which belong to a dictionary with a size of 26,000. MSE-OCMS with $f^{*}=1$ and $l$-OCMS are evaluated for $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, while RHR, HE, and CMS+HE are evaluated for all the metrics. Results are plotted in Fig. 4.
MSE-OCMS and $l$-OCMS outperform other algorithms in terms of $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, and their accuracy aligns with the calculated values. HE and CMS+HE have identical accuracy in all the metrics, and they perform the worst in $l_{1}$ / $l_{2}$ losses. RHR has the same order of accuracy as $l$-OCMS regarding $l_{1}$ / $l_{2}$ losses. However, it performs the worst in $\widehat{MSE}$. This is expected because Kosarak has one value whose frequency is above 0.4, which negatively impacts the MSE of other values sharing the same mod with it (see Section 4.3 for a detailed explanation).
7 Conclusion
This paper revisits the private Count-mean Sketch (CMS) algorithm, corrects the errors in the expectation and variance calculations in the original CMS paper, and derives the optimized parameters to reduce all kinds of loss functions. The optimized CMS is proven theoretically and empirically to be the only algorithm optimized in the worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss, while its communication cost is logarithmic to the dictionary size. Moreover, this paper demonstrates that many state-of-the-art algorithms are equivalent to CMS. Finally, we discuss the necessity of randomness in CMS and show that both public and private randomness have the same communication cost.
Appendix A Derivation of the Expectation
Let’s focus on the term $\hat{y}^{(i)}[h_{j^{(i)}}(x)]$ in the definition of $f(x)$ in Eq. (2), which can be expanded as
$\hat{y}^{(i)}[h_{j^{(i)}}(x)]=R(h_{j^{(i)}}(X^{(i)}))[h_{j^{(i)}}(x)]\\ =\sum_{j\in[k]}\mathbf{1}\{j=j^{(i)}\}R(h_{j}(X^{(i)}))[h_{j}(x)],$ (24)
where $\mathbf{1}\{*\}$ returns one if * is true. Given that $h_{j^{(i)}}$ is uniformly sampled from $\mathcal{H}$, we have
$E[\hat{y}^{(i)}[h_{j^{(i)}}(x)]]\\ =\sum_{j\in[k]}E\bigl{[}\mathbf{1}\{j=j^{(i)}\}R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{]}\\ =\sum_{j\in[k]}\frac{1}{k}c_{j}(X^{(i)},x),$ (25)
where $c_{j}(X^{(i)},x)$ indicates whether $X^{(i)}$ and $x$ collide in the hash function $h_{j}$, and $E\bigl{[}R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{]}=c_{j}(X^{(i)},x)$ is derived from Property 2.1. Substituting the above equation into Eq. (2), we have
$E[\hat{f}(x)]=\frac{m}{n(m-1)}\sum_{i\in[n]}E\bigl{[}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{]}-\frac{1}{m-1}\\ =\frac{m}{n(m-1)}\sum_{i\in[n]}\sum_{j\in[k]}\frac{c_{j}(X^{(i)},x)}{k}-\frac{1}{m-1}\\ =\frac{m}{m-1}[f(x)+\sum_{x^{\prime}\in[d]\backslash x}\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}f(x^{\prime})]-\frac{1}{m-1},$ (26)
and Theorem 3.2 is proved.
Substituting $\forall x,x^{\prime}:\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}=\frac{1}{m}$ into the above equation, we have
$E[\hat{f}(x)]=\frac{m}{m-1}[f(x)+\sum_{x^{\prime}\in[d]\backslash x}\frac{1}{m}f(x^{\prime})]-\frac{1}{m-1}\\ =\frac{m}{m-1}[f(x)+(1-f(x))\frac{1}{m}]-\frac{1}{m-1}=f(x).$ (27)
Regarding its inverse, whose condition is that there exists an $x^{\prime}$ such that $\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}\neq\frac{1}{m}$. If we construct a dataset containing only $x^{\prime}$ and $x$, then we have
$E[\hat{f}(x)]=\frac{m}{m-1}[f(x)+(1-f(x))\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}]\\ -\frac{1}{m-1}\neq f(x),$ (28)
so Corollary 3.1 is proved.
Appendix B Derivation of the Variance
Let’s start with the variance of an individual response $\hat{y}^{(i)}[h_{j^{(i)}}(x)]$:
$Var\bigl{(}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{)}\\ =E\bigl{[}\bigl{(}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{)}^{2}\bigr{]}-\bar{c}(X^{(i)},x)^{2}$ (29)
where $E\bigl{[}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{]}=\bar{c}(X^{(i)},x)$ is proven in Eq. (25). Then,
$E\bigl{[}\bigl{(}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{)}^{2}\bigr{]}\\ =\sum_{j\in[k]}E\bigl{[}\bigl{(}\mathbf{1}\{j=j^{(i)}\}R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{)}^{2}\bigr{]}\\ =\sum_{j\in[k]}E[\mathbf{1}\{j=j^{(i)}\}]E\bigl{[}\bigl{(}R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{)}^{2}\bigr{]}\\ =\frac{1}{k}\sum_{j\in[k]}E\bigl{[}\bigl{(}R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{)}^{2}\bigr{]}\\ \stackrel{{\scriptstyle\text{(a)}}}{{=}}\frac{1}{k}\sum_{j\in[k]}\biggl{(}Var\bigl{(}R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{)}+c_{j}(X^{(i)},x)\biggr{)}$ (30)
where =$\stackrel{{\scriptstyle\text{(a)}}}{{=}}$ holds because $E\bigl{[}R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{]}^{2}=c_{j}(X^{(i)},x)^{2}=c_{j}(X^{(i)},x)$. Integrating the above equation into Eq. (29), we have
$Var\bigl{(}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{)}=\frac{1}{k}\sum_{j\in[k]}Var\bigl{(}R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{)}\\ +\bar{c}(X^{(i)},x)-\bar{c}(X^{(i)},x)^{2}.$ (31)
Now, we will study the variance of $\sum_{i\in[n]}\hat{y}^{(i)}[h_{j^{(i)}}(x)]$:
$Var\bigl{(}\sum_{i\in[n]}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{)}=\sum_{i\in[n]}Var\bigl{(}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{)}\\ +\sum_{i_{1}\neq i_{2}}Cov\bigl{(}\hat{y}^{(i_{1})}[h_{j^{(i_{1})}}(x)],\hat{y}^{(i_{2})}[h_{j^{(i_{2})}}(x)]\bigr{)}.$ (32)
Focusing on the covariance and expanding $\hat{y}^{(i)}[h_{j^{(i)}}(x)]$ as Eq. (24), we have
$Cov\bigl{(}\hat{y}^{(i_{1})}[h_{j^{(i_{1})}}(x)],\hat{y}^{(i_{2})}[h_{j^{(i_{2})}}(x)]\bigr{)}=\\ Cov\bigl{(}\sum_{j\in[k]}\mathbf{1}\{j=j^{(i_{1})}\}R(h_{j}(X^{(i_{1})}))[h_{j}(x)],\\ \sum_{j\in[k]}\mathbf{1}\{j=j^{(i_{2})}\}R(h_{j}(X^{(i_{2})}))[h_{j}(x)]\bigr{)}.$ (33)
Given that $Cov(X,Y)=E[XY]-E[X]E[Y]$, we calculate:
$E\bigl{[}\sum_{j_{1}\in[k]}\sum_{j_{2}\in[k]}\mathbf{1}\{j_{1}=j^{(i_{1})}\}\mathbf{1}\{j_{2}=j^{(i_{2})}\}\\ R(h_{j_{1}}(X^{(i_{1})}))[h_{j_{1}}(x)]R(h_{j_{2}}(X^{(i_{2})}))[h_{j_{2}}(x)]\bigr{]}\\ \stackrel{{\scriptstyle\text{(a)}}}{{=}}\sum_{j_{1}\in[k]}\sum_{j_{2}\in[k]}E[\mathbf{1}\{j_{1}=j^{(i_{1})}\}]E[\mathbf{1}\{j_{2}=j^{(i_{2})}\}]\times\\ E[R(h_{j_{1}}(X^{(i_{1})}))[h_{j_{1}}(x)]]E[R(h_{j_{2}}(X^{(i_{2})}))[h_{j_{2}}(x)]]\\ =\sum_{j_{1}\in[k]}\sum_{j_{2}\in[k]}\frac{1}{k^{2}}c_{j_{1}}(X^{(i_{1})},x)c_{j_{2}}(X^{(i_{2})},x)\\ =\bar{c}(X^{(i_{1})},x)\bar{c}(X^{(i_{2})},x),$ (34)
where =$\stackrel{{\scriptstyle\text{(a)}}}{{=}}$ holds because the assignment of a hash function is simply a uniform sampling, which is independent of everything, and $R(h_{j_{1}}(X^{(i_{1})}))[h_{j_{1}}(x)]$ is also independent of $R(h_{j_{2}}(X^{(i_{2})}))[h_{j_{2}}(x)]$ given Property 2.2.
Utilizing Eq. (25) and (34), we have
$Cov\bigl{(}\hat{y}^{(i_{1})}[h_{j^{(i_{1})}}(x)],\hat{y}^{(i_{2})}[h_{j^{(i_{2})}}(x)]\bigr{)}\\ =\bar{c}(X^{(i_{1})},x)\bar{c}(X^{(i_{2})},x)-\bar{c}(X^{(i_{1})},x)\bar{c}(X^{(i_{2})},x)=0.$ (35)
As a result, we have
[…]
and subsequently, we have
$Var(\hat{f}(x))=\frac{m^{2}}{(m-1)^{2}n}\sum_{x^{\prime}\in[d]}\biggl{[}\biggl{(}\frac{1}{k}\sum_{j\in[k]}\\ Var\bigl{(}R(h_{j}(x^{\prime}))[h_{j}(x)]\bigr{)}\biggr{)}+\bar{c}(x^{\prime},x)-\bar{c}(x^{\prime},x)^{2}\biggr{]}f(x^{\prime}),$ (37)
which proves Eq. (4).
When considering Property 2.3, we have $Var\bigl{(}R(h_{j}(x^{\prime}))[h_{j}(x)]\bigr{)}=Var(R|=)$ if $h_{j}(x^{\prime})=h_{j}(x)$. Otherwise, it equals $Var(R|\neq)$. When $x^{\prime}=x$, they will collide in every hash function, so $\bar{c}(x^{\prime},x)=1$ and $Var\bigl{(}R(h_{j}(x^{\prime}))[h_{j}(x)]\bigr{)}=Var(R|=)$ in this case. On the other hand, when $x^{\prime}\neq x$, we have
$\frac{1}{k}\sum_{j\in[k]}Var\bigl{(}R(h_{j}(x^{\prime}))[h_{j}(x)]\bigr{)}\\ =\frac{1}{k}\sum_{j\in[k]}[c_{j}(x^{\prime},x)Var(R|=)+(1-c_{j}(x^{\prime},x))Var(R|\neq)]\\ =\bar{c}(x^{\prime},x)Var(R|=)+(1-\bar{c}(x^{\prime},x))Var(R|\neq).$ (38)
Therefore, when substituting Eq. (38) for $x^{\prime}\neq x$ and the aforementioned parameters with $x=x^{\prime}$ into Eq. (4), we have
$Var(\hat{f}(x))=\frac{m^{2}}{(m-1)^{2}n}\biggl{[}\\ \sum_{x^{\prime}\in[d]\backslash x}f(x^{\prime})\biggl{(}\bar{c}(x^{\prime},x)Var(R|=)+(1-\bar{c}(x^{\prime},x))\\ Var(R|\neq)+\bar{c}(x^{\prime},x)(1-\bar{c}(x^{\prime},x))\biggr{)}+f(x)Var(R|=)\biggr{]}.$ (39)
Thus, Eq. 5 is proved.
If $\bar{c}(x^{\prime},x)=\frac{1}{m}$, it becomes
$Var(\hat{f}(x))=\frac{m}{(m-1)^{2}n}\biggl{[}\\ \sum_{x^{\prime}\in[d]\backslash x}f(x)\biggl{(}Var(R|=)+(m-1)\\ Var(R|\neq)+\frac{1-m}{m}\biggr{)}+mf(x)Var(R|=)\biggr{]}\\ =\frac{m}{(m-1)^{2}n}\biggl{[}\bigl{(}1-f(x)\big{)}\biggl{(}Var(R|=)+(m-1)\\ Var(R|\neq)+\frac{1-m}{m}\biggr{)}+mf(x)Var(R|=)\biggr{]},$ (40)
which proves Eq. (6).
Appendix C Proof of Theorem 3.5
Recall the definition of $\hat{f}(x)$ in Eq. (2), which is equivalent to the sum of bounded random variables. Define $Y^{(i)}=\frac{m}{m-1}\hat{y}^{(i)}[h_{j}^{(i)}(x)]-\frac{1}{m-1}$, and we have $n\hat{f}(x)=\sum_{i}Y^{(i)}$. Note that $\frac{Y^{(i)}-A}{B-A}$ is a Bernoulli random variable where
$A=\min\hat{f}(x)=\frac{m}{m-1}\frac{e^{-1}}{e^{\epsilon}-1}-\frac{1}{m-1}$
$B=\max\hat{f}(x)=\frac{m}{m-1}\frac{e^{\epsilon}+m-2}{e^{\epsilon}-1}-\frac{1}{m-1}.$
Here, we assume that CMS uses RR to perturb the hashed values. Also note that we only consider unbiased CMS, i.e., $E[\hat{f}(x)]=f(x)$. Using the Chernoff (upper) bound, we have $\forall\delta\in[0,1]$
$Pr[\sum_{i}\frac{Y^{(i)}-A}{B-A}\geq(1+\delta)n\frac{f(x)-A}{B-A}]\\ \leq\exp(-\frac{n}{3}\frac{f(x)-A}{B-A}\delta^{2}).$ (41)
Define $\delta=\frac{\alpha\sqrt{Var(\hat{f}(x))}}{f(x)-A}$. Given that $\forall\delta\in[0,1]$, it is equivalent to requiring $\alpha\in[0,\sqrt{\frac{ne^{\epsilon}}{m-1}}]$. Thus, Eq. (41) can be rewritten as
$Pr[\hat{f}(x)\geq f(x)+\alpha\sqrt{Var(\hat{f}(x))}]\\ \leq\exp(-\frac{n}{3}\frac{Var(\hat{f}(x))}{(B-A)(f(x)-A)}\alpha^{2})$ (42)
$Var(\hat{f}(x))$ is derived in Eq. (45), and $\frac{Var(\hat{f}(x))}{(B-A)(f(x)-A)}$ decreases with $f(x)$ (one can verify it using derivatives). Calculating the minimum at $f(x)=1$, we have
$\frac{Var(\hat{f}(x))}{(B-A)(f(x)-A)}\geq\frac{m-1}{e^{\epsilon}+m-1}.$
As a result,
$Pr[\hat{f}(x)\geq f(x)+\alpha\sqrt{Var(\hat{f}(x))}]\\ \leq\exp(-\frac{n}{3}\frac{m-1}{e^{\epsilon}+m-1}\alpha^{2}).$ (43)
The same proof is also applicable to $\hat{f}(x)\leq f(x)-\alpha\sqrt{Var(\hat{f}(x))}$. Therefore,
$Pr[|\hat{f}(x)-f(x)|\geq+\alpha\sqrt{Var(\hat{f}(x))}]\\ \leq 2\exp(-\frac{n}{3}\frac{m-1}{e^{\epsilon}+m-1}\alpha^{2}).\qed$ (44)
Appendix D Proof regarding Optimized CMS
We first explore the optimized $m$ for the CMS+RR. Substituting Eq. (12) into Eq. (6), we have
$Var(\hat{f}(x)_{RR})=\frac{1-f(x)}{n(m-1)}+\\ \frac{m[(1-f(x))(e^{\epsilon}-1)(2-m)+me^{\epsilon}]}{n(m-1)(e^{\epsilon}-1)^{2}}$ (45)
Let’s first study the worst-case MSE of CMS+RR, which is equivalent to the worst-case variance $Var(\hat{f}(x)_{RR})$. Observe that the variance is linear to $f(x)$, we have $\max_{x}MSE(\hat{f}(x))=\max\{Var(\hat{f}(x)_{RR}|f(x)=0),Var(\hat{f}(x)_{RR}|f(x)=1)\}$. When $f(x)=1$, we have
$\frac{\partial Var(\hat{f}(x)_{RR}|f(x)=1)}{\partial f(x)}=\frac{m(m-2)e^{\epsilon}}{n(m-1)^{2}(e^{\epsilon}-1)^{2}},$
implying that $Var(\hat{f}(x)_{RR}|f(x)=1)$ increases monotonically with $m$ when $m\geq 2$. On the other hand, when $f(x)=0$, we have
$\frac{\partial Var(\hat{f}(x)_{RR}|f(x)=0)}{\partial f(x)}=\frac{(m-e^{\epsilon}-1)(m+e^{\epsilon}-1)}{n(m-1)(e^{\epsilon}-1)^{2}},$
which indicates a minimum at $m=e^{\epsilon}+1$ given the range $m\geq 2$. However, when $m=e^{\epsilon}+1$, $Var(\hat{f}(x)_{RR}|f(x)=1)$ is always larger than $Var(\hat{f}(x)_{RR}|f(x)=0)$. Given that $Var(\hat{f}(x)_{RR}|f(x)=1)$ decreases as $m$ decreases, while $Var(\hat{f}(x)_{RR}|f(x)=0)$ increases when $m$ decreases within the range $[2,e^{\epsilon}+1]$, there exists an $m$ making them equal to each other. Solving
$Var(\hat{f}(x)_{RR}|f(x)=0)=Var(\hat{f}(x)_{RR}|f(x)=1),$
we have $m=e^{\epsilon/2}+1$, and the corresponding variance is
$Var(\hat{f}(x)_{RR})=\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}.$
When $12\leq m<e^{\epsilon/2}+$ 1, we have $Var(\hat{f}(x)_{RR}|f(x)=1)<Var(\hat{f}(x)_{RR}|f(x)=0)$. Also, $Var(\hat{f}(x)_{RR}|f(x)=0)$ decreases with $m$ in this range. Therefore, when $12\leq m\leq e^{\epsilon/2}+$ 1, we have
$\max_{f(x)}Var(\hat{f}(x)_{RR})>\max_{f(x)}Var(\hat{f}(x)_{RR}|m=e^{\epsilon/2}+1)\\ =\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}.$
That is, $m$ should never be smaller than $e^{\epsilon/2}+1$.
Define $f^{*}$ to be the upper bound of $f(x)$, i.e., $\forall x:f(x)\leq f^{*}$. When $e^{\epsilon/2}+1\leq m\leq e^{\epsilon}+1$, we have proven that $Var(\hat{f}(x)_{RR})$ increases with $f(x)$, so the goal becomes reducing $Var(\hat{f}(x)_{RR}|f(x)=f^{*})$. The optimized $m$ can be calculated by
$\frac{\partial Var(\hat{f}(x)_{RR}|f(x)=f^{*})}{\partial m}=0,$
which results in
$m=1+\frac{\Delta_{MSE}}{f^{*}e^{\epsilon}+1-f^{*}},$ (46)
where
$\Delta_{MSE}=e^{\epsilon/2}\sqrt{[(1-f^{*})e^{\epsilon}+f^{*}][f^{*}e^{\epsilon}+(1-f^{*})]}.$
The corresponding variance is
$Var^{*}(\hat{f}(x)_{RR})=\frac{2(\Delta_{MSE}+e^{\epsilon})}{n(e^{\epsilon}-1)^{2}}$
Notice that Eq. (46) is valid only when $e^{\epsilon/2}+1\leq m\leq e^{\epsilon}+1$, which translates to $f^{*}\leq\frac{1}{2}$. When $f^{*}>\frac{1}{2}$, we have $m<1+e^{\epsilon/2}$, which has been proven to be worse than $m=1+e^{\epsilon/2}$. Therefore, we have
$m=\begin{cases}1+\Delta_{MSE}&\text{ if }f^{*}\leq\frac{1}{2}\\ 1+e^{\epsilon/2}&\text{ if }\frac{1}{2}<f^{*}\leq 1,\end{cases}$
which corresponds to
$Var^{*}(\hat{f}(x)_{RR})=\begin{cases}\frac{2(\Delta_{MSE}+e^{\epsilon})}{n(e^{\epsilon}-1)^{2}}&\text{ if }f^{*}\leq\frac{1}{2}\\ \frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}&\text{ if }\frac{1}{2}<f^{*}\leq 1\end{cases}.$
When it comes to symmetry and asymmetry RAPPOR, since their $Var(a)[a]$ and $Var(a)[b]$ are independent of $m$, the corresponding $Var(\hat{f}(x))$ decreases when $m$ increases. Thus, when $m$ is large enough, we have
$Var(\hat{f}(x)_{RP})\rightarrow\frac{1}{n}[Var(R|\neq)(1-f(x))+\\ Var(R|=)f(x)].$ (47)
For asymmetric RAPPOR, we realize $Var(\hat{f}(x)_{aRP})>Var^{*}(\hat{f}(x)_{RR})$ when $f^{*}>0$, so the CMS with asymmetric RAPPOR cannot serve as the optimized CMS. For symmetric RAPPOR, $Var(\hat{f}(x)_{sRP})>Var^{*}(\hat{f}(x)_{RR})$ when $0\leq f^{*}<\frac{1}{2}$, and $Var(\hat{f}(x)_{sRP})=Var^{*}(\hat{f}(x)_{RR})$ when $\frac{1}{2}\leq f^{*}\leq 1$. Thus, the CMS with symmetric RAPPOR is also ruled out. As a result, Theorem 15 is proved. Corollary 3.4 is proved automatically when substituting $f^{*}=1$ into Theorem 15.
Now, we will explore the optimized parameters for $l_{2}$ loss. By considering only the pairwise independence of $\mathcal{H}$, we integrate Eq. (6) into Eq. (11):
$l_{2}(\hat{f})=\sum_{x\in[d]}\frac{m}{n(m-1)^{2}}\bigl{[}\bigl{(}Var(R|=)+(m-1)Var(R|\neq)\\ +\frac{m-1}{m}\bigr{)}(1-f(x))+mVar(R|=)f(x)\bigr{]}\\ =\frac{m}{n(m-1)^{2}}\bigl{[}\bigl{(}Var(R|=)+(m-1)Var(R|\neq)\\ +\frac{m-1}{m}\bigr{)}(d-1)+mVar(R|=)\bigr{]}$ (48)
We will first use RR to perturb the hashed values. Substituting Eq. (12) into Eq. (48), we obtain
$l_{2}(\hat{f}_{RR})=\frac{(d+e^{\epsilon}-1)m^{2}}{n(m-1)(e^{\epsilon}-1)^{2}}+\\ \frac{(d-1)(2m+e^{\epsilon}-1)}{n(m-1)(e^{\epsilon}-1)}.$ (49)
By studying its derivative with respect to $m$, we can find its minimum when
$m=1+\frac{\Delta_{l}}{e^{\epsilon}+d-1},$
where
$\Delta_{l}=\sqrt{(e^{\epsilon}+d-1)(de^{\epsilon}-e^{\epsilon}+1)}e^{\epsilon/2}.$
In practice, we choose the closest integer to $1+\frac{\Delta_{l}}{e^{\epsilon}+d-1}$ as $m$. The corresponding minimum value is
$\min_{m}l_{2}(\hat{f}_{RR})=\frac{2(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}}.$ (50)
If the size of candidates $d$ is large enough, we have
$l_{2}(\hat{f}_{RR})\rightarrow\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}},$ (51)
and the corresponding $m\rightarrow 1+e^{\epsilon}$.
Similar to the worst-case MSE, we could prove that the CMS with symmetric RAPPOR will always have larger $l_{2}$ loss than the CMS+RR. For asymmetric RAPPOR, its $l_{2}$ loss is the same as that of CMS+RR only when $d$ is large enough. However, its communication cost is too high because RAPPOR requires sending a vector of $m$ bits to the server, and Eq. (47) requires $m\rightarrow\infty$. To determine the necessary value of $m$, we solve the following equation for asymmetric RAPPOR:
$l_{2}(\hat{f}_{aRP})=dVar(\hat{f}(x)_{aRP})=(1+\tau)\frac{4de^{\epsilon}}{(e^{\epsilon}-1)^{2}},$
where $\tau$ is a small number like 0.01. $m$ is solved as
$m=\Omega(\frac{(e^{\epsilon}-1)^{2}}{(1+\tau)e^{\epsilon}}).$ (52)
On the other hand, CMS+RR only requires $\log_{2}(1+e^{\epsilon})$ bits to output the perturbed result. Thus, CMS+RR is always preferred when optimizing $l_{2}$ loss due to its low communication cost. Now, we have proven the $l_{2}$-loss part in Theorem 3.7 and Corollary 3.5.
For $l_{1}$ loss, given the Cauchy-Schwarz inequality, we have $l_{1}\leq\sqrt{d\>l_{2}}$. Combining this with Eq. (50) leads to
$l_{1}(\hat{f})\leq\sqrt{\frac{2d(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}}}.$ (53)
Define $\lceil l_{1}\rceil$ to be the upper bound of $l_{1}$. If the size of candidates $d$ is large enough, we have
$\lceil l_{1}(\hat{f})\rceil\rightarrow\frac{2de^{\epsilon/2}}{\sqrt{n}e^{\epsilon}-1}.$ (54)
Similar to $l_{2}$ loss, the corresponding $m\rightarrow 1+e^{\epsilon}$ (in practice, we choose the closest integer). Thus, the $l_{1}$-loss part in Theorem 3.7 and Corollary 3.5 are proven.
Appendix E Proof regarding Imperfect Hashing
Let’s start by proving Theorem 3.8. Observe $\frac{\partial Var(g(x\mid\mathcal{H}^{\prime}))}{\partial m^{\prime}}$, and note that it is linear to $f(x)$. Considering the derivative at $f(x)=1$ and $f(x)=0$ yields:
$-\max\{\frac{2m^{\prime}Var(R|=)}{n(m^{\prime}-1)^{3}},\\ \frac{(m^{\prime}-1)(Var(R|\neq)+1)+(m^{\prime}+1)Var(R|=)}{n(m^{\prime}-1)^{3}}\}\\ \leq\frac{\partial Var(g(x|\mathcal{H}^{\prime}))}{\partial m^{\prime}}<0.$ (55)
Observing that $Var(R|=)\geq Var(R|\neq)$ is satisfied by the LDP mechanism such as randomized response and RAPPOR, we have
$-\frac{2m^{\prime}Var(R|=)+m^{\prime}-1}{n(m^{\prime}-1)^{3}}\leq\frac{\partial Var(g(x|\mathcal{H}^{\prime}))}{\partial m^{\prime}}<0.$ (56)
At the same time, we have $m^{\prime}=m-\frac{m}{(2q+1)^{2}}$. Denote $Var(\hat{f}(x))$ as the variance of the CMS using perfect hashing. We want $Var(g(x|\mathcal{H}^{\prime}))$ to approach $Var(\hat{f}(x))$, which is formulated as $Var(g(x|\mathcal{H}^{\prime}))\leq Var(\hat{f}(x))(1+\tau)$, where $\tau$ is a small number like 0.01. This can be translated to
$\frac{2mVar(R|=)+m-1}{n(m-1)^{3}}\frac{m}{(2q+1)^{2}}\leq\tau Var(\hat{f}(x)),$
which can be organized as
$\frac{2mVar(R|=)+m-1}{n(m-1)^{3}}\frac{m}{\tau Var(\hat{f}(x))}\leq(2q+1)^{2},$ (57)
thus proving Theorem 3.8. When $m=1+e^{\epsilon/2}$, $Var(\hat{f}(x))$ becomes $\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$, and the left-hand side of Eq. (57) decreases with $\epsilon$. Thus, we can substitute $\epsilon=0$ into Eq. (57), and we have $2q+1>\sqrt{\frac{1}{\tau}}$. Corollary 3.6 is proved.
If $\forall x:Var(g(x))\leq(1+\tau)Var(\hat{f}(x))$, then $l_{2}(g)\leq(1+\tau)l_{2}(\hat{f}(x))$. When $m=1+e^{\epsilon}$, $Var(\hat{f}(x))$ increases with $f(x)$, so we have $Var(\hat{f}(x)|f(x)=0)\leq Var(\hat{f}(x))$. Thus, if
$\frac{2mVar(R|=)+m-1}{n(m-1)^{3}}\frac{m}{\tau Var(\hat{f}(x)|f(x)=0)}\\ \leq(2q+1)^{2}$ (58)
is satisfied, Eq. (57) will also be satisfied. Substituting $\frac{4e^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ into $Var(\hat{f}(x)|f(x)=0)$ in Eq. (58) and setting $m=1+e^{\epsilon}$, we also have the left-hand side of Eq. (58) decreasing with $\epsilon$. Thus, this inequality is still satisfied when $\epsilon=0$, which derives $2q+1>\sqrt{\frac{1}{\tau}}$. Thus, Corollary 3.7 is proved.
Now we will prove Theorem 3.9. First, by calculating the derivative of $Pr(h^{\prime}(x_{1})=h^{\prime}(x_{2}))$ with respect to $r$ to find the maximum, we have
$Pr(h^{\prime}(x_{1})=h^{\prime}(x_{2}))\leq\frac{1}{m}+\frac{1}{4mq(q+1)},$ (59)
which is equivalent to $\bar{c}(x,x^{\prime})\leq\frac{1}{m}+\frac{1}{4mq(q+1)}$.
Next, consider the $MSE(\hat{f}(x))$ in Eq. (7). We aim for $MSE(\hat{f}(x))$ to approach $Var(\hat{f}(x))$, where $\hat{f}(x)$ denotes the estimation with perfect pairwise independent hashing. Regarding the bias term of $MSE(\hat{f}(x))$, we have
$E[\hat{f}(x|\mathcal{H}^{\prime})]-f(x)=\frac{m}{m-1}(\bar{c}(x,x^{\prime})-\frac{1}{m})\sum_{x^{\prime}\in[d]\backslash x}f(x^{\prime})\\ \leq\frac{m}{m-1}\frac{1}{4mq(q+1)}=\frac{1}{4(m-1)q(q+1)}.$ (60)
Another term in $MSE(\hat{f}(x))$ is $Var(\hat{f}(x))$, which is linear to $\bar{c}(x,x^{\prime})$ and $\bar{c}(x,x^{\prime})^{2}$, and all the parameters of these two terms are positive. Therefore,
$\frac{Var[\hat{f}(x|\mathcal{H}^{\prime})]}{Var[\hat{f}(x)]}\leq\max\{\frac{\bar{c}(x,x^{\prime})}{1/m},\frac{\bar{c}(x,x^{\prime})^{2}}{1/m^{2}}\}\\ =\frac{(2q+1)^{4}}{16q^{2}(q+1)^{2}}.$ (61)
To satisfy
$MSE(f(x)|\mathcal{H}^{\prime})-MSE(\hat{f}(x))\leq 2\tau Var[\hat{f}(x)],$
it is equivalent to requiring
$(E[\hat{f}(x|\mathcal{H}^{\prime})]-f(x))^{2}<\tau Var[\hat{f}(x)]$
$\frac{Var[\hat{f}(x|\mathcal{H}^{\prime})]}{Var[\hat{f}(x)]}<1+\tau$
Substituting Eqs. (60) and (61) into the above equations proves Theorem 3.9.
Appendix F Proof regarding Correcting the Original CMS
F.1 Proof of Theorem 3.10
Consider each $c_{j}(x,x^{\prime})$ as an independent Bernoulli random variable with probability $\frac{1}{m}$ of being one. Given that $E[c_{j}(x,x^{\prime})]=\frac{1}{m}$ and $Var(c_{j}(x,x^{\prime}))=\frac{m-1}{m}$, we have
[…]
and
$\mathop{Var}_{\forall c_{j}(x,x^{\prime})}[E[\hat{f}(x)]]=\frac{m^{2}}{(m-1)^{2}}\biggl{(}\sum_{x^{\prime}\in[d]\backslash x}\frac{m-1}{m^{2}k}f(x^{\prime})^{2}\biggr{)}\\ =\frac{1}{(m-1)k}\sum_{x^{\prime}\in[d]\backslash x}f(x^{\prime})^{2}.$ (63)
Thus, Theorem 3.10 is proved.
F.2 Proof of Theorem 22
Given Markov’s inequality, Eq. (22) is equivalent to requiring
$\frac{\mathop{Var}_{\forall c_{j}(x,x^{\prime})}(E[\hat{f}(x)])}{\tau Var(\hat{f}(x))}\leq\delta.$ (64)
Theorem 3.10 has demonstrated that
$\mathop{Var}_{\forall c_{j}(x,x^{\prime})}(E[\hat{f}(x)])\leq\frac{1}{(m-1)k}$ (65)
The original CMS uses RAPPOR or Hadamard Encoding (HE) to perturb the hash values in Step 2, so their variance can be calculated using Eq. (5), which can be scaled down as
$Var(\hat{f}(x))>\frac{1}{n}\min\{Var(R|=),Var(R|\neq)\}\\ =\frac{Var(R|\neq)}{n},$ (66)
given that $Var(R|=)>Var(R|\neq)$ in both RAPPOR and HE. Combining Eq. (65) and (66), we have
$\frac{\mathop{Var}_{\forall c_{j}(x,x^{\prime})}(E[\hat{f}(x)])}{Var(\hat{f}(x))}<\frac{n}{(m-1)kVar(R|\neq)}.$
Substituting the above inequality into Eq. (64), the requirement of Eq. (22) is converted as
$\frac{n}{(m-1)kVar(R|\neq)}<\delta,$ (67)
which is the same as the condition of Theorem 22.
Appendix G Decoding Hadamard Encoding
Define $H$ to be a scaled Walsh Hadamard matrix, where $H[i,j]=(-1)^{i\cdot j}$ with $\cdot$ representing bitwise multiplication. Hadamard encoding employs a LDP mechanism to perturb $H[X^{(i)}+1,j^{(i)}]$ of each object, where $X^{(i)}$ and $j^{(i)}$ are as defined in Section 3. [5] has proved that
$\hat{f}_{H}(x)=\sum_{i\in[n]}\hat{H}[X^{(i)}+1,j^{(i)}]H[j^{(i)},x+1]$
unbiasedly estimates the frequency of $f(x)$. Here, $\hat{H}[X^{(i)}+1,j^{(i)}]=R(H[X^{(i)}+1,j^{(i)}])$, with $R$ denoting the LDP reconstruction process as detailed in Section 2.2.
Interpreting $H[j^{(i)},:]$ as a hash function, $\hat{H}[X^{(i)}+1,j^{(i)}]H[j^{(i)},x+1]$ yields +1 if both $X^{(i)}$ and $x$ hash to the same value, and -1 otherwise. This behavior is analogous to $\hat{y}^{(i)}[h_{j^{i}}(x)]$ in Eq. (2). However, $\hat{y}^{(i)}[h_{j^{i}}(x)]$ returns zero when $X^{(i)}$ and $x$ hash to different values. By considering the constants $\frac{m}{m-1}$ and $\frac{1}{m-1}$ in Eq. (2), and setting $m=2$, we derive $12\hat{y}^{(i)}[h_{j^{i}}(x)]-$ 1, which produces results identical to $\hat{H}[X^{(i)}+1,j^{(i)}]H[j^{(i)},x+1]$. Consequently, the decoding process for Hadamard encoding is equivalent to Eq. (2).
Appendix H Data Source of Section 5.1
We start with the accuracy. Since HE is equivalent to unbiased CMS+RR with $m=2$, we use Eq. (45) to compute its $\widehat{MSE}$, $l_{1}$ loss, and $l_{2}$ losses. The computed $\widehat{MSE}$ is the same as in [11]. The $l_{1}$ and $l_{2}$ losses of RHR were taken from [12], but it did not calculate the $\widehat{MSE}$. Section 4.3 shows that RHR is a skewed version of unbiased CMS, and the $\widehat{MSE}$ of RHR could be worse than that from the CMS+RR with $m=2$ (i.e., HE), so the $\widehat{MSE}$ of RHR is $\Omega\left(\left(\frac{e^{\epsilon}+1}{e^{\epsilon}-1}\right)^{2}\right)$. OLH is equivalent to the unbiased CMS with $m=1+e^{\epsilon}$, so we use Eq. (45) and follow Section 3.4 to calculate its $l_{1}$ loss, $l_{2}$ loss, and $\widehat{MSE}$. Section 4.5 shows that CMS+HE is equivalent to the unbiased CMS+RR with $m=2$ (i.e., HE), so its $l_{1}$ loss, $l_{2}$ loss, and $\widehat{MSE}$ are the same as HE. The $l_{1}$ and $l_{2}$ losses of SS are presented in [10]. Its $\widehat{MSE}$ is derived using Eq. (2) in [9] with
$p=\frac{ke^{\epsilon}}{ke^{\epsilon}+d-k}$
$q=\frac{ke^{\epsilon}}{ke^{\epsilon}+d-k}\frac{k-1}{d-1}+\frac{d-k}{ke^{\epsilon}+d-k}\frac{k}{d-1}$
$k=d/(e^{\epsilon}+1)$
$f=1$
where $k$ represents the size of a subset in SS. $f=1$ because the $Var(\hat{f}(x))$ increases with $f(x)$ for SS.
The $\widehat{MSE}$ of a-RP and s-RP is calculated using Eq. (2) in [9] as well. Both cases where $f(x)=0$ and $f(x)=1$ are calculated, and the worst-case variance is the maximum between them.
Next, we present the data source regarding communication cost. We first compute the cost of OCMS, which requires the server and a client to sync on the assigned hash function and the perturbed response $z^{(i)}$. Section 3.5 shows that a hash function is comprised of two random numbers, each ranging from 1 to $\Theta(d)$, so $2\log_{2}d$ bits are sufficient to represent one hash function. Then, we need another $\log_{2}m$ bits to sync the perturbed response $z^{(i)}$. Given that OCMS has either $m=1+e^{\epsilon/2}$ or $m=1+e^{\epsilon}$ depending on the optimization goal, the bits to transfer $z^{(i)}$ are close to $\epsilon/2$ and $\epsilon$, respectively. Thus, the overall communication cost of MSE-OCMS and $l$-OCMS is $22\log_{2}d+\epsilon/$ 2 and $2\log_{2}d+\epsilon$, respectively.
The costs of HE and RHR are taken from their respective literature [11] and [12]. It is notable that we do not use the communication cost from their public-randomness scheme because they fail to consider the cost of syncing the public randomness, as pointed out in Section 5.3. For OLH, its hash function maps $[d]$ to $[1+e^{\epsilon}]$, and each mapping is mutually independent, so it requires $d\log_{2}(1+e^{\epsilon})\approx d\epsilon$ bits to represent all the mappings. The original CMS [7] did not consider pairwise independence, so its hash function mapping from $[d]$ to $[m]$ requires $d\log_{2}m$ bits to communicate. SS randomly selects $\frac{d}{e^{\epsilon}+1}$ values from $[d]$, and this subset selection requires $\frac{d}{e^{\epsilon}+1}$ bits to represent. RAPPOR generates a one-hot vector with size $d$, which requires $d$ bits to transfer.
Appendix I Upper Bound of the worst-case MSE estimator
The upper bound of $\overline{MSE}(\hat{f})$ is set to be $[1+\frac{2}{t}\bigl{(}\sqrt{t\log(20|\mathbf{x}|)}+\log(20|\mathbf{x}|)\bigr{)}]V$ based on the following theorem:
Theorem I.1.
If $\forall x\in\mathbf{x}:E[(\hat{f}(x)-f(x))^{2}]\leq V$, where $V$ is constant, we have
$Pr\bigl{(}\overline{MSE}(\hat{f})\geq[1+\frac{2}{t}\bigl{(}\sqrt{t\log(20|\mathbf{x}|)}+\\ \log(20|\mathbf{x}|)\bigr{)}]V\bigr{)}\leq 0.05,$
where $t$ is the experiment rounds.
Proof: Similar to Appendix C, we define $Y^{(i)}=\frac{m}{m-1}\hat{y}^{(i)}[h_{j}^{(i)}(x)]-\frac{1}{m-1}$, and $\frac{Y^{(i)}-A}{B-A}$ is a Bernoulli random variable (see Appendix C for the definition of $A$ and $B$). All the $X^{(i)}$ can be placed into two groups based on whether $X^{(i)}=x$. If $X^{(i)}=x$, $\frac{Y^{(i)}-A}{B-A}\sim\text{Ber}\left(\frac{e^{\epsilon}}{e^{\epsilon}+m-1}\right)$, where $\text{Ber}(p)$ denotes the Bernoulli distribution with probability $p$. On the other hand, if $X^{(i)}\neq x$, $\frac{Y^{(i)}-A}{B-A}\sim\text{Ber}\left(\frac{1}{m}\frac{e^{\epsilon}}{e^{\epsilon}+m-1}+\frac{m-1}{m}\frac{1}{e^{\epsilon}+m-1}\right)$. For convenience, we denote $p_{1}=\frac{e^{\epsilon}}{e^{\epsilon}+m-1}$ and $p_{2}=\frac{1}{m}\frac{e^{\epsilon}}{e^{\epsilon}+m-1}+\frac{m-1}{m}\frac{1}{e^{\epsilon}+m-1}$.
Subsequently, $\frac{n(\hat{f}(x)-A)}{B-A}$ is equivalent to the sum of two binomial random variables, where the first is sampled from $\text{BN}(nf(x),p_{1})$ and the second is sampled from $\text{BN}(n(1-f(x)),p_{2})$, where $\text{BN}(n,p)$ denotes the binomial distribution with $n$ trials each having a success probability $p$.
When $n$ is large enough, both binomial random variables can be approximated as Gaussian random variables, which are sampled from $\mathcal{N}(nf(x)p_{1},np_{1}(1-p_{1}))$ and $\mathcal{N}(n(1-f(x))p_{2},np_{2}(1-p_{2}))$ respectively, where $\mathcal{N}(\mu,\sigma^{2})$ denotes a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$. Since the sum of these two Gaussian random variables is also a Gaussian random variable, $\hat{f}(x)$ is approximated as
$\frac{n(\hat{f}(x)-A)}{B-A}\sim\mathcal{N}(nf(x)p_{1}+n(1-f(x))p_{2},\\ nf(x)p_{1}(1-p_{1})+n(1-f(x))p_{2}(1-p_{2})).$
One can verify that the above equation is equivalent to
$\frac{n(\hat{f}(x)-A)}{B-A}\sim\mathcal{N}(\frac{n(f(x)-A)}{B-A},\frac{n^{2}Var(\hat{f}(x))}{(B-A)^{2}}),$
which can be refactored as
$\frac{\hat{f}(x)-f(x)}{\sqrt{Var(\hat{f}(x))}}\sim\mathcal{N}(0,1).$
Thus, $\frac{\hat{f}(x)-f(x)}{\sqrt{Var(\hat{f}(x))}}$ of each experiment can be considered a standard Gaussian random variable. There are $t$ rounds of the experiment; using the Laurent-Massart bound [23], we have
$Pr\biggl{(}\sum_{t}[(\frac{\hat{f}(x)_{t}-f(x)}{\sqrt{Var(\hat{f}(x))}})^{2}-1]\\ \geq 2(\sqrt{t\alpha}+\alpha)\biggr{)}\leq e^{-\alpha}.$
Define $S(x)=\frac{1}{t}(\hat{f}(x)_{t}-f(x))^{2}$, and we have
$Pr\biggl{(}S(x)\geq[1+\frac{2}{t}(\sqrt{t\alpha}+\alpha)]Var(\hat{f}(x))\biggr{)}\leq e^{-\alpha}.$
Given our assumption that $E[(\hat{f}(x)-f(x))^{2}]=Var(\hat{f}(x))\leq V$ is identical for $\forall x\in\mathbf{x}$, we have
$Pr\biggl{(}S(x)\geq[1+\frac{2}{t}(\sqrt{t\alpha}+\alpha)]V\biggr{)}\leq e^{-\alpha}.$
Note that $\forall x:S(x)<(1+\frac{2}{t}(\sqrt{t\alpha}+\alpha))V$ is equivalent to $\max_{x}S(x)<(1+\frac{2}{t}(\sqrt{t\alpha}+\alpha))V$. Given that $\overline{MSE}(\hat{f})=\max_{x}S(x)$, we have
$Pr\bigl{(}\overline{MSE}(\hat{f})<[1+\frac{2}{t}(\sqrt{t\alpha}+\alpha)]V\bigr{)}>(1-e^{-\alpha})^{|\mathbf{x}|},$ (68)
which decreases with $|\mathbf{x}|$. Even when $|\mathbf{x}|\rightarrow\infty$, if $e^{-\alpha}|\mathbf{x}|=\frac{1}{20}$, we still have $(1-e^{-\alpha})^{|\mathbf{x}|}\approx 0.95$. Thus, we have $\alpha=\log(20|\mathbf{x}|)$. Then, we obtain
$Pr\bigl{(}\overline{MSE}(\hat{f})<[1+\frac{2}{t}\bigl{(}\sqrt{t\log(20|\mathbf{x}|)}+\\ \log(20|\mathbf{x}|)\bigr{)}]V\bigr{)}>0.95.\qed$

### v2

@@
-Count-mean Sketch as an Optimized Framework for Frequency Estimation with Local Differential Privacy
+Improving Count-Mean Sketch as the Leading Locally Differentially Private Frequency Estimator for Large Dictionaries
 Abstract
-This paper identifies that a group of state-of-the-art locally-differentially-private (LDP) algorithms for frequency estimation are equivalent to the private Count-Mean Sketch (CMS) algorithm with different parameters.
-Therefore, we revisit the private CMS, correct errors in the original CMS paper regarding expectation and variance, modify the CMS implementation to eliminate existing bias, and explore optimized parameters for CMS to achieve optimality in reducing the worst-case mean squared error (MSE), $l_{1}$ loss, and $l_{2}$ loss.
+This paper identifies that a group of latest locally-differentially-private (LDP) algorithms for frequency estimation, including all the Hadamard-matrix-based algorithms, are equivalent to the private Count-Mean Sketch (CMS) algorithm with different parameters.
+Therefore, we revisit the private CMS, correct errors in the original CMS paper regarding expectation and variance, modify the CMS implementation to eliminate existing bias, and optimize CMS using randomized response (RR) as the perturbation method.
+The optimized CMS with RR is shown to outperform CMS variants with other known perturbations in reducing the worst-case mean squared error (MSE), $l_{1}$ loss, and $l_{2}$ loss.
 Additionally, we prove that pairwise-independent hashing is sufficient for CMS, reducing its communication cost to the logarithm of the cardinality of all possible values (i.e., a dictionary).
-As a result, the aforementioned optimized CMS is proven theoretically and empirically to be the only algorithm optimized for reducing the worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss when dealing with a very large dictionary.
+As a result, the optimized CMS with RR is proven theoretically and empirically as the leading algorithm for reducing the aforementioned loss functions when dealing with a very large dictionary.
 Furthermore, we demonstrate that randomness is necessary to ensure the correctness of CMS, and the communication cost of CMS, though low, is unavoidable despite the randomness being public or private.
@@
 However, frequency calculation becomes challenging when data is protected, and their values are not directly accessible to the analyst.
-One of the most popular approaches to protect data is differential privacy [1], which is considered the de facto standard for extracting statistics from a dataset while protecting individual data privacy.
-Differential privacy has a local version [2], known as Local Differential Privacy (LDP), which requires enough perturbation on individual data such that an analyst or attacker cannot gain significant knowledge about the individual data values despite their prior knowledge [3].
+One of the most popular approaches to protect data is differential privacy (DP) [1], which is considered the de facto standard for extracting statistics from a dataset while preserving individual data privacy.
+DP introduces sufficient noise to the statistical outputs such that the participation of any individual’s data cannot be confidently inferred from the query results.
+However, DP still requires individuals to upload their data to a server before extracting statistics, posing risks of eavesdropping during data transmission or data breaches on the server side [2].
+To address these concerns, Local Differential Privacy (LDP) [3] has been proposed, which requires enough perturbation on individual data on the client side before sending them to the server for statistical analysis, theoretically preventing eavesdropping or data breaches.
 However, LDP also adds more challenges to calculating statistics from a dataset, including frequency estimation.
-A major challenge of LDP-preserving frequency estimation is that traditional mechanisms, such as randomized response [4], result in the variance of the estimated frequency increasing linearly with the number of all possible values in a dataset (referred to as a dictionary), making the estimation infeasible if the dictionary is too large.
-Several algorithms [5, 6, 7, 8, 9, 10, 11, 12, 13] have been developed to address this problem, which will be discussed in the Previous Work (Section 1.1).
+A major challenge of LDP-preserving frequency estimation is that traditional mechanisms, such as randomized response [4], result in the precision of the estimated frequency increasing linearly with the number of all possible values in a dataset (referred to as a dictionary), making the estimation infeasible if the dictionary is too large.
+Several algorithms [5, 6, 7, 8, 9, 10, 11, 12, 13] have been developed to address this problem, and their precision is proven to be independent of the dictionary size, though some of them still have communication cost linear to the dictionary size.
+These algorithms will be further introduced in the Previous Work (Section 1.1).
 Many of these algorithms (e.g., [5, 9, 14, 12]) are found to be equivalent to the Count-Mean Sketch (CMS) algorithm (see Section 4 for more details).
@@
 The LDP version of CMS was originally developed by [7], but it had minor issues in calculating the expectation and variance (see Section 3.7) and did not explore the optimized parameters for CMS.
-Therefore, our work provides a corrected calculation of expectation and variance (Sections 3.1 and 3.2) and explores the optimized parameters of CMS (Section 3.4), which optimally reduce the worst-case mean squared error ($\widehat{MSE}$), $l_{1}$ loss, and $l_{2}$ loss as follows:
-$\widehat{MSE}^{*}=\begin{cases}\frac{2(\Delta_{MSE}+e^{\epsilon})}{n(e^{\epsilon}-1)^{2}}&\text{ if }f^{*}\leq\frac{1}{2}\\ \frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}&\text{ if }\frac{1}{2}<f^{*}\leq 1\end{cases}$
-$l_{1}^{*}\leq\sqrt{\frac{2d(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}}}\xrightarrow{d\rightarrow\infty}\frac{2de^{\epsilon/2}}{\sqrt{n}(e^{\epsilon}-1)}$
-$l_{2}^{*}=\frac{2(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}}\xrightarrow{d\rightarrow\infty}\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$
-with
-$\Delta_{MSE}=e^{\epsilon/2}\sqrt{[(1-f^{*})e^{\epsilon}+f^{*}][f^{*}e^{\epsilon}+(1-f^{*})]}$
-$\Delta_{l}=e^{\epsilon/2}\sqrt{(e^{\epsilon}+d-1)(de^{\epsilon}-e^{\epsilon}+1)}$
-where $\epsilon$, $n$, and $d$ are the privacy factor, dataset size, and dictionary size, respectively; $f^{*}$ is defined as $\forall x\in\mathbf{x}:f(x)\leq f^{*}$ where $\mathbf{x}$ are the values of interest; * denotes the optimality.
+Therefore, our work provides a corrected calculation of expectation and variance (Sections 3.1 and 3.2) and optimize CMS using randomized response as the perturbation method (referred to as OCMS+RR).
+OCMS+RR is shown to outperform CMS variants with other known perturbations in reducing the worst-case mean squared error ($\widehat{MSE}$), $l_{1}$ loss, and $l_{2}$ loss (Section 3.4).
 Moreover, our work proves that pairwise-independent hash functions are sufficient for CMS, and Section 3.5 proposes a practical approach to construct a pairwise-independent hashing family given any range of a hash function.
-The pairwise independence can reduce the communication cost to the logarithm of the dictionary size.
-As a result, Section 5.1 compares various state-of-the-art algorithms and concludes that our optimized CMS is the only algorithm optimized in the worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss while its communication cost is only logarithmic relative to the dictionary size.
-Additionally, Section 5.2 proves the necessity of randomness in CMS and provides a counterexample against the existing deterministic algorithm.
+The pairwise independence can reduce the communication cost to the logarithm of the dictionary size, allowing OCMS+RR to be applied to datasets with very large dictionaries.
+The pseudocode of OCMS+RR is presented in Section 3.6 (source code available at https://github.com/mingen-pan/CountMeanSketch).
+When the dictionary size is much larger than the privacy factor $\epsilon$ defined by LDP, the aforementioned loss functions (defined in Section 2.2) are reduced by OCMS+RR as
+$\widehat{MSE}\leq\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}$
+$l_{1}\xrightarrow{d\gg e^{\epsilon}}\frac{2de^{\epsilon/2}}{\sqrt{n}(e^{\epsilon}-1)}$
+$l_{2}\xrightarrow{d\gg e^{\epsilon}}\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$
+where $n$, and $d$ are the dataset size and dictionary size, respectively.
+As a result, Section 5.1 compares existing algorithms with OCMS+RR and concludes that our OCMS+RR is the leading algorithm for reducing the worst-case MSE ($\widehat{MSE}$), $l_{1}$ loss, and $l_{2}$ loss, while its communication cost is only logarithmic relative to the dictionary size.
+Additionally, Section 5.2 proves the necessity of randomness in CMS and provides a counterexample to an attempt to de-randomize a CMequivalent algorithm by [14].
 Section 5.3 shows that the communication cost of CMS is unavoidable, despite the randomness being public or private.
-Section 6 empirically demonstrates that the optimized CMS outperforms other algorithms.
+Section 6 empirically demonstrates that OCMS+RR outperforms other algorithms.
 1.1 Previous Work
-RAPPOR [6] is the first algorithm to estimate frequency with accuracy independent of dictionary size.
+RAPPOR [6] is the first algorithm to estimate frequency with precision independent of dictionary size.
 It represents the original value as a one-hot vector and applies LDP perturbation to every bit of the array.
 However, the communication cost is proportional to the dictionary size, as the one-hot vector is the same size as the dictionary.
-[15] developed another version of RAPPOR, which hashes an original value to a smaller range than the dictionary size, such that the one-hot encoding results in a much smaller vector.
+An alternative version of RAPPOR [15] hashes an original value into a smaller range than the dictionary size, resulting in a much smaller one-hot encoded vector.
 This version uses gradient descent to search for one possible solution for the original frequencies.
 However, this mapping is equivalent to dimension reduction in linear algebra, so the found solution is not guaranteed to be close to the original frequencies.
-To tackle the large dictionary, [5] developed an algorithm that transforms an original value to a column of a Hadamard matrix, uniformly samples one element from the column, and perturbs it.
+To tackle the large dictionary, an algorithm was developed by [5], which transforms an original value to a column of a Hadamard matrix, uniformly samples one element from the column, and perturbs it.
 The original frequencies can be reconstructed from the transformed values by applying the Hadamard matrix again.
-The accuracy of this approach is independent of the dictionary size, and the communication cost is the logarithm of the dictionary size.
+The precision of this approach is independent of the dictionary size, and the communication cost is the logarithm of the dictionary size.
 The Apple Privacy team [7] modified the Count-mean Sketch (CMS) algorithm to be LDP.
@@
 In addition to CMS, other members of the Count Sketch family have also been adapted to LDP (e.g., [8, 13]).
+Wang et al.
 [9] summarized existing approaches and proposed a variant of RAPPOR and a local hashing method, which were claimed to be optimal in terms of MSE.
 However, [9] assumed the estimated frequencies to be always close to zero, so their optimality did not hold when the original frequency approaches one.
-At the same time, [10] proposed a Subset Selection algorithm to estimate frequency with optimized accuracy in $l_{1}$ / $l_{2}$ losses, but its communication cost is also proportional to the dictionary size when the privacy requirement is high.
+Simultaneously, the Subset Selection algorithm [10] was proposed to estimate frequency with leading precision in $l_{1}$ / $l_{2}$ losses, but its communication cost is also proportional to the dictionary size when the privacy requirement is high.
+Acharya et al.
 [11] later integrated the Hadamard matrix into the subset selection and reduced the communication cost to the logarithm of the dictionary size.
-[12] further improved this method by applying the Hadamard transform twice.
-Though its performance can achieve order-optimal in $l_{1}$ / $l_{2}$ losses, it may perform worse than the Hadamard Response in terms of MSE (as discussed in Section 4.3).
+This method was further improved by [12], which applies the Hadamard transform twice.
+Though its performance can achieve the state-of-the-art order of precision in $l_{1}$ / $l_{2}$ losses, it may perform worse than the Hadamard Response in terms of MSE (as discussed in Section 4.3).
 2 Background
@@
 Local Differential Privacy (LDP) was initially proposed in [3] to ensure that querying an object does not significantly disclose its actual value.
+Consider a variable $V$ that is perturbed by a mechanism $M$:
 Definiton 2.1.
-A mechanism $M$ is $\epsilon$-LDP if and only if $\forall(v,v^{\prime})\in\mathcal{V}^{2}$, and $\forall u\in\mathcal{U}$,
-$Pr(M(v)=u|v)\leq e^{\epsilon}Pr(M(v^{\prime})=u|v^{\prime}).$
-While there are other variations of LDP, this paper focuses solely on the most widely-used form mentioned above.
-Given the above definition, $P$ is defined as a $|\mathcal{U}|\times|\mathcal{V}|$ matrix, where $P[u,v]$ represents the probability of the LDP mechanism $M$ generating $u$ given the input $v$.
+$M$ is $\epsilon$-LDP if and only if $\forall v,v^{\prime}\in\mathcal{V}$, and $\forall u\in\mathcal{U}$,
+$Pr(M(V)=u|V=v)\leq e^{\epsilon}Pr(M(V)=u|V=v^{\prime}),$
+where $\mathcal{V}$ and $\mathcal{U}$ represent the dictionary (all possible values) of $V$ and the range of $M$, respectively.
+This paper may abbreviate $\left(M\left(V\right)|V=v\right)$ as $M(v)$ when the variable is not the focus and its value is known.
+Given the above definition, $P$ is defined as a $|\mathcal{U}|\times|\mathcal{V}|$ matrix, where $P[u,v]=Pr(M(v)=u)$.
 Here $u$ and $v$ are represented by an index between $1..|\mathcal{U}|$ and $1..|\mathcal{V}|$, respectively.
-Suppose there exists another matrix $Q$ of shape $|\mathcal{V}|\times|\mathcal{U}|$, which satisfies $I=QP$ and $I$ is an identity matrix.
-The decoding process of $M$ is defined to output $Q[:,u]$ given the output $u$ from $M$.
-The reconstruction process (perturbation and decoding) is denoted as $R$, whose output is an estimator of the input $v$, formulated as:
+We consider only cases where $|\mathcal{U}|\geq|\mathcal{V}|$ and $P$ is full rank, i.e., $\text{rank}(P)=|\mathcal{V}|$.
+This condition should hold for all meaningful LDP mechanisms, as the expected frequencies of the perturbed values correspond to a linear transformation of the original frequencies, and a rank-deficient transformation would prevent the reconstruction of the original frequencies from the perturbed values.
+Widely-used algorithms (e.g., [4, 6]) satisfy this condition.
+Subsequently, a matrix $Q$ is constructed as $(P^{T}P)^{-1}P^{T}$, which satisfies $I=QP$ with $I$ being an identity matrix.
+Given an output $u$ from $M$, its decoding is defined to output $Q[:,u]$, i.e., the $u$-th column of $Q$.
+Thus, the reconstruction process (perturbation and decoding) of a variable $V$ is defined as $R(V)=Q[:,M(V)]$.
+Similarly, $\left(R\left(V\right)|V=v\right)$ is abbreviated as $R(v)$ when the variable is not the focus and its value is known.
+An example of the reconstruction is presented in Appendix A.
+The randomness of $R(v)$ comes from $M(v)$, and each possible value of $R(v)$ is a vector of size $|\mathcal{V}|$ satisfying:
 Property 2.1.
 $\begin{gathered}E[R(v)[v]]=1\\ \forall v^{\prime}\neq v:E[R(v)[v^{\prime}]]=0\end{gathered}$
+where $E$ denotes expectation.
 Proof: represent any $u$ from $M(v)$ as a one-hot vector of size $|\mathcal{U}|$, where the $u$-th element is one.
@@
 Meanwhile, we require the LDP mechanism $M$ to be stateless, meaning that its output depends only on its input. This is formulated as
-$Pr(M(v)=u|v)=Pr(M(v)=u|v,\theta),$
-where $\theta$ represents any other information (except the randomness generator used by $M$). Therefore, the outputs of $M$ from two inputs are mutually independent, which is formulated as:
+$Pr(M(V)=u|V=v)=Pr(M(V)=u|V=v,\theta),$
+where $\theta$ represents any other information (except the randomness generator used by $M$).
+Considering two variables $V$ and $V^{\prime}$, whose values are determined prior to perturbation, the stateless requirement ensures that their reconstruction process $R(V)$ and $R(V^{\prime})$ are mutually independent random variables, formulated as
 Property 2.2.
-If $v$ and $v^{\prime}$ are two inputs into $M$, the reconstruction of them from the outputs of $M$ satisfies $\forall(a,b)\in\mathcal{V}^{2}:Cov(R(v)[a]),R(v^{\prime})[b])=0$.
-Most widely-used algorithms (e.g., randomized response [4] and RAPPOR [6]) satisfy this property. Additionally, we observe that these LDP mechanisms also exhibit symmetry in variance, formulated as:
+The covariance of $R(V)$ and $R(V^{\prime})$ satisfies $\forall v,v^{\prime},v_{a},v_{b}\in\mathcal{V}:Cov(R(V)[v_{a}],R(V^{\prime})[v_{b}]|V=v,V^{\prime}=v^{\prime})=0$.
+Widely-used algorithms (e.g., [4, 6]) satisfy this property. Additionally, we observe that these LDP mechanisms also exhibit symmetry in variance, denoted as $Var$, formulated as:
 Property 2.3.
-When $\forall v,v^{\prime}\in\mathcal{V}^{2}$ and $v\neq v^{\prime}$, $Var(R(v)[v^{\prime}])$ are identical, denoted as $Var(R|\neq)$; $\forall v\in\mathcal{V}$, $Var(R(v)[v])$ are also identical, denoted as $Var(R|=)$.
+Every element of $\{Var(R(v)[v^{\prime}])|v,v^{\prime}\in\mathcal{V},v\neq v^{\prime}\}$ has an identical value, denoted as $Var(R|\neq)$; every element of $\{Var(R(v)[v])|v\in\mathcal{V}\}$ has an identical value, denoted as $Var(R|=)$.
 2.2 Frequency Estimation
-Let’s define some symbols: $X^{(i)}$ is the $i$-th object of an arbitrary dataset of size $n$; the possible values of all the objects comprise an dictionary $[d]$, where $d$ is its size, and $\forall i\in[n]:X^{(i)}\in[d]$.
+Let’s define some symbols: $X^{(i)}$ is the $i$-th object of an arbitrary dataset of size $n$; the possible values of all the objects comprise a dictionary $[d]$, where $d$ is its size, and $\forall i\in[n]:X^{(i)}\in[d]$.
 The frequency of value $x$ is defined as $f(x)=\frac{\sum_{i\in[n]}\mathbf{1}\{X^{(i)}=x\}}{n}$.
-The LDP workflow of frequency estimation usually proceeds as follows:
-1. Each object $X^{(i)}$ is perturbed by an LDP mechanism, which outputs $Y^{(i)}$.
-2. The server collects and aggregates the outputs as $\hat{f}(x)=A_{x}([Y^{(i)}]_{i\in[n]})$, if the server is interested in $f(x)$. The aggregation can be run for every $x\in[d]$.
-$\hat{f}(x)$ is an estimator of $f(x)$. Existing literature usually uses $l_{1}$ and $l_{2}$ losses to measure its overall accuracy, which are defined as follows:
+The LDP workflow of frequency estimation typically proceeds as follows:
+1. Each object $X^{(i)}$ is perturbed by an LDP mechanism, which outputs $\tilde{X}^{(i)}$.
+2.
+An analyst collects and aggregates the outputs as $\hat{f}(x)=A_{x}([\tilde{X}^{(i)}]_{i\in[n]})$, if the analyst is interested in $f(x)$, where $A_{x}$ denotes the aggregation.
+The aggregation can be performed for every $x\in[d]$.
+A potential scenario involves a group of clients sending their perturbed responses to a server, while their unperturbed responses constitute a dataset.
+The server then estimates the frequency of the unperturbed responses.
+$\hat{f}(x)$ in Step 2 is an estimator of $f(x)$. Existing literature usually uses $l_{1}$ and $l_{2}$ losses to measure its overall precision, which are defined as follows:
 $\begin{gathered}l_{1}(\hat{f})=\max_{\forall\mathbf{X}:|\mathbf{X}|=n}E[\sum_{x\in[d]}|\hat{f}(x)-f(x)|]\\ l_{2}(\hat{f})=\max_{\forall\mathbf{X}:|\mathbf{X}|=n}E[\sum_{x\in[d]}(\hat{f}(x)-f(x))^{2}]\end{gathered}$
 where $\mathbf{X}=[X^{(i)}]_{i\in[n]}$, and $\forall\mathbf{X}$ assumes the size $n$ is constant.
-These two losses measure the overall accuracy of the estimation.
-However, in many cases, an analyst is also interested in the accuracy of the frequency estimation of an individual value.
-The mean squared error (MSE) is commonly used to measure the accuracy of an estimation, reflecting both the variance and bias of the estimation.
+These two losses measure the overall precision of the estimation.
+However, in many cases, an analyst is also interested in the precision of the frequency estimation of an individual value.
+The mean squared error (MSE), formulated as $MSE(\hat{f}(x))=E[(\hat{f}(x)-f(x))^{2}]$, is commonly used to measure the precision of an estimation, reflecting both the variance and bias of the estimation.
 Since an analyst is assumed to have no knowledge of the actual dataset, they need to assume the worst case when calculating the MSE, which is defined as
-$\widehat{MSE}(\hat{f})=\max_{\forall\mathbf{X}:|\mathbf{X}|=n}\max_{x\in\mathbf{x}}E[(\hat{f}(x)-f(x))^{2}],$ (1)
+$\widehat{MSE}(\hat{f})=\max_{\forall\mathbf{X}:|\mathbf{X}|=n}\max_{x\in\mathbf{x}}MSE(\hat{f}(x)),$ (1)
 where $\widehat{MSE}$ denotes the upper bound of MSE, i.e., the worst case; $\mathbf{x}$ denotes the values of interest, which is $[d]$ by default.
-If the estimator $\hat{f}$ is unbiased, the worse-case MSE is the same as the worst-case variance, which is also positively correlated to the confidence interval (CI).
+If the estimator $\hat{f}$ is unbiased, the worst-case MSE is the same as the worst-case variance, which is also positively correlated to the confidence interval (CI).
 Thus, an unbiased estimator optimized for $\widehat{MSE}$ is also optimized for CI.
@@
 3.
-Collect all the responses $(z^{(i)})_{i\in[n]}$, decode them, and acquire the corresponding $\hat{y}^{(i)}$ as an estimator of $y^{(i)}$, i.e., $\hat{y}^{(i)}=R(y^{(i)})$, where $R$ is defined in Section 2.1.
+Collect all the responses $(z^{(i)})_{i\in[n]}$, decode them following Section 2.1 as $(\hat{y}^{(i)})_{i\in[n]}$, where each $\hat{y}^{(i)}$ is an estimator of $y^{(i)}$, i.e., $\hat{y}^{(i)}=R(y^{(i)})$, with $R$ defined in Section 2.1.
 Notice that each $\hat{y}^{(i)}$ is a vector of size $m$.
@@
 $\hat{f}(x)=\frac{m}{n(m-1)}\bigl{[}\sum_{i\in[n]}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{]}-\frac{1}{m-1}.$ (2)
-The only output that depends on the value of $X^{(i)}$ is $z^{(i)}$, which is perturbed by an $\epsilon$-LDP mechanism.
+Notably, $\hat{f}(x)$ may produce values outside the interval $[0,1]$ due to perturbations in the collected data.
+One straightforward solution is to truncate $\hat{f}(x)$ to $[0,1]$, i.e.
+$\min(1,\max(0,\hat{f}(x)))$, which also improves precision by eliminating deviations beyond the bounds.
+More advanced algorithms, such as projection onto a simplex, can also be used to enforce $\hat{f}(x)\in[0,1]$ and improve precision, as discussed in [6].
+Regarding privacy guarantee, the only output that depends on the value of $X^{(i)}$ is $z^{(i)}$, which has been perturbed by an $\epsilon$-LDP mechanism.
 The postprocessing of the outputs from an LDP mechanism does not change its privacy guarantee [16], so we have proved:
@@
 The expectation of $\hat{f}(x)$ is
-$E[\hat{f}(x)]\\ =\frac{m}{m-1}[f(x)+\sum_{x^{\prime}\in[d]\backslash x}\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}f(x^{\prime})]-\frac{1}{m-1}$ (3)
+$E[\hat{f}(x)]=\frac{m}{m-1}[f(x)\\ +\sum_{x^{\prime}\in[d]\backslash x}\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}f(x^{\prime})]-\frac{1}{m-1}$ (3)
 where $f(x)$ is the frequency of value $x$; $[d]\backslash x$ represents the set of $[d]$ excluding $x$.
 Corollary 3.1.
-If and only if $\forall x,x^{\prime}:\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}=\frac{1}{m}$,
+If and only if $\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}=\frac{1}{m}$ is satisfied for all $x\neq x^{\prime}$, then
 $E[\hat{f}(x)]=f(x).$
-The proof of Theorem 3.2 and Corollary 3.1 is presented in Appendix A.
-It is notable that $\forall x,x^{\prime}:\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}=\frac{1}{m}$ is equivalent to $\Pr_{h\in\mathcal{H}}(h(x)=h(x^{\prime}))=\frac{1}{m}$, where $h\in\mathcal{H}$ represents uniform sampling, which is also the definition of a pairwise-independent hashing family.
+The proof of Theorem 3.2 and Corollary 3.1 is presented in Appendix D.
+It is notable that $\forall x\neq x^{\prime}:\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}=\frac{1}{m}$ is equivalent to $\Pr_{h\in\mathcal{H}}(h(x)=h(x^{\prime}))=\frac{1}{m}$, where $h\in\mathcal{H}$ represents uniform sampling, which is also the definition of a pairwise-independent hashing family.
 Thus, Corollary 3.1 can be rephrased as:
@@
 If and only if $\mathcal{H}$ is pairwise independent, i.e., $\Pr_{h\in\mathcal{H}}(h(x)=h(x^{\prime}))=\frac{1}{m}$, we have $E[\hat{f}(x)]=f(x)$.
-The pairwise independence requires the size of $\mathcal{H}$ to be $d^{2}$, so the communication cost of sending one hash function becomes $2\log_{2}(d)$.
-The creation of such a hashing family is presented in Section 3.5.
 3.2 Variance
@@
 If $R$ satisfies Properties 2.1 and 2.2, we have
-$Var(\hat{f}(x))=\frac{m^{2}}{(m-1)^{2}n}\sum_{x^{\prime}\in[d]}\biggl{[}\biggl{(}\frac{1}{k}\sum_{j\in[k]}\\ Var\bigl{(}R(h_{j}(x^{\prime}))[h_{j}(x)]\bigr{)}\biggr{)}+\bar{c}(x^{\prime},x)-\bar{c}(x^{\prime},x)^{2}\biggr{]}f(x^{\prime}),$ (4)
+$Var(\hat{f}(x))=\frac{m^{2}}{(m-1)^{2}n}\sum_{x^{\prime}\in[d]}\biggl{[}\biggl{(}\frac{1}{k}\sum_{j\in[k]}Var\bigl{(}\\ R(h_{j}(x^{\prime}))[h_{j}(x)]\bigr{)}\biggr{)}+\bar{c}(x^{\prime},x)-\bar{c}(x^{\prime},x)^{2}\biggr{]}f(x^{\prime}),$ (4)
 where $\bar{c}(x^{\prime},x)=\sum_{j\in[k]}\frac{c_{j}(x^{\prime},x)}{k}$.
@@
 If $R$ satisfies Properties 2.1 - 2.3, we have
-$Var(\hat{f}(x))=\frac{m^{2}}{(m-1)^{2}n}\biggl{[}\\ \sum_{x^{\prime}\in[d]\backslash x}f(x^{\prime})\biggl{(}\bar{c}(x^{\prime},x)Var(R|=)+(1-\bar{c}(x^{\prime},x))\\ Var(R|\neq)+\bar{c}(x^{\prime},x)(1-\bar{c}(x^{\prime},x))\biggr{)}+f(x)Var(R|=)\biggr{]},$ (5)
+$Var(\hat{f}(x))=\frac{m^{2}}{(m-1)^{2}n}\biggl{[}\sum_{x^{\prime}\in[d]\backslash x}f(x^{\prime})\\ \biggl{(}\bar{c}(x^{\prime},x)Var(R|=)+(1-\bar{c}(x^{\prime},x))Var(R|\neq)+\\ \bar{c}(x^{\prime},x)(1-\bar{c}(x^{\prime},x))\biggr{)}+f(x)Var(R|=)\biggr{]},$ (5)
 where $Var(R|=)$ and $Var(R|\neq)$ are defined in Property 2.3.
@@
 $Var(\hat{f}(x))\\ =\frac{m}{(m-1)^{2}n}\biggl{[}\bigl{(}1-f(x)\big{)}\biggl{(}Var(R|=)+(m-1)\\ Var(R|\neq)+\frac{m-1}{m}\biggr{)}+mf(x)Var(R|=)\biggr{]}.$ (6)
-The lemma, theorem, and corollary of this section are proven in Appendix B.
-3.3 Accuracy Analysis
-The mean squared error (MSE) measures the accuracy of $\hat{f}(x)$, which is defined as
+The lemma, theorem, and corollary of this section are proven in Appendix E.
+3.3 Accuracy and Precision Analysis
+The mean squared error (MSE) measures the accuracy and precision of $\hat{f}(x)$, which is defined as
 $MSE(\hat{f}(x))$ $=E[(\hat{f}(x)-f(x))^{2}]$
@@
 and
-$MSE(g(x))=\frac{1}{n(1-\bar{c}(x,x^{\prime}))^{2}}\biggl{[}\\ (1-f(x))\biggl{(}\bar{c}(x^{\prime},x)Var(R|=)+(1-\bar{c}(x^{\prime},x))\\ Var(R|\neq)+\bar{c}(x^{\prime},x)(1-\bar{c}(x^{\prime},x))\biggr{)}+f(x)Var(R|=)\biggr{]}.$ (10)
+$MSE(g(x))=\frac{1}{n(1-\bar{c}(x,x^{\prime}))^{2}}\biggl{[}(1-f(x))\\ \biggl{(}\bar{c}(x^{\prime},x)Var(R|=)+(1-\bar{c}(x^{\prime},x))Var(R|\neq)\\ +\bar{c}(x^{\prime},x)(1-\bar{c}(x^{\prime},x))\biggr{)}+f(x)Var(R|=)\biggr{]}.$ (10)
 The formal proof is similar to that of the original estimator $\hat{f}(x)$, so the proof is omitted here.
 The rationale is that the original estimator $\hat{f}(x)$ assumes $\bar{c}(x,x^{\prime})$ being $\frac{1}{m}$, but here we integrate $\bar{c}(x,x^{\prime})$ into the estimation and replace $\frac{1}{m}$ with $\bar{c}(x,x^{\prime})$.
-Note that Theorem 10 is not applicable to the original CMS [7] because its $\bar{c}(x,x^{\prime})$ is not constant. We will analyze its accuracy in Section 3.6.
-Besides MSE, one may also use the confidence interval (CI) of $\hat{f}(x)$ to measure accuracy.
+Besides MSE, one may also use the confidence interval (CI) of $\hat{f}(x)$ to measure precision.
 If $\hat{f}(x)$ is unbiased, its CI is the same as its concentration bound.
@@
 $Pr\bigl{[}|\hat{f}(x)-f(x)|\geq\alpha\sqrt{Var(\hat{f}(x))}\bigr{]}\\ \leq 2\exp(-\frac{\alpha^{2}}{3}\cdot\frac{m-1}{e^{\epsilon}+m-1}).$
-This theorem is proven in Appendix C. Note that CMS+RR will be proven to be optimized in Section 3.4, so this theorem is applicable directly to the optimized CMS.
-3.4 Optimized CMS
-This section explores the optimized parameters of CMS to reduce the $l_{1}$ loss, $l_{2}$ loss, and the worst-case MSE defined in Section 2.2.
-It is notable that $l_{2}$ loss is connected to MSE as follows:
+This theorem is proven in Appendix F. Notably, this theorem focuses solely on CMS+RR, as Section 3.4 will demonstrate that it outperforms CMS implementations with other known perturbations.
+3.4 Optimizing CMS
+This section explores the parameters of CMS to reduce the $l_{1}$ loss, $l_{2}$ loss, and the worst-case MSE defined in Section 2.2. It is notable that $l_{2}$ loss is connected to MSE as follows:
 $l_{2}(\hat{f})=\max_{\forall\mathbf{X}:|\mathbf{X}|=n}\sum_{x\in[d]}E[(\hat{f}(x)-f(x))^{2}]\\ =\max_{\forall\mathbf{X}:|\mathbf{X}|=n}\sum_{x\in[d]}MSE(\hat{f}(x)).$ (11)
@@
 Assuming that the privacy factor $\epsilon$ and the dictionary $[d]$ are given, the undetermined variables in this equation are only the LDP process $R$ and the range of a hash function $m$.
-We start with the LDP process $R$.
-[9] summarized the state-of-the-art frequency estimation algorithms at that time.
-When considering different privacy regimes and original frequencies, i.e., $f(x)$, we found that randomized response (RR), symmetric RAPPOR (s-RP), and asymmetric RAPPOR (a-RP) are optimized in some privacy regimes.
-Though more optimized algorithms exist, their performance is either as optimized as these algorithms in their respective regimes, or they can be converted to CMS+RR (as shown in Section 4).
-Thus, we only consider RR, s-RP, and a-RP as candidates for the LDP process $R$.
-The variance of the these mechanisms are listed below:
-$Var(RR(a)[b])=\\ \frac{1}{(e^{\epsilon}-1)^{2}}\begin{cases}e^{\epsilon}(m-1)&\text{ if }a=b\\ e^{\epsilon}+m-2&\text{ if }a\neq b,\end{cases}$ (12)
-$Var(aRP(a)[b])=\\ \frac{1}{(e^{\epsilon}-1)^{2}}\begin{cases}(e^{\epsilon}+1)^{2}&\text{ if }a=b\\ 4e^{\epsilon}&\text{ if }a\neq b,\end{cases}$ (13)
-and
-$\forall a,b:Var(sRP(a)[b])=\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}.$ (14)
-Then, we explore the optimized $m$ for different LDP processes and compare their loss functions and the worst-case MSE.
-We start with the worst-case MSE.
-Here, we allow an analyst to have prior knowledge of the upper bound of $f(x)$, and we derive:
+We first consider randomized response (RR) as the LDP process, whose variance is as follows:
+$Var(RR(v)[v^{\prime}])=\\ \frac{1}{(e^{\epsilon}-1)^{2}}\begin{cases}e^{\epsilon}(m-1)&\text{ if }v=v^{\prime}\\ e^{\epsilon}+m-2&\text{ if }v\neq v^{\prime}.\end{cases}$ (12)
+Now, we optimize $m$ for CMS+RR, starting with the worst-case MSE. Here, we assume that an analyst has prior knowledge of the upper bound of $f(x)$ and derive:
 Theorem 3.6.
-If the upper bound of $f(x)$ is known as $\forall x\in\mathbf{x}:f(x)\leq f^{*}$, where $\mathbf{x}$ are the values of interest, the worst-case MSE among $\mathbf{x}$ is optimized as
+If $\forall x\in\mathbf{x}:f(x)\leq f^{*}$ is guaranteed to be correct by prior knowledge, where $\mathbf{x}$ are the values of interest, the worst-case MSE among $\mathbf{x}$ for CMS+RR is optimized as
 $\widehat{MSE}(\hat{f})=\begin{cases}\frac{2(\Delta_{MSE}+e^{\epsilon})}{n(e^{\epsilon}-1)^{2}}&\text{ if }f^{*}\leq\frac{1}{2}\\ \frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}&\text{ if }\frac{1}{2}<f^{*}\leq 1\end{cases},$
@@
 $\Delta_{MSE}=e^{\epsilon/2}\sqrt{[(1-f^{*})e^{\epsilon}+f^{*}][f^{*}e^{\epsilon}+(1-f^{*})]},$
-which can be achieved when the LDP process $R$ is randomized response, and $m$ is the closest integer to
+and the hash-function range $m$ is the closest integer to
 $\begin{cases}1+\frac{\Delta_{MSE}}{f^{*}e^{\epsilon}+1-f^{*}}&\text{ if }f^{*}\leq\frac{1}{2}\\ 1+e^{\epsilon/2}&\text{ if }\frac{1}{2}<f^{*}\leq 1.\end{cases}$ (13)
-The theorem is proven in Appendix D. Notably, $f^{*}$ is prior knowledge of an analyst and could be larger than $\max_{x}f(x)$, which is unknown to the analyst. Subsequently, we have:
+The theorem is proven in Appendix B.
+Notably, $f^{*}$ is prior knowledge of an analyst and could be larger than $\max_{x}f(x)$, which is unknown to the analyst.
+When $f^{*}\geq\frac{1}{2}$, it has the same worst-case MSE as $f^{*}=1$, which is guaranteed to be correct.
+On the other hand, if $f^{*}<\frac{1}{2}$ and the prior knowledge happens to be incorrect, i.e., $\max_{x}f(x)>f^{*}$, we demonstrate that
+Theorem 3.7.
+Given the optimized CMS+RR in Theorem 13, if $\delta=\max_{x}f(x)-f^{*}>0$ and $f^{*}<\frac{1}{2}$,
+$\widehat{MSE}(\hat{f})=\widehat{MSE}(\hat{f}|f^{*})+\frac{\delta(1-2f^{*})e^{\epsilon}}{n\Delta_{MSE}}\\ \leq\widehat{MSE}(\hat{f}|f^{*})+\frac{\delta(1-2f^{*})}{n}.$ (14)
+where $\widehat{MSE}(\hat{f})$ and $\widehat{MSE}(\hat{f}|f^{*})$ are the actual worst-case MSE and the worst-case MSE assuming $\max_{x}f(x)\leq f^{*}$, respectively.
+The theorem is proven in Appendix B.
+Since $\widehat{MSE}(\hat{f}|f^{*})\geq\widehat{MSE}(\hat{f}|f^{*}=0)=\frac{4e^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$, as long as $\delta(1-2f^{*})=O(\frac{4e^{\epsilon}}{n(e^{\epsilon}-1)^{2}})$, we still have $\widehat{MSE}(\hat{f})=O(\widehat{MSE}(\hat{f}|f^{*}))$.
+Therefore, if the prior knowledge $f^{*}$ is guaranteed to be correct or its error is bounded by the above condition, considering the prior knowledge $f^{*}$ in the $\widehat{MSE}$ calculation is recommended.
+Section 6.2 provides an example of utilizing $f^{*}$.
+Otherwise, one can always assume $f^{*}=1$ for simplicity.
+Hereinafter, we assume no prior knowledge of $f(x)$ unless explicitly mentioned, i.e., $f^{*}=1$ by default.
+Next, we optimize the $l_{1}$ and $l_{2}$ losses:
+Theorem 3.8.
+The $l_{2}$ loss and the upper bound of $l_{1}$ loss of CMS+RR are optimized as
+$l_{2}^{*}(\hat{f})=\frac{2(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}},$
+and
+$l_{1}^{*}(\hat{f})\leq\sqrt{\frac{2d(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}}},$
+respectively, where
+$\Delta_{l}=e^{\epsilon/2}\sqrt{(e^{\epsilon}+d-1)(de^{\epsilon}-e^{\epsilon}+1)},$
+and $m$ is the closest integer to $1+\frac{\Delta_{l}}{e^{\epsilon}+d-1}$.
 Corollary 3.4.
-If the upper bound of $f(x)$ is unknown, i.e., $f^{*}=1$, the worst-case MSE is optimized as
-$\widehat{MSE}(\hat{f})=\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}},$
-which can be achieved when $R$ is randomized response, and $m$ is the closest integer to $1+e^{\epsilon/2}$
-Hereinafter, we assume no prior knowledge of $f(x)$ unless explicitly mentioned, i.e., $f^{*}=1$ by default.
-On the other hand, if an analyst is interested in some rare values, and the corresponding $f^{*}$ is around zero, we can substitute $f^{*}=0$ into Theorem 15 and obtain $m=1+e^{\epsilon}$, which should be sufficient to approximate the optimized $m$.
-Its worst-case variance is derived as:
-$\widehat{MSE}(\hat{f})=\frac{1}{n}[f^{*}+\frac{4e^{\epsilon}}{(e^{\epsilon}-1)^{2}}].$ (16)
-We can see $m=1+e^{\epsilon}$ works well if $f^{*}$ is not a dominant term in Eq. (16).
-Next, we derive the optimized $l_{1}$ and $l_{2}$ losses:
-Theorem 3.7.
-The $l_{1}$ and $l_{2}$ losses are optimized as
-$l_{1}^{*}(\hat{f})\leq\sqrt{\frac{2d(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}}},$
+When $d\gg e^{\epsilon}$, the optimized $l_{2}$ loss and the upper bound of the optimized $l_{1}$ loss of CMS+RR approach
+$l_{2}^{*}(\hat{f})\rightarrow\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}},$
 and
-$l_{2}^{*}(\hat{f})=\frac{2(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}},$
-respectively, where
-$\Delta_{l}=e^{\epsilon/2}\sqrt{(e^{\epsilon}+d-1)(de^{\epsilon}-e^{\epsilon}+1)}.$
-Both optimalities are achieved when the LDP process $R$ is randomized response, and $m$ is the closest integer to $1+\frac{\Delta_{l}}{e^{\epsilon}+d-1}$.
-Corollary 3.5.
-When the dictionary size $d$ is large enough, the optimized $l_{1}$ and $l_{2}$ losses approach
 $\lceil l_{1}\rceil^{*}(\hat{f})\rightarrow\frac{2de^{\epsilon/2}}{\sqrt{n}(e^{\epsilon}-1)},$
-and
-$l_{2}^{*}(\hat{f})\rightarrow\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}},$
 respectively, where $\lceil l_{1}\rceil$ represents the upper bound of the $l_{1}$ loss, and the corresponding $m\rightarrow 1+e^{\epsilon}$.
-The theorem and corollary above are proven in Appendix D.
-It should be noted that the CMS using symmetric or asymmetric RAPPOR as its LDP process cannot serve as the optimized CMS for the following reasons: (1) their performance in reducing worst-case MSE or $l_{1}$ / $l_{2}$ losses is outperformed by CMS+RR in some conditions, and (2) they have very high communication costs.
-Please refer to Appendix D for more details.
-One caveat of the setting to minimize $l_{1}$ / $l_{2}$ losses is that it will lead to a worst-case MSE larger than the optimized value. Assuming $d$ is large enough, and $m=1+e^{\epsilon}$, we have
-$\widehat{MSE}(\hat{f})=Var(\hat{f}(x)_{RR}|f(x)=1)=\frac{(e^{\epsilon}+1)^{2}}{n(e^{\epsilon}-1)^{2}},$ (17)
-which is always larger than the optimized $\widehat{MSE}$, i.e., $\frac{e^{\epsilon/2}}{n(e^{\epsilon}-1)^{2}}$.
+The theorem and corollary above are proven in Appendix B.
+One caveat is that the CMS+RR optimized for $l_{1}$ / $l_{2}$ losses is not optimized for worst-case MSE.
+Assuming $d$ is large enough, and $m=1+e^{\epsilon}$, we have
+$\widehat{MSE}(\hat{f})=Var(\hat{f}(x)_{RR}|f(x)=1)=\frac{(e^{\epsilon}+1)^{2}}{n(e^{\epsilon}-1)^{2}},$
+which is always larger than the optimized $\widehat{MSE}$ of CMS+RR, i.e., $\frac{e^{\epsilon/2}}{n(e^{\epsilon}-1)^{2}}$.
 What is worse, even when $\epsilon$ is very large, it only converges to $\frac{1}{n}$, not zero.
-Therefore, before an analyst estimates the frequencies of a dataset, they need to determine whether to focus on (1) the overall accuracy (i.e., $l_{1}$ / $l_{2}$ losses) or (2) the worst-case accuracy of the individual estimations (i.e., $\widehat{MSE}$).
-Since optimizing $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses requires different $m$, we will hereinafter refer to the CMS optimized for $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses as MSE-OCMS and $l$-OCMS, respectively.
-Additionally, the optimized CMS will be compared with the existing algorithms in Section 5.1.
+Therefore, before an analyst estimates the frequencies of a dataset, they need to determine whether to focus on (1) the overall precision (i.e., $l_{1}$ / $l_{2}$ losses) or (2) the worst-case precision of the individual estimations (i.e., $\widehat{MSE}$).
+Since optimizing $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses requires different $m$, we will hereinafter refer to the CMS+RR optimized for $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses as MSE-OCMS+RR and $l$-OCMS+RR, respectively.
+Additionally, OCMS+RR will be compared with the existing algorithms in Section 5.1.
+Now, we evaluate the CMS using other existing LDP processes.
+A study [9] summarized the latest frequency estimation algorithms at that time.
+When considering different privacy regimes, they found that randomized response (RR), symmetric RAPPOR (RP), and asymmetric RAPPOR (a-RP) achieved state-of-the-art precision in some privacy regimes.
+Though more advanced algorithms exist, they are either as precise as these algorithms in their respective regimes, or they can be converted to CMS+RR (as shown in Section 4).
+Thus, in addition to RR, we will also evaluate CMS using RP and a-RP as the LDP process.
+Since the precision of these two mechanisms is independent of their input cardinality $m$, and $Var(\hat{f}(x))$ decreases with $m$ when $R$ is independent of $m$, the loss functions of CMS with RAPPOR mechanisms are optimized when $m\rightarrow\infty$.
+Appendix G derives their values as follows:
+CMS+ RR $sRP$ $aRP$
+$\widehat{MSE}(\hat{f}|$ $f^{*}<\frac{1}{2})$ $\frac{2(\Delta_{MSE}+e^{\epsilon})}{n(e^{\epsilon}-1)^{2}}$ $\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}$ $\frac{f^{*}(e^{\epsilon}-1)^{2}+4^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$
+$\widehat{MSE}(\hat{f}|$ $f^{*}\geq\frac{1}{2})$ $\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}$ $\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}$ $\frac{f^{*}(e^{\epsilon}-1)^{2}+4^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$
+$l_{2}$ $\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$ $\frac{de^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}$ $\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$
+CMS using RAPPOR mechanisms are less preferred for the following reasons: (1) their performance in reducing worst-case MSE or $l_{1}$ / $l_{2}$ losses is outperformed by CMS+RR in some conditions (highlighted in red in the table), and (2) they have very high communication costs (see Appendix G for details).
 3.5 Unbiased CMS with Imperfect Hashing
-An optimized CMS is built upon an unbiased CMS, which requires its hashing to be pairwise independent.
-However, the construction of a $K$-wise-independent hashing family requires the range of the hashed values (i.e., $[m]$) to comprise a finite field [17].
+OCMS+RR is built upon an unbiased CMS, which requires its hashing to be pairwise independent.
+However, the construction of a $K$-wise-independent hashing family requires the range of the hashed values (i.e., $[m]$) to comprise a finite field (Construction 3.32 of [17]).
 A finite field must have $p^{l}$ elements [18], where $p$ is a prime number and $l\in\mathbb{N}^{+}$.
@@
 In practice, we can still construct an approximately pairwise-independent hashing family for any $m$.
-Consider a perfect pairwise independent hashing family with $p=2$ and $2^{l}\geq d$:
-=$\mathcal{H}=\{h(x)|\forall a_{i}\in[2^{l}]:h(x)=a_{0}+a_{1}x\}$ (18)
-where $a_{0}$, $a_{1}$, and $x$ are elements in a finite field $\mathbb{F}(2^{l})$, so the addition and multiplication in Eq.
+Consider a perfect pairwise independent hashing family with $p^{l}>d$:
+$\mathcal{H}_{pi}=\{h(x)=a_{0}+a_{1}x\mid a_{0},a_{1}\in[p^{l}]\}.$ (15)
+where $a_{0}$, $a_{1}$, and $x$ are elements in a finite field $\mathbb{F}(p^{l})$, so the addition and multiplication in Eq.
 (15) are all field operations.
-The range of any $h(x)$ is also the field $\mathbb{F}(2^{l})$.
-Given two random variables $a_{0}$ and $a_{1}$, the size of $\mathcal{H}$ is $2^{2l}=\Theta(d^{2})$.
-Subsequently, we construct a new hashing family based on $\mathcal{H}$:
-=$\mathcal{H}^{\prime}=\{h^{\prime}(x)|h(x)\bmod m\}.$
-Since $\forall h\in\mathcal{H}$, $\forall y\in[m]$: $Pr(h(x)=y)=2^{-l}$, and assuming $2^{l}=qm+r$, where $q\in\mathbb{N}$ and $r\in\mathbb{N}$, we will have $r$ hashed values with $Pr(h^{\prime}(x)=y)=\frac{q+1}{2^{l}}$ and the remaining $m-r$ hashed values with $Pr(h^{\prime}(x)=y)=\frac{q}{2^{l}}$.
+The range of any $h(x)$ is also the field $\mathbb{F}(p^{l})$.
+One approach to construct $p^{l}$ is to find a prime number $p^{\prime}>d$ and set $l=1$, i.e., $p^{l}=p^{\prime}$.
+In this case, the field operations of $\mathbb{F}(p^{\prime})$ simply require every operation to modulo $p^{\prime}$ [19].
+However, finding a prime number when $d$ is very large can be time-consuming.
+On the other hand, we can find any $l$ such that $2^{l}>d$, and set $p=2$, and the field operations of $\mathbb{F}(2^{l})$ are introduced in [19].
+Subsequently, we construct a new hashing family based on $\mathcal{H}_{pi}$:
+$\mathcal{H}_{api}=\{h^{\prime}(x)=h(x)\bmod m\mid h(x)\in\mathcal{H}_{pi}\}.$
+where $api$ represents approximately pairwise independent.
+Notably, the finite field size $p^{l}$ should also be no less than $m$.
+Since $\forall h\in\mathcal{H}_{pi}$, $\forall y\in[m]$: $Pr(h(x)=y)=p^{-l}$, and assuming $p^{l}=qm+r$, where $q\in\mathbb{N}$ and $r\in\mathbb{N}$, we will have $r$ hashed values with $Pr(h^{\prime}(x)=y)=\frac{q+1}{p^{l}}$ and the remaining $m-r$ hashed values with $Pr(h^{\prime}(x)=y)=\frac{q}{p^{l}}$.
 Thus, we have:
@@
 $\forall x_{1},x_{2}\in[d]^{2}$:
-$Pr_{h^{\prime}\in\mathcal{H}^{\prime}}(h^{\prime}(x_{1})=h^{\prime}(x_{2}))=r(\frac{q+1}{2^{l}})^{2}+(m-r)(\frac{q}{2^{l}})^{2}\\ =\frac{(2q+1)r+mq^{2}}{(mq+r)^{2}},$ (19)
-where $h^{\prime}\in\mathcal{H}^{\prime}$ represents uniform sampling.
-Notice that $Pr_{h^{\prime}\in\mathcal{H}^{\prime}}(h^{\prime}(x_{1})=h^{\prime}(x_{2}))=\bar{c}(x_{1},x_{2})$, which is a constant and can be substituted into the $g(x)$ in Theorem 10, where $m^{\prime}=\frac{(mq+r)^{2}}{(2q+1)r+mq^{2}}$.
+$Pr_{h^{\prime}\in\mathcal{H}_{api}}(h^{\prime}(x_{1})=h^{\prime}(x_{2}))\\ =r(\frac{q+1}{p^{l}})^{2}+(m-r)(\frac{q}{p^{l}})^{2}=\frac{(2q+1)r+mq^{2}}{(mq+r)^{2}},$ (16)
+where $h^{\prime}\in\mathcal{H}_{api}$ represents uniform sampling.
+Notice that $Pr_{h^{\prime}\in\mathcal{H}_{api}}(h^{\prime}(x_{1})=h^{\prime}(x_{2}))=\bar{c}(x_{1},x_{2})$, which is a constant and can be substituted into the $g(x)$ in Theorem 10, where $m^{\prime}=\frac{(mq+r)^{2}}{(2q+1)r+mq^{2}}$.
 As a result, $g(x)$ is an unbiased estimator of $f(x)$.
@@
 we have
-$Var(g(x|\mathcal{H}^{\prime}))<(1+\tau)Var(\hat{f}(x)),$
-where $g(x|\mathcal{H}^{\prime})$ and $\hat{f}(x)$ represent the $g(x)$ with $\mathcal{H}^{\prime}$ and a standard estimator of $f(x)$ with a pairwise-independent hashing family.
-Consider the optimized CMS for the worst-case MSE (Corollary 3.4), we can derive:
+$Var(g(x|\mathcal{H}_{api}))<(1+\tau)Var(\hat{f}(x)),$
+where $g(x|\mathcal{H}_{api})$ and $\hat{f}(x)$ represent the $g(x)$ with $\mathcal{H}_{api}$ and a standard estimator of $f(x)$ with a pairwise-independent hashing family.
+Applying the theorem to OCMS+RR, we can derive:
 Corollary 3.5.
-When $2q+1>\sqrt{\frac{1}{\tau}}$ and $m=1+e^{\epsilon/2}$, we have
+When $2q+1>\sqrt{\frac{1}{\tau}}$, we have
 $MSE(g)$ $<(1+\tau)\widehat{MSE}(\hat{f})$
-Corollary 3.7.
-When $2q+1>\sqrt{\frac{1}{\tau}}$ and $m=1+e^{\epsilon}$, we have
 $l_{2}(g)$ $<(1+\tau)l_{2}^{*}(\hat{f})$
-The condition of these two corollaries is independent of $\epsilon$. If $\tau=0.01$, we have $q>4.5$ to satisfy the two corollaries above.
-On the other hand, one may prefer using the original frequency estimator $\hat{f}(x)$ in Eq.
-(2) instead of $g(x)$, while still employing the imperfect hashing.
-The following theorem will demonstrate how large $q$ is required to make the deviation from $MSE(\hat{f}(x))$ negligible.
-Theorem 3.9.
-When
-$\frac{1}{4(m-1)q(q+1)}\leq\sqrt{\tau Var(\hat{f}(x))}$
-and
-$\frac{8q^{2}+8q+1}{16q^{2}(q+1)^{2}}\leq\tau,$
-we have
-$MSE(f(x)|\mathcal{H}^{\prime})<(1+2\tau)MSE(\hat{f}(x)),$
-where $MSE(\hat{f}(x))$ represents the MSE of the $f(x)$ with the perfect hashing.
-Notably, $Var(\hat{f}(x))=\Theta(1/n)$, so $q=\Theta(n^{1/4})$. That is, the size of $\mathcal{H}^{\prime}$ has to increase with the size of the dataset.
-All the theorems and corollaries in this section are proven in Appendix E.
-3.6 Correcting the original CMS
-In the original CMS paper [7], CMS randomly generates $k$ hash functions, and each hash function $h_{j}(x)$ is generated by sampling $d$ mutually independent random variables, each uniformly distributed among $[m]$.
-Subsequently, $\forall(x,x^{\prime})\in[d]^{2},c_{j}(x,x^{\prime})$ can be considered as a Bernoulli random variable with a probability of $\frac{1}{m}$ of being one.
+for MSE-OCMS+RR and $l$-OCMS+RR, respectively.
+The theorem and corollary are proven in Appendix H.
+Notably, the condition of this corollary is independent of $\epsilon$.
+If $\tau=0.01$, we have $q>4.5$ to satisfy the two corollaries above.
+Therefore, one could use $g(x)$ to approximate $\hat{f}(x)$ if $p^{l}\geq 5m$.
+Since $a_{0}$ and $a_{1}$ of $\mathcal{H}_{api}$ can be any numbers in $\mathbb{F}(p^{l})$, we have $|\mathcal{H}_{api}|=(p^{l})^{2}=p^{2\lceil\log_{p}\max\{d+1,5m\}\rceil}$, and $2\lceil\log_{2}\max\{d+1,5m\}\rceil$ bits are sufficient to represent a hash function in $\mathcal{H}_{api}$ if $p=2$.
+3.6 Implementation
+Algorithm 1 OCMS+RR (Part 1)
+1:$d$: dictionary size
+2:$\epsilon$: privacy factor
+3:Mode := either MSE-OCMS+RR or $l$-OCMS+RR
+4:$f^{*}$: prior knowledge on the upper bound of $f(x)$. Default value is one.
+[…]
+6:
+7:function FiniteFieldSize($d$, $m$)
+8:     $v\coloneq\max\{d+1,5m\}$ // see Corollary 3.5 for why $5m$
+[…]
+[…]
+11:     end if
+12:     return $2^{\lceil\log_{2}(v)\rceil}$
+13:end function
+14:
+15:function HashRange($\epsilon,d,f^{*}$, Mode)
+16:     if Mode is MSE-OCMS+RR then
+17:         if $f^{*}\geq\frac{1}{2}$ then
+18:              return round( $e^{\epsilon/2}+1$ )
+19:         end if
+20:         $\Delta\coloneq e^{\epsilon/2}\sqrt{[(1-f^{*})e^{\epsilon}+f^{*}][f^{*}e^{\epsilon}+(1-f^{*})]}$
+21:         return round( $\frac{\Delta}{f^{*}e^{\epsilon}+1-f^{*}}+1$ )
+22:     else if Mode is $l$-OCMS+RR then
+23:         $\Delta\coloneq e^{\epsilon/2}\sqrt{(e^{\epsilon}+d-1)(de^{\epsilon}-e^{\epsilon}+1)}$
+24:         return round( $\frac{\Delta}{e^{\epsilon}+d-1}+1$ )
+25:     end if
+26:end function
+27:
+28:$m\coloneq$ :- HashRange($\epsilon,d,f^{*}$, Mode)
+29:$p_{\mathcal{H}}\coloneq$ :- FiniteFieldSize($d,m$) // i.e., $p^{l}$ in Section 3.5.
+30:
+31:function Hash($a_{0},a_{1},x,m,p_{\mathcal{H}}$)
+[…]
+33:         $v\coloneq(a_{0}+a_{1}x)\mod{p_{\mathcal{H}}}$
+34:     else
+[…]
+[…]
+37:     end if
+38:     return $v\mod{m}$
+39:end function
+40:
+41:// return $Q[v,u]$ in Section 2.1.
+42:function Decode($u$: perturbed value, $v$: value of interest, $m$, $\epsilon$)
+43:     if $u=v$ then
+44:         return $\frac{e^{\epsilon}+m-1}{e^{\epsilon}-1}$
+45:     end if
+46:     return $\frac{-1}{e^{\epsilon}-1}$
+47:end function
+48:
+49:function Client($x$: true value of a client, $X^{(i)}$)
+50:     Uniformly Sample $a_{0}$ and $a_{1}$ from $[0,p_{\mathcal{H}}-1]$.
+51:     $y\coloneq$ :- Hash($a_{0},a_{1},x,m,p_{\mathcal{H}}$) // i.e., $h_{j^{(i)}}(X^{(i)})$
+52:     return RandomizedResponse($y$, $m$, $\epsilon$), $a_{0},a_{1}$
+53:end function
+Algorithm 2 OCMS+RR (Part 2)
+54:function Server($x$: value of interest, $\mathbf{C}$: all responses from the clients)
+[…]
+56:     for $z,a_{0},a_{1}$ in $\mathbf{C}$ do
+57:         $h_{x}\coloneq$ :- Hash($a_{0},a_{1},x,m,p_{\mathcal{H}}$) // i.e., $h_{j^{(i)}}(x)$
+58:         $\hat{y}\coloneq$ :- Decode($z$, $h_{x}$, $m$, $\epsilon$) // i.e., $\hat{y}^{(i)}[h_{j^{(i)}}(x)]$
+59:         $\hat{y}_{tot}$ += $\hat{y}$
+60:     end for
+61:     $n\coloneq$ :- Length($\mathbf{C}$)
+62:     $q,\;r\coloneq\lfloor\>p_{\mathcal{H}}/m\>\rfloor,\;\;p_{\mathcal{H}}\mod{m}$
+63:     $m^{\prime}\coloneq\frac{(mq+r)^{2}}{(2q+1)r+mq^{2}}$
+64:     return $\frac{m^{\prime}}{n(m^{\prime}-1)}\hat{y}_{tot}-\frac{1}{m^{\prime}-1}$
+65:end function
+OCMS+RR (Algorithm 1) is implemented to not only adapt the optimized hash range $m$ proposed in Section 3.4, but also the approximately pairwise independent hashing family $\mathcal{H}_{api}$ and the corresponding estimator $g(x)$ in Section 3.5.
+The Client function perturbs each object $X^{(i)}$, and their perturbed values, along with their hashing functions, are sent to the server.
+The server then calls Server to estimate the frequency of any $x$, i.e., $f(x)$.
+In particular, when the dictionary size $d$ is small, i.e., $d<2^{64}-59$, the algorithm constructs the hashing family with $p^{l}$ as a prime number.
+Otherwise, it computes the minimum $2^{l}>d$ as $p^{l}$.
+The source code of OCMS+RR is available at https://github.com/mingen-pan/CountMeanSketch).
+3.7 Issues in the original CMS
+Most frequency estimators ensure that the estimated frequency converges to the actual frequency as the dataset size increases, provided each object is queried once.
+However, this section demonstrates that the estimation from the original CMS exhibits a constant bias, even with an infinitely large dataset.
+The original CMS randomly generates $k$ hash functions, and each hash function $h_{j}(x)$ is generated by sampling $d$ mutually independent random variables, each uniformly distributed among $[m]$.
+Subsequently, $\forall x,x^{\prime}\in[d]$ with $x\neq x^{\prime}$, $c_{j}(x,x^{\prime})$ can be considered as a Bernoulli random variable with a probability of $\frac{1}{m}$ of being one.
 Since the $k$ hash functions are generated independently, $\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}$ is equivalent to a binomial random variable, whose expectation is $\frac{1}{m}$.
 However, the probability of the output of a binomial distribution exactly equaling its expectation is very low, so $\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}$ will likely deviate from $\frac{1}{m}$ to some extent (see Slud’s Inequality [20]).
-Thus, $E[\hat{f}(x)]$ will likely deviate from $f(x)$ to some extent too, i.e., $\hat{f}(x)$ has bias.
-Notably, this bias is independent of the size of the dataset, so it will still exist even when an infinite amount of data is given for a single run of CMS.
-This conclusion is also applicable to the non-private Count Sketch families, which are also proven to have bias [20] if they query a dataset only once.
+Given Corollary 3.1, $E[\hat{f}(x)]$ will likely deviate from $f(x)$ to some extent too, i.e., $\hat{f}(x)$ has bias.
 To evaluate the degree of bias, when considering $\forall c_{j}(x,x^{\prime})$ as random variables, we can derive:
@@
 $\mathop{Var}_{\forall c_{j}(x,x^{\prime})}(E[\hat{f}(x)])=\frac{1}{(m-1)k}\sum_{x^{\prime}\in[d]\backslash x}f(x^{\prime})^{2}.$ (18)
-The proof is presented in Appendix F.1.
+The proof is presented in Appendix I.
 The variance indicates that the bias is $O(\frac{1}{\sqrt{(m-1)k}})$.
@@
 We found that they treated $c_{j}(x,x^{\prime})$ as independent random variables for every pair of $X^{(i)}$ and $X^{(i^{\prime})}$, which is incorrect because all pairs with the same values $x$ and $x^{\prime}$ will always have the same $c_{j}(x,x^{\prime})$.
-Eq. (7) demonstrates how the bias of $\hat{f}(x)$ contributes to its MSE. The following theorem demonstrates how to prevent the bias from being a dominant term in $MSE(\hat{f}(x))$:
-Theorem 3.11.
-If
-$\frac{\delta\tau Var(R|\neq)}{n}>\frac{1}{(m-1)k}$
-is satisfied, where $\tau$ is a small number, like 0.1, we have
-$Pr\biggl{(}(E[\hat{f}(x)]-f(x))^{2}\leq\tau Var(\hat{f}(x))\biggr{)}\geq 1-\delta.$ (22)
-The proof is presented in Appendix F.2.
-This implies that the number of hash functions $k$ should be $\Omega(\frac{n}{m})$ to ensure that the bias of $f(x)$ is negligible.
-However, this constraint was not mentioned in the original paper.
 The original paper also has an issue in its variance calculation, where they required the hashing family $\mathcal{H}$ to be 3-wise-independent because they assumed any $X^{(i_{1})}$ and $X^{(i_{2})}$ were correlated, which was problematic.
 In contrast, the non-private Count Sketch [21] only requires pairwise independence, which is consistent with our work.
-Notably, the LDP perturbation only affects the accuracy after the hashing, so it should not alter the pairwise-independence requirement for the hashing family.
-4 Connection with Existing Frequency Estimation Algorithms
-This section demonstrates that various frequency estimation algorithms proposed in the previous literature are equivalent or closely connected to CMS.
+Notably, the LDP perturbation only affects the precision after the hashing, so it should not alter the pairwise-independence requirement for the hashing family.
+4 Connection with Existing Frequency Estimators
+This section demonstrates that various frequency estimators proposed in the previous literature are equivalent or closely connected to CMS.
 First, we define the Transformation-and-Sampling Framework (TnS Framework) as follows:
@@
 (TnS Framework)
-1. Transform the original value $x$ to a vector of elements $M[x]$, i.e., a row of a matrix $M$.
-2. Uniformly sample an element from $M[x]$, denoted as $M[x][j]$
-3. Perturb the sampled element $M[x][j]$ with some LDP mechanism.
+1. Transform the original value $x$ to a vector of elements $M[x]$, i.e., a row of a matrix $M$, where the $j$-th element belongs to the finite set $\mathcal{V}_{j}$.
+2. Uniformly sample an element from the vector $M[x]$, denoted as $M[x][j]$.
+3. Perturb the sampled element $M[x][j]$ within $\mathcal{V}_{j}$ using some LDP mechanism.
 The above process is equivalent to (1) uniformly sampling a column from $M$, denoted as $M[:,j]$, (2) mapping the original value $x$ to $M[x,j]$, and (3) perturbing it.
-The column here functions as a hash function, and all the columns of $M$ comprise a hashing family.
+The column here serves as a hash function, and all the columns of $M$ comprise a hashing family.
 Thus, we have proved:
@@
 It is evident that HE belongs to the TnS framework, so it is equivalent to CMS.
-Additionally, Appendix G proves that decoding process of HE is equivalent to Eq.
+Additionally, Appendix J proves that the decoding process of HE is equivalent to Eq.
 (2).
@@
 4.3 Recursive Hadamard Response
-Recursive Hadamard Response (RHR) [12] is an improvement of Hadamard Encoding, increasing accuracy by communicating more bits.
+Recursive Hadamard Response (RHR) [12] is an improvement of Hadamard Encoding, increasing precision by communicating more bits.
 It (1) splits the original value $x$ into $x//2^{b-1}$ (floor division) and $x\bmod 2^{b-1}$; (2) creates a Hadamard matrix with a size larger than $d//2^{b-1}$, (3) select a row from the Hadamard matrix with an index equal to $(x//2^{b-1})+1$, and (4) uniformly samples an element from the selected row as $s$, and (5) perturbs $s\times(x\bmod 2^{b-1})$ using randomized response.
@@
 In CMS, any pair of inputs have the same collision probability of $1/m$, which is $2^{-b}$ given $2^{b}$ hashed values.
-Thus, RHR is a skewed version of CMS.
+Thus, RHR is a skewed version of CMS+RR.
 If the dataset is randomly generated, i.e., every pair of objects may have the same mod with a probability of $2^{b-1}$, then the hash collision probability becomes $2^{-b}$.
 Note that [12] sets $b=\epsilon$ if the dictionary is large enough.
-This is equivalent to $m=e^{\epsilon}$ in CMS, which is close but not identical to the optimized $m$ used by OCMS, so the accuracy of RHR at most has the same order as OCMS.
+This is equivalent to $m=e^{\epsilon}$ in CMS, which is close but not identical to the optimized $m$ used by OCMS+RR, so the precision of RHR at most has the same order as OCMS+RR.
 However, we can always generate a dataset with all values having the same mod when divided by $2^{b-1}$.
@@
 Therefore, this algorithm is equivalent to CMS.
-However, [9] did not compute the optimized $\widehat{MSE}$ correctly as they only considered $f(x)=0$, but Section 3.4 demonstrates that both $f(x)=0$ and $f(x)=1$ could be the maximum points of the MSE.
+The original work [9] assumed $f(x)\approx 0$ when optimizing the MSE of their algorithm, but Section 3.4 demonstrates that both $f(x)=0$ and $f(x)=1$ could be the maximum points of the MSE.
 Moreover, they did not optimize the communication cost of sending a hash function to the server, which will be discussed in Section 5.1.
@@
 Suppose the first CMS hashes the original values to $[m_{1}]$, and its perturbation algorithm is another CMS hashing its input to $[m_{2}]$.
-Given the pairwise independence, the collision probability $\bar{c}(x,x^{\prime})$ of two distinct input values is $\frac{m_{1}+m_{2}-1}{m_{1}m_{2}}$.
+Given pairwise independence, the collision probability $\bar{c}(x,x^{\prime})$ of two distinct input values is $\frac{m_{1}+m_{2}-1}{m_{1}m_{2}}$.
 Therefore, the recursive CMS is equivalent to a regular CMS with $m=\frac{m_{1}m_{2}}{m_{1}+m_{2}-1}$.
@@
 4.6 RAPPOR
-The encoding of RAPPOR is equivalent to CMS if each object in RAPPOR belongs to its own cohort and there is only one hash function in each cohort.
-Assuming there is sufficient data and we can ignore the uncertainty from the LDP mechanism and hash-function assignment, the encoding of RAPPOR and CMS can be considered a linear transformation from $\mathcal{R}^{d}$ (frequencies of original values) to $\mathcal{R}^{km}$ (frequencies of the hashed values).
-For the unbiased CMS (Corollary 3.1), we have $km=d^{2}m>d$, so CMS can derive the original frequencies without bias.
+The encoding of RAPPOR [15] is equivalent to CMS if each object in RAPPOR belongs to its own cohort and there is only one hash function in each cohort.
+The hashing of RAPPOR and CMS can be considered a linear transformation from $\mathcal{R}^{d}$ (frequencies of original values) to $\mathcal{R}^{km}$ (frequencies of the hashed values).
+For an unbiased CMS (Corollary 3.1), we have $km=d^{2}m>d$, so CMS can derive the original frequencies without bias.
 However, RAPPOR utilizes gradient descent to find the original frequencies and does not require $km\geq d$.
-[15] even chooses small $km$ on purpose to reduce computation.
+The original RAPPOR[15] even chooses small $km$ on purpose to reduce computation.
 Unfortunately, this is problematic because multiple values in a high-dimension simplex will map to the same value in a low-dimension simplex, and the value RAPPOR finds may not be the original value.
 5 Discussion
-5.1 Accuracy, Communication Cost, and Optimality
-Tables I and II compare the accuracy and communication cost among various frequency estimation algorithms, respectively.
-We will demonstrate that the optimized CMS is also the optimal algorithm when considering both accuracy and communication cost.
+5.1 Precision and Communication Cost
+Tables I and II compare the precision and communication cost among various frequency estimators, respectively.
+We will demonstrate that OCMS+RR is the leading algorithm when considering both precision and communication cost.
 $l_{1}$ $l_{2}$ $\widehat{MSE}$
@@
 OLH $\frac{2de^{\epsilon/2}}{e^{\epsilon}-1}$ $\frac{4de^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ $(\frac{e^{\epsilon}+1}{e^{\epsilon}-1})^{2}$
-OCMS $\frac{2de^{\epsilon/2}}{e^{\epsilon}-1}$ $\frac{4de^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ $\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$
+OCMS+RR $\frac{2de^{\epsilon/2}}{e^{\epsilon}-1}$ $\frac{4de^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ $\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$
 CMS+HE $\frac{d(e^{\epsilon}+1)}{e^{\epsilon}-1}$ $\frac{d(e^{\epsilon}+1)^{2}}{(e^{\epsilon}-1)^{2}}$ $(\frac{e^{\epsilon}+1}{e^{\epsilon}-1})^{2}$
-SS $\sqrt{\frac{8}{\pi}}\frac{de^{\epsilon/2}}{e^{\epsilon}-1}$ $\frac{4de^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ $(\frac{e^{\epsilon}+1}{e^{\epsilon}-1})^{2}$
+SS $\frac{2de^{\epsilon/2}}{e^{\epsilon}-1}$ $\frac{4de^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ $(\frac{e^{\epsilon}+1}{e^{\epsilon}-1})^{2}$
 a-RP $\frac{2de^{\epsilon/2}}{e^{\epsilon}-1}$ $\frac{4de^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ $(\frac{e^{\epsilon}+1}{e^{\epsilon}-1})^{2}$
-s-RP $\frac{de^{\epsilon/4}}{e^{\epsilon/2}-1}$ $\frac{de^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$ $\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$
-TABLE I: Accuracy comparison among frequency-estimation algorithms
-  • Note: HE, RHR, OLH, OCMS, CMS+HE, SS, a-RP, and s-RP represent Hadamard Encoding, Recursive Hadamard Response, Optimized Local Hashing, Optimized CMS, CMS with Hadamard Encoding, Subset Selection, asymmetric RAPPOR, and symmetric RAPPOR, respectively.
-Here, OCMS represents MSE-OCMS and $l$-OCMS when comparing $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively.
-The data source is presented in Section H.
+RP $\frac{de^{\epsilon/4}}{e^{\epsilon/2}-1}$ $\frac{de^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$ $\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$
+TABLE I: Precision comparison among frequency-estimation algorithms when $d\gg e^{\epsilon}$.
+  • Note: HE, RHR, OLH, OCMS+RR, CMS+HE, SS, a-RP, and RP represent Hadamard Encoding, Recursive Hadamard Response, Optimized Local Hashing, Optimized CMS+RR, CMS with Hadamard Encoding, Subset Selection, asymmetric RAPPOR, and symmetric RAPPOR, respectively.
+Here, OCMS+RR represents MSE-OCMS+RR and $l$-OCMS+RR when comparing $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively.
+The data source is presented in Appendix C.
+  • *: the original value from [10] is problematic, and Appendix C recalculates the value as shown.
 Communication Cost (bits)
@@
 OLH $d\epsilon$
-MSE-OCMS $2\log d+\frac{\epsilon}{2}$
-$l$-OCMS $2\log d+\epsilon$
+MSE-OCMS+RR $\max\{2\log d+\frac{\epsilon}{2},\frac{3}{2}\epsilon+6\}$
+$l$-OCMS+RR $\max\{2\log d+\epsilon,3\epsilon+6\}$
 original CMS $\Omega(d\log m)$
@@
 TABLE II: Communication Cost among frequency-estimation algorithms
-  • Note: see Table I for the meaning of abbreviations. The communication cost only accounts for the cost between the server and one client. The data source is presented in Section H.
-Table I shows that the optimized CMS (OCMS) is optimal in $l_{2}$ loss and the worst-case MSE, even considering the constant term.
-Its $l_{1}$ loss is order-optimal.
-Since HE, OLH, and CMS+HE are equivalent to unbiased CMSs without the optimized parameters, it is expected that they are not optimal in at least one dimension.
-Only OCMS and s-RP are optimal in the worst-case MSE, but Table II shows that s-RP requires $d$ bits to communicate while OCMS only requires $2\log_{2}d+\frac{\epsilon}{2}$ bits.
-If the communication cost is required to be $O(\log d)$, only RHR has the same order of accuracy as OCMS in $l_{1}$ and $l_{2}$ losses.
-However, empirical study (Section 6) demonstrates that OCMS outperforms RHR in all datasets, consistent with the theoretical analysis in Section 4.3.
-Therefore, OCMS is the only algorithm optimized for the worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss while communicating efficiently.
+  • Note: see Table I for the meaning of abbreviations. The communication cost only accounts for the cost between the server and one client. The data source is presented in Appendix C.
+Table I shows that the OCMS+RR is leading in reducing $l_{1}$ loss, $l_{2}$ loss, and the worst-case MSE, even considering the constant term.
+Since HE, OLH, and CMS+HE are equivalent to unbiased CMS+RR without the optimized parameters, it is expected that they are not leading in at least one dimension.
+Only OCMS+RR and RP are leading in the worst-case MSE, but Table II shows that RP requires $d$ bits to communicate while OCMS+RR only requires $\max\{2\log d+\frac{\epsilon}{2},\frac{3}{2}\epsilon+6\}$ bits.
+If the communication cost is required to be $O(\log d)$, only RHR has the same order of precision as OCMS+RR in $l_{1}$ and $l_{2}$ losses.
+However, empirical study (Section 6) demonstrates that OCMS+RR outperforms RHR in all datasets, consistent with the theoretical analysis in Section 4.3.
+Therefore, OCMS+RR is the only leading algorithm for the worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss while communicating efficiently.
 5.2 Necessity of Randomness
-[11] proposed an algorithm to convert the Hadamard Response (a special case of CMS) into a deterministic algorithm, arguing that the randomness could be avoided, thereby decreasing the communication cost.
+An algorithm was proposed by [11] to convert the Hadamard Response (a special case of CMS) into a deterministic algorithm, arguing that the randomness could be avoided, thereby decreasing the communication cost.
 However, we will show that one can always create a counterexample causing the frequency estimation to have a constant bias.
-Assume that we are interested in the frequency of value $x$, and $f(x)$ is set to zero (no $x$ in the dataset).
+Assume that we are interested in the frequency of value $x$, and $f(x)$ happens to be zero (no $x$ in the dataset).
 Given any hash function (denoted as $h_{j}$ for convenience), we denote $x_{j}$ as the value colliding with $x$, i.e., $c_{j}(x,x_{j})=1$.
@@
 Theorem 5.1.
-If the assignment of hash functions to the objects, i.e., $\forall i:h_{j^{(i)}}$, is constant in CMS, one can always construct a counterexample making
+If the assignment of hash functions to the objects, $\{h_{j^{(i)}}|i\in n\}$, is constant in CMS, there exists a dataset resulting in
 $E[\hat{f}(x)]-f(x)\neq 0.$
@@
 Here, we focus only on the randomness in assigning hash functions and assume that the LDP process $R$ always uses its own randomness.
-In Section 3.2, the result of $Var(\hat{f}(x))$ is built on $Cov(\hat{y}^{(i_{1})}[h_{j^{(i_{1})}}(x)],\hat{y}^{(i_{2})}[h_{j^{(i_{2})}}(x)])=0$ for any $i_{1}$ and $i_{2}$, which is further derived from the assumption that the hash-function assignments $\mathbf{1}(j^{(i_{1})}=j)$ and $\mathbf{1}(j^{(i_{2})}=j)$ are pairwise independent.
-[…]
+By examining the proof involving hash-function assignments in Appendices D and E, we realize $Var(\hat{f}(x))$ is calculated based on $Cov(\hat{y}^{(i_{1})}[h_{j^{(i_{1})}}(x)],\hat{y}^{(i_{2})}[h_{j^{(i_{2})}}(x)])=0$ for any $i_{1}$ and $i_{2}$, which is further derived from the assumption that the hash-function assignments $\mathbf{1}(j^{(i_{1})}=j)$ and $\mathbf{1}(j^{(i_{2})}=j)$ are pairwise independent.
+Recall that hash-function assignment uniformly samples from $\mathcal{H}$ for each object.
+This process maps $n$ distinct inputs (object IDs) to $n$ repeatable elements in $[k]$ (recall $k=|\mathcal{H}|$), which can be considered as hashing, denoted as $s$.
+We require $s$ to be uniformly sampled from a pairwise-independent hashing family $\mathcal{S}$, which guarantees the pairwise independence of $\mathbf{1}(j^{(i_{1})}=j)$ and $\mathbf{1}(j^{(i_{2})}=j)$ for all $i_{1}$, $i_{2}$, and $j$.
+If we follow Section 3.5 to construct a hashing family mapping from $[n]$ to $[k]$, the parameter $p^{l}$ should exceed $\max\{n,k\}$, which requires $2\lceil\log_{2}(\max\{n,k\}+1)\rceil$ bits to represent one of its mappings.
+We have proved:
 Theorem 5.2.
-$\lceil\max\{4\log_{2}d,2\log_{2}n\}\rceil$ random bits are enough to generate all the hash-function assignments $[j^{(i)}]_{i\in[n]}$ while ensuring their pairwise independence, which is the prerequisite of Eq.
+$2\lceil\log_{2}(\max\{n,k\}+1)\rceil$ random bits are sufficient to generate all the hash-function assignments $[j^{(i)}]_{i\in[n]}$ while ensuring their pairwise independence, which is the prerequisite of Eq.
 (4).
@@
 Previous literature (e.g., [5, 14, 12]) had a strong interest in public randomness, because it was believed that public randomness could save communication costs.
-However, we argue that public and private randomness make no difference for CMS, and no communication cost can be saved, for the following reasons: the server and a client (i.e., an object in CMS) have to sync on which hash function is assigned to the client, and they need to either sync (1) the identifier of a hash function, or (2) the assignment generator.
-There are at least $d^{2}$ hash functions in $\mathcal{H}$, so identifying one of them requires $2\log_{2}d$ bits.
-Meanwhile, Theorem 5.2 shows that at least $4\log_{2}d$ bits are required to represent the assignment generator.
+However, we argue that public and private randomness make no difference for an unbiased CMS, and no communication cost can be saved, for the following reasons: the server and a client (i.e., an object in CMS) have to sync on which hash function is assigned to the client, and they need to either sync (1) the identifier of a hash function, or (2) the assignment generator.
+There are at least $(d+1)^{2}$ hash functions in a pairwise independent $\mathcal{H}_{pi}$, so identifying one of them requires $2\log_{2}d$ bits.
+Meanwhile, Theorem 5.2 shows that at least $2\log_{2}k$ bits are required to represent the assignment generator, where $k$ is the hashing family size, so this is equivalent to requiring at least $4\log_{2}d$ bits.
 Thus, syncing the identifier will be less expensive, and $2\log_{2}d$ bits are needed to sync, regardless of which side generates the assignment.
@@
 The same problem also exists in the public-randomness algorithm in [14].
-The only scenario where public randomness could save communication costs is when all the clients are co-located, referred to as a client group.
+The only scenario where public randomness could save communication costs is when all the clients are clocated, referred to as a client group.
 The traffic inside a client group is assumed to be free.
@@
 In this case, the total communication cost between the server and the client group is $\lceil\max\{4\log_{2}d,2\log_{2}n\}\rceil+n\epsilon$.
-If the client group still uses private randomness, they will return $n(\log d+\epsilon)$ bits, which is larger than the cost from public randomness.
+If the client group still uses private randomness, they will return $n(2\log_{2}d+\epsilon)$ bits, which is larger than the cost from public randomness.
 However, the client-group assumption is contradictory to the motivation of LDP, which usually assumes the clients are distributed.
@@
 6 Experiment
-This section empirically compares the optimized CMS with the algorithms listed below, using both synthetic and real-world datasets:
+This section empirically compares the OCMS+RR (implemented following Section 3.6) with the algorithms listed below, using both synthetic and real-world datasets:
 1. Hadamard Encoding (HE): the first practical frequency estimation algorithm.
@@
 3. Recursive Hadamard Response (RHR): the state-of-the-art version of the Hadamard Response family.
-Other algorithms listed in Section 5.1 are not included due to their very high communication cost, which is impractical with a very large dictionary.
+Other algorithms listed in Section 5.1 are not included due to their very high communication cost, which is impractical for a very large dictionary.
 In the following experiments, we use this estimator to estimate $\widehat{MSE}$:
@@
 However, $\overline{MSE}(\hat{f})$ is biased because $\frac{1}{t}\sum_{t}(\hat{f}(x)-f(x))^{2}$ of each $x$ is equivalent to a random variable fluctuating around $E[(\hat{f}(x)-f(x))^{2}]$, and the max operator will choose the largest deviated value.
-Appendix I provides an upper bound of $\overline{MSE}(\hat{f})$.
+Appendix K provides an upper bound of $\overline{MSE}(\hat{f})$.
 As long as $\overline{MSE}(\hat{f})$ is between the upper bound and the calculated $\widehat{MSE}(\hat{f})$ (both bounds are plotted), we will consider the experimental $\overline{MSE}(\hat{f})$ consistent with the theoretical value from Theorem 13.
@@
 All the experiments below will be run 100 times for each privacy factor $\epsilon$ ranging from 1 to 5.
-6.1 Verification of Optimality and Equivalence
+6.1 Verification of Precision and Equivalence
 We use a synthetic dataset sampled from a specific Zipf’s distribution, which has a frequency distribution as
@@
 The $\widehat{MSE}$, $l_{1}$ loss, and $l_{2}$ loss are evaluated based on the first 100 most frequent values (assuming they are known in advance).
-MSE-OCMS with $f^{*}=1$ (i.e., no prior knowledge on $f(x)$) and $l$-OCMS are evaluated for $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, while RHR, HE, and CMS+HE are evaluated for all the metrics.
+MSE-OCMS+RR with $f^{*}=1$ (i.e., no prior knowledge on $f(x)$) and $l$-OCMS+RR are evaluated for $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, while RHR, HE, and CMS+HE are evaluated for all the metrics.
 Results are plotted in Fig.
 4.
-MSE-OCMS and $l$-OCMS outperform other algorithms in terms of $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, and their accuracy aligns with the calculated values.
+MSE-OCMS+RR and $l$-OCMS+RR outperform other algorithms in terms of $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, and their precision aligns with the calculated values.
 HE and CMS+HE have identical performance in all the metrics, supporting our theory that these two algorithms are equivalent.
@@
 Through several simulations of this distribution, we are confident that the probability of $\max_{x}f(x)\leq 0.01$ is high.
-Thus, we set $f^{*}=0$ for MSE-OCMS and verify if its $\widehat{MSE}$ satisfies Eq.
-(16).
-Meanwhile, $l$-OCMS is evaluated for $l_{1}$ / $l_{2}$ losses.
-RHR, HE, and CMS+HE are also evaluated for $\widehat{MSE}$ $l_{1}$ loss, and $l_{2}$ loss.
-Results are plotted in Fig.
-4.
-Similar to the above, MSE-OCMS and $l$-OCMS outperform other algorithms in terms of $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, and their accuracy aligns with the calculated values.
-HE and CMS+HE have identical accuracy in all the metrics, and they perform the worst.
-The accuracy of RHR lies between OCMS and the other algorithms.
+Thus, we set $f^{*}=0$ for MSE-OCMS+RR and its calculated $\widehat{MSE}$ is derived from Eq.
+(14) as:
+$\widehat{MSE}(\hat{f})=\frac{1}{n}[\max_{x}f(x)+\frac{4e^{\epsilon}}{(e^{\epsilon}-1)^{2}}].$
+Meanwhile, $l$-OCMS+RR is evaluated for $l_{1}$ / $l_{2}$ losses. RHR, HE, and CMS+HE are also evaluated for $\widehat{MSE}$, $l_{1}$ loss, and $l_{2}$ loss. Results are plotted in Fig. 4.
+Similar to the above, MSE-OCMS+RR and $l$-OCMS+RR outperform other algorithms in terms of $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, and their precision aligns with the calculated values.
+HE and CMS+HE have identical precision in all the metrics, and they perform the worst.
+The precision of RHR lies between OCMS+RR and the other algorithms.
 6.3 Real-world Dataset
@@
 It contains over 170,000 entries, all of which belong to a dictionary with a size of 26,000.
-MSE-OCMS with $f^{*}=1$ and $l$-OCMS are evaluated for $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, while RHR, HE, and CMS+HE are evaluated for all the metrics.
+MSE-OCMS+RR with $f^{*}=1$ and $l$-OCMS+RR are evaluated for $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, while RHR, HE, and CMS+HE are evaluated for all the metrics.
 Results are plotted in Fig.
 4.
-MSE-OCMS and $l$-OCMS outperform other algorithms in terms of $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, and their accuracy aligns with the calculated values.
-HE and CMS+HE have identical accuracy in all the metrics, and they perform the worst in $l_{1}$ / $l_{2}$ losses.
-RHR has the same order of accuracy as $l$-OCMS regarding $l_{1}$ / $l_{2}$ losses.
+MSE-OCMS+RR and $l$-OCMS+RR outperform other algorithms in terms of $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively, and their precision aligns with the calculated values.
+HE and CMS+HE have identical precision in all the metrics, and they perform the worst in $l_{1}$ / $l_{2}$ losses.
+RHR has the same order of precision as $l$-OCMS+RR regarding $l_{1}$ / $l_{2}$ losses.
 However, it performs the worst in $\widehat{MSE}$.
@@
 7 Conclusion
-This paper revisits the private Count-mean Sketch (CMS) algorithm, corrects the errors in the expectation and variance calculations in the original CMS paper, and derives the optimized parameters to reduce all kinds of loss functions.
-The optimized CMS is proven theoretically and empirically to be the only algorithm optimized in the worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss, while its communication cost is logarithmic to the dictionary size.
-Moreover, this paper demonstrates that many state-of-the-art algorithms are equivalent to CMS.
+This paper revisits the private CMS algorithm, corrects errors in the expectation and variance calculations in the original CMS paper, and optimizes CMS with randomized response to reduce all kinds of loss functions.
+The optimized CMS is proven theoretically and empirically to be the leading algorithm for reducing the worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss, when requiring the communication cost to be logarithmic to the dictionary size.
+Moreover, this paper demonstrates that many existing algorithms are equivalent to CMS.
 Finally, we discuss the necessity of randomness in CMS and show that both public and private randomness have the same communication cost.
-Appendix A Derivation of the Expectation
+Appendix A Example on Reconstruction Process
+Suppose a dictionary $\mathbb{V}$ has three possible values: 1, 2, and 3; and another dictionary $\mathbb{U}$ has five possible values: $A,B,C,D$, and $E$.
+A mechanism $M$ perturbs a variable $V\in\mathbb{V}$ and outputs a value $U\in\mathbb{U}$.
+The probability that $M$ outputs $U=u$ given $V=v$ is randomly assigned as
+$u$ $v$ 1 2 3
+A 0.133 0.143 0.192
+B 0.031 0.212 0.658
+C 0.287 0.155 0.039
+D 0.161 0.254 0.060
+E 0.389 0.236 0.051
+which is equivalent to matrix $P$. Subsequently, the inverse matrix $Q$ is derived as
+$v$ $u$ A B C D E
+1 0.366 -0.088 2.289 -3.629 2.267
+2 0.010 -0.393 -1.928 6.625 -1.294
+3 0.332 1.560 0.476 -1.996 0.240
+If $M$ perturbs $V$ and outputs $U=C$, the decoding process will return the column of $Q$ corresponding to $C$, which is $[2.289,-1.928,0.476]$.
+If $V=1$, the probability that $M(V)$ outputs $U=C$ is 0.287.
+The same probability applies to the event that the reconstruction process $R(V)$, i.e., $Q[:M(V)]$, generates the corresponding column.
+Denote the column outputted by the reconstruction as $\tilde{V}$.
+One can verify that the expectation of $\tilde{V}$ equals $[1,0,0]$ when $V=1$, which is consistent with Property 2.1.
+Appendix B Proof regarding Optimizing CMS with RR
+We first explore the optimized $m$ for CMS + RR. Substituting Eq. (12) into Eq. (6), we have
+$Var(\hat{f}(x)_{RR})=\frac{1-f(x)}{n(m-1)}+\\ \frac{m[(1-f(x))(e^{\epsilon}-1)(2-m)+me^{\epsilon}]}{n(m-1)(e^{\epsilon}-1)^{2}}$ (20)
+Let us first study the worst-case MSE of CMS+RR, which is equivalent to the worst-case variance $Var(\hat{f}(x)_{RR})$.
+Observing that the variance is linear to $f(x)$, we have $\max_{x}MSE(\hat{f}(x))=\max\{Var(\hat{f}(x)_{RR}|f(x)=0),Var(\hat{f}(x)_{RR}|f(x)=1)\}$.
+When $f(x)=1$, we have
+$\frac{\partial Var(\hat{f}(x)_{RR}|f(x)=1)}{\partial f(x)}=\frac{m(m-2)e^{\epsilon}}{n(m-1)^{2}(e^{\epsilon}-1)^{2}},$
+implying that $Var(\hat{f}(x)_{RR}|f(x)=1)$ increases monotonically with $m$ when $m\geq 2$. On the other hand, when $f(x)=0$, we have
+$\frac{\partial Var(\hat{f}(x)_{RR}|f(x)=0)}{\partial f(x)}=\frac{(m-e^{\epsilon}-1)(m+e^{\epsilon}-1)}{n(m-1)(e^{\epsilon}-1)^{2}},$
+which indicates a minimum at $m=e^{\epsilon}+1$ given the range $m\geq 2$.
+However, when $m=e^{\epsilon}+1$, $Var(\hat{f}(x)_{RR}|f(x)=1)$ is always larger than $Var(\hat{f}(x)_{RR}|f(x)=0)$.
+Given that $Var(\hat{f}(x)_{RR}|f(x)=1)$ decreases as $m$ decreases, while $Var(\hat{f}(x)_{RR}|f(x)=0)$ increases when $m$ decreases within the range $[2,e^{\epsilon}+1]$, there exists an $m$ making them equal to each other.
+Solving
+$Var(\hat{f}(x)_{RR}|f(x)=0)=Var(\hat{f}(x)_{RR}|f(x)=1),$
+we have $m=e^{\epsilon/2}+1$, and the corresponding variance is
+$Var(\hat{f}(x)_{RR})=\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}.$
+When $m\in[2,e^{\epsilon/2}+1]$, we have $Var(\hat{f}(x)_{RR}|f(x)=1)<Var(\hat{f}(x)_{RR}|f(x)=0)$.
+Also, $Var(\hat{f}(x)_{RR}|f(x)=0)$ decreases with $m$ in this range.
+Therefore,when $m\in[2,e^{\epsilon/2}+1]$, we have
+$\max_{f(x)}Var(\hat{f}(x)_{RR})>\max_{f(x)}Var(\hat{f}(x)_{RR}|m=e^{\epsilon/2}+1)\\ =\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}.$
+That is, $m$ should never be smaller than $e^{\epsilon/2}+1$.
+Define $f^{*}$ as the upper bound of $f(x)$, i.e., $\forall x:f(x)\leq f^{*}$.
+When $m\in[e^{\epsilon/2}+1,e^{\epsilon}+1]$, we have proven that $Var(\hat{f}(x)_{RR})$ increases with $f(x)$, so the goal becomes reducing $Var(\hat{f}(x)_{RR}|f(x)=f^{*})$.
+The optimized $m$ can be calculated by
+$\frac{\partial Var(\hat{f}(x)_{RR}|f(x)=f^{*})}{\partial m}=0,$
+which results in
+$m=1+\frac{\Delta_{MSE}}{f^{*}e^{\epsilon}+1-f^{*}},$ (21)
+where
+$\Delta_{MSE}=e^{\epsilon/2}\sqrt{[(1-f^{*})e^{\epsilon}+f^{*}][f^{*}e^{\epsilon}+(1-f^{*})]}.$
+The corresponding variance is
+$Var^{*}(\hat{f}(x)_{RR})=\frac{2(\Delta_{MSE}+e^{\epsilon})}{n(e^{\epsilon}-1)^{2}}$
+Notice that Eq.
+(21) is valid only when $m\in[e^{\epsilon/2}+1,e^{\epsilon}+1]$, which translates to $f^{*}\leq\frac{1}{2}$.
+When $f^{*}>\frac{1}{2}$, we have $m<1+e^{\epsilon/2}$, which has been proven to be worse than $m=1+e^{\epsilon/2}$.
+Therefore, we have
+$m=\begin{cases}1+\Delta_{MSE}&\text{ if }f^{*}\leq\frac{1}{2}\\ 1+e^{\epsilon/2}&\text{ if }\frac{1}{2}<f^{*}\leq 1,\end{cases}$
+which corresponds to
+$Var^{*}(\hat{f}(x)_{RR})=\begin{cases}\frac{2(\Delta_{MSE}+e^{\epsilon})}{n(e^{\epsilon}-1)^{2}}&\text{ if }f^{*}\leq\frac{1}{2}\\ \frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}&\text{ if }\frac{1}{2}<f^{*}\leq 1\end{cases}.$
+Regarding Theorem 3.7, we observe that $Var(\hat{f}(x)_{RR})$ in Eq. (20) is linear to $f(x)$, so
+$Var(\hat{f}(x)_{RR}|f(x)=f^{*}+\delta)-Var(\hat{f}(x)_{RR}|f(x)=f^{*})\\ =\frac{(m-1)^{2}-e^{\epsilon}}{(m-1)(e^{\epsilon}-1)n}\delta.$
+Substituting $m=1+\Delta_{MSE}$ into the above equation and simplifying it, we have:
+$Var(\hat{f}(x)_{RR}|f(x)=f^{*}+\delta)-Var(\hat{f}(x)_{RR}|f(x)=f^{*})\\ =\frac{\delta(1-2f^{*})e^{\epsilon}}{n\Delta_{MSE}}.$
+Since $Var(\hat{f}(x)_{RR})$ increases with $f(x)$ given $m=1+\Delta_{MSE}$, we have
+$\widehat{MSE}(\hat{f}(x))=Var(\hat{f}(x)_{RR}|f(x)=f^{*}+\delta)\\ =\widehat{MSE}(\hat{f}|f^{*})+\frac{\delta(1-2f^{*})e^{\epsilon}}{n\Delta_{MSE}},$
+which proves the equality sign of Theorem 3.7.
+Given that $\frac{e^{\epsilon}}{\Delta_{MSE}}$ decreases when $f^{*}\in[0,\frac{1}{2}]$, we have $\frac{e^{\epsilon}}{\Delta_{MSE}}\leq 1$, which proves the $\leq$” sign of Theorem 3.7.
+Now, we will explore the optimized parameters for $l_{2}$ loss. By considering only the pairwise independence of $\mathcal{H}$, we integrate Eq. (6) into Eq. (11):
+$l_{2}(\hat{f})=\sum_{x\in[d]}\frac{m}{n(m-1)^{2}}\bigl{[}\bigl{(}Var(R|=)+(m-1)Var(R|\neq)\\ +\frac{m-1}{m}\bigr{)}(1-f(x))+mVar(R|=)f(x)\bigr{]}\\ =\frac{m}{n(m-1)^{2}}\bigl{[}\bigl{(}Var(R|=)+(m-1)Var(R|\neq)\\ +\frac{m-1}{m}\bigr{)}(d-1)+mVar(R|=)\bigr{]}$ (22)
+Applying RR and Substituting Eq. (12) into Eq. (22), we obtain
+$l_{2}(\hat{f}_{RR})=\frac{(d+e^{\epsilon}-1)m^{2}}{n(m-1)(e^{\epsilon}-1)^{2}}+\frac{(d-1)(2m+e^{\epsilon}-1)}{n(m-1)(e^{\epsilon}-1)}.$
+By studying its derivative with respect to $m$, we can find its minimum when
+$m=1+\frac{\Delta_{l}}{e^{\epsilon}+d-1},$
+where
+$\Delta_{l}=\sqrt{(e^{\epsilon}+d-1)(de^{\epsilon}-e^{\epsilon}+1)}e^{\epsilon/2}.$
+In practice, we choose the closest integer to $1+\frac{\Delta_{l}}{e^{\epsilon}+d-1}$ as $m$.
+The corresponding optimized $l_{2}$ is the one in Theorem 3.8.
+Given the Cauchy-Schwarz inequality, we have $l_{1}\leq\sqrt{d\>l_{2}}$, which derives the optimized $l_{1}$ in Theorem 3.8.
+Corollary 3.4 is proven accordingly when applying $d\gg e^{\epsilon}$.
+Appendix C Data Source of Section 5.1
+We start with the precision.
+Since HE is equivalent to unbiased CMS+RR with $m=2$, we use Eq.
+(20) to compute its $\widehat{MSE}$, $l_{1}$ loss, and $l_{2}$ losses.
+The computed $\widehat{MSE}$ is the same as in [11].
+The $l_{1}$ and $l_{2}$ losses of RHR were taken from [12], but it did not calculate the $\widehat{MSE}$.
+Section 4.3 shows that RHR is a skewed version of unbiased CMS, and the $\widehat{MSE}$ of RHR could be worse than that from the CMS+RR with $m=2$ (i.e., HE), so the $\widehat{MSE}$ of RHR is $\Omega\left(\left(\frac{e^{\epsilon}+1}{e^{\epsilon}-1}\right)^{2}\right)$.
+OLH is equivalent to the unbiased CMS with $m=1+e^{\epsilon}$, so we use Eq.
+(20) and follow Section 3.4 to calculate its $l_{1}$ loss, $l_{2}$ loss, and $\widehat{MSE}$.
+Section 4.5 shows that CMS+HE is equivalent to the unbiased CMS+RR with $m=2$ (i.e., HE), so its $l_{1}$ loss, $l_{2}$ loss, and $\widehat{MSE}$ are the same as HE.
+The $l_{1}$ and $l_{2}$ losses of SS are presented in [10], but their calculation of $l_{1}$ is incorrect: (1) they approximate $\sum_{x}\in[d]|\hat{f}(x)-f(x)|$ as a normal distribution, whereas the central limit theorem requires each $|\hat{f}(x)-f(x)|$ to be i.i.d, and (2) they derived the result based on the Mean Absolute Deviation of the approximated Gaussian distribution, which is equivalent to using $|\sum_{x}\hat{f}(x)-f(x)|$ to represent $\sum_{x}|\hat{f}(x)-f(x)|$.
+The correct way to compute $l_{1}$ is to use Cauchy–Schwarz inequality, i.e., $l_{1}\leq\sqrt{d\>l_{2}}$, which provides a tight upper bound of $l_{1}$ as $\frac{2de^{\epsilon/2}}{e^{\epsilon}-1}$.
+The $\widehat{MSE}$ of SS is derived using Eq.
+(2) in [9] with
+$p=\frac{ke^{\epsilon}}{ke^{\epsilon}+d-k}$
+$q=\frac{ke^{\epsilon}}{ke^{\epsilon}+d-k}\frac{k-1}{d-1}+\frac{d-k}{ke^{\epsilon}+d-k}\frac{k}{d-1}$
+$k=d/(e^{\epsilon}+1)$
+$f=1$
+where $k$ represents the size of a subset in SS. $f=1$ because the $Var(\hat{f}(x))$ increases with $f(x)$ for SS.
+The $\widehat{MSE}$ of a-RP and RP is calculated using Eq. (2) in [9] as well. Both cases where $f(x)=0$ and $f(x)=1$ are calculated, and the worst-case variance is the maximum between them.
+Next, we present the data source regarding communication cost.
+We first compute the cost of OCMS+RR, which requires the server and a client to sync on the assigned hash function and the perturbed response $z^{(i)}$.
+Section 3.5 shows that $2\lceil\log_{2}\max\{d+1,5m\}\rceil$ bits are required to represent an approximately pairwise independent hash function.
+Also, we need another $\log_{2}m$ bits to sync the perturbed response $z^{(i)}$.
+Given that OCMS+RR has either $m=1+e^{\epsilon/2}$ or $m=1+e^{\epsilon}$ depending on the optimization goal, the overall communication costs of MSE-OCMS+RR and $l$-OCMS+RR are approximately $\max\{2\log_{2}d+\epsilon/2,\frac{3\epsilon}{2}+6\}$ and $\max\{2\log_{2}d+\epsilon/2,3\epsilon+6\}$, respectively.
+The costs of HE and RHR are taken from their respective literature [11] and [12].
+It is notable that we do not use the communication cost from their public-randomness scheme because they fail to consider the cost of syncing the public randomness, as pointed out in Section 5.3.
+For OLH, its hash function maps $[d]$ to $[1+e^{\epsilon}]$, and each mapping is mutually independent, so it requires $d\log_{2}(1+e^{\epsilon})\approx d\epsilon$ bits to represent all the mappings.
+The original CMS [7] did not consider pairwise independence, so its hash function mapping from $[d]$ to $[m]$ requires $d\log_{2}m$ bits to communicate.
+SS randomly selects $\frac{d}{e^{\epsilon}+1}$ values from $[d]$, and this subset selection requires $\frac{d}{e^{\epsilon}+1}$ bits to represent.
+RAPPOR generates a one-hot vector with size $d$, which requires $d$ bits to transfer.
+Appendix D Derivation of the Expectation
 Let’s focus on the term $\hat{y}^{(i)}[h_{j^{(i)}}(x)]$ in the definition of $f(x)$ in Eq. (2), which can be expanded as
@@
 where $\mathbf{1}\{*\}$ returns one if * is true. Given that $h_{j^{(i)}}$ is uniformly sampled from $\mathcal{H}$, we have
-$E[\hat{y}^{(i)}[h_{j^{(i)}}(x)]]\\ =\sum_{j\in[k]}E\bigl{[}\mathbf{1}\{j=j^{(i)}\}R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{]}\\ =\sum_{j\in[k]}\frac{1}{k}c_{j}(X^{(i)},x),$ (25)
+$E[\hat{y}^{(i)}[h_{j^{(i)}}(x)]]=\sum_{j\in[k]}E\bigl{[}\mathbf{1}\{j=j^{(i)}\}\\ R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{]}=\sum_{j\in[k]}\frac{1}{k}c_{j}(X^{(i)},x),$ (24)
 where $c_{j}(X^{(i)},x)$ indicates whether $X^{(i)}$ and $x$ collide in the hash function $h_{j}$, and $E\bigl{[}R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{]}=c_{j}(X^{(i)},x)$ is derived from Property 2.1.
@@
 and Theorem 3.2 is proved.
-Substituting $\forall x,x^{\prime}:\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}=\frac{1}{m}$ into the above equation, we have
+Substituting $\forall x,x^{\prime}$, where $x\neq x^{\prime}$ and $\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}=\frac{1}{m}$ into the above equation, we have
 $E[\hat{f}(x)]=\frac{m}{m-1}[f(x)+\sum_{x^{\prime}\in[d]\backslash x}\frac{1}{m}f(x^{\prime})]-\frac{1}{m-1}\\ =\frac{m}{m-1}[f(x)+(1-f(x))\frac{1}{m}]-\frac{1}{m-1}=f(x).$ (26)
@@
 so Corollary 3.1 is proved.
-Appendix B Derivation of the Variance
+Appendix E Derivation of the Variance
 Let’s start with the variance of an individual response $\hat{y}^{(i)}[h_{j^{(i)}}(x)]$:
-$Var\bigl{(}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{)}\\ =E\bigl{[}\bigl{(}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{)}^{2}\bigr{]}-\bar{c}(X^{(i)},x)^{2}$ (29)
-where $E\bigl{[}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{]}=\bar{c}(X^{(i)},x)$ is proven in Eq. (25). Then,
+$Var\bigl{(}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{)}\\ =E\bigl{[}\bigl{(}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{)}^{2}\bigr{]}-\bar{c}(X^{(i)},x)^{2},$ (28)
+where $E\bigl{[}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{]}=\bar{c}(X^{(i)},x)$ given Eq. (24). Then,
 $E\bigl{[}\bigl{(}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{)}^{2}\bigr{]}\\ =\sum_{j\in[k]}E\bigl{[}\bigl{(}\mathbf{1}\{j=j^{(i)}\}R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{)}^{2}\bigr{]}\\ =\sum_{j\in[k]}E[\mathbf{1}\{j=j^{(i)}\}]E\bigl{[}\bigl{(}R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{)}^{2}\bigr{]}\\ =\frac{1}{k}\sum_{j\in[k]}E\bigl{[}\bigl{(}R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{)}^{2}\bigr{]}\\ \stackrel{{\scriptstyle\text{(a)}}}{{=}}\frac{1}{k}\sum_{j\in[k]}\biggl{(}Var\bigl{(}R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{)}+c_{j}(X^{(i)},x)\biggr{)}$ (29)
@@
 $\frac{1}{k}\sum_{j\in[k]}Var\bigl{(}R(h_{j}(x^{\prime}))[h_{j}(x)]\bigr{)}\\ =\frac{1}{k}\sum_{j\in[k]}[c_{j}(x^{\prime},x)Var(R|=)+(1-c_{j}(x^{\prime},x))Var(R|\neq)]\\ =\bar{c}(x^{\prime},x)Var(R|=)+(1-\bar{c}(x^{\prime},x))Var(R|\neq).$ (37)
-Therefore, when substituting Eq. (38) for $x^{\prime}\neq x$ and the aforementioned parameters with $x=x^{\prime}$ into Eq. (4), we have
-$Var(\hat{f}(x))=\frac{m^{2}}{(m-1)^{2}n}\biggl{[}\\ \sum_{x^{\prime}\in[d]\backslash x}f(x^{\prime})\biggl{(}\bar{c}(x^{\prime},x)Var(R|=)+(1-\bar{c}(x^{\prime},x))\\ Var(R|\neq)+\bar{c}(x^{\prime},x)(1-\bar{c}(x^{\prime},x))\biggr{)}+f(x)Var(R|=)\biggr{]}.$ (39)
-Thus, Eq. 5 is proved.
-If $\bar{c}(x^{\prime},x)=\frac{1}{m}$, it becomes
-$Var(\hat{f}(x))=\frac{m}{(m-1)^{2}n}\biggl{[}\\ \sum_{x^{\prime}\in[d]\backslash x}f(x)\biggl{(}Var(R|=)+(m-1)\\ Var(R|\neq)+\frac{1-m}{m}\biggr{)}+mf(x)Var(R|=)\biggr{]}\\ =\frac{m}{(m-1)^{2}n}\biggl{[}\bigl{(}1-f(x)\big{)}\biggl{(}Var(R|=)+(m-1)\\ Var(R|\neq)+\frac{1-m}{m}\biggr{)}+mf(x)Var(R|=)\biggr{]},$ (40)
-which proves Eq. (6).
-Appendix C Proof of Theorem 3.5
+Therefore, when substituting Eq.
+(37) for $x^{\prime}\neq x$ and the aforementioned parameters with $x=x^{\prime}$ into Eq.
+(4), we derived Eq.
+(5).
+If $\bar{c}(x^{\prime},x)=\frac{1}{m}$, it becomes Eq.
+(6).
+Appendix F Proof of Theorem 3.5
 Recall the definition of $\hat{f}(x)$ in Eq.
@@
 $Pr[|\hat{f}(x)-f(x)|\geq+\alpha\sqrt{Var(\hat{f}(x))}]\\ \leq 2\exp(-\frac{n}{3}\frac{m-1}{e^{\epsilon}+m-1}\alpha^{2}).\qed$ (41)
-Appendix D Proof regarding Optimized CMS
-We first explore the optimized $m$ for the CMS+RR. Substituting Eq. (12) into Eq. (6), we have
-$Var(\hat{f}(x)_{RR})=\frac{1-f(x)}{n(m-1)}+\\ \frac{m[(1-f(x))(e^{\epsilon}-1)(2-m)+me^{\epsilon}]}{n(m-1)(e^{\epsilon}-1)^{2}}$ (45)
-Let’s first study the worst-case MSE of CMS+RR, which is equivalent to the worst-case variance $Var(\hat{f}(x)_{RR})$.
-Observe that the variance is linear to $f(x)$, we have $\max_{x}MSE(\hat{f}(x))=\max\{Var(\hat{f}(x)_{RR}|f(x)=0),Var(\hat{f}(x)_{RR}|f(x)=1)\}$.
-When $f(x)=1$, we have
-$\frac{\partial Var(\hat{f}(x)_{RR}|f(x)=1)}{\partial f(x)}=\frac{m(m-2)e^{\epsilon}}{n(m-1)^{2}(e^{\epsilon}-1)^{2}},$
-implying that $Var(\hat{f}(x)_{RR}|f(x)=1)$ increases monotonically with $m$ when $m\geq 2$. On the other hand, when $f(x)=0$, we have
-$\frac{\partial Var(\hat{f}(x)_{RR}|f(x)=0)}{\partial f(x)}=\frac{(m-e^{\epsilon}-1)(m+e^{\epsilon}-1)}{n(m-1)(e^{\epsilon}-1)^{2}},$
-which indicates a minimum at $m=e^{\epsilon}+1$ given the range $m\geq 2$.
-However, when $m=e^{\epsilon}+1$, $Var(\hat{f}(x)_{RR}|f(x)=1)$ is always larger than $Var(\hat{f}(x)_{RR}|f(x)=0)$.
-Given that $Var(\hat{f}(x)_{RR}|f(x)=1)$ decreases as $m$ decreases, while $Var(\hat{f}(x)_{RR}|f(x)=0)$ increases when $m$ decreases within the range $[2,e^{\epsilon}+1]$, there exists an $m$ making them equal to each other.
-Solving
-$Var(\hat{f}(x)_{RR}|f(x)=0)=Var(\hat{f}(x)_{RR}|f(x)=1),$
-we have $m=e^{\epsilon/2}+1$, and the corresponding variance is
-$Var(\hat{f}(x)_{RR})=\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}.$
-When $12\leq m<e^{\epsilon/2}+$ 1, we have $Var(\hat{f}(x)_{RR}|f(x)=1)<Var(\hat{f}(x)_{RR}|f(x)=0)$.
-Also, $Var(\hat{f}(x)_{RR}|f(x)=0)$ decreases with $m$ in this range.
-Therefore, when $12\leq m\leq e^{\epsilon/2}+$ 1, we have
-$\max_{f(x)}Var(\hat{f}(x)_{RR})>\max_{f(x)}Var(\hat{f}(x)_{RR}|m=e^{\epsilon/2}+1)\\ =\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}.$
-That is, $m$ should never be smaller than $e^{\epsilon/2}+1$.
-Define $f^{*}$ to be the upper bound of $f(x)$, i.e., $\forall x:f(x)\leq f^{*}$.
-When $e^{\epsilon/2}+1\leq m\leq e^{\epsilon}+1$, we have proven that $Var(\hat{f}(x)_{RR})$ increases with $f(x)$, so the goal becomes reducing $Var(\hat{f}(x)_{RR}|f(x)=f^{*})$.
-The optimized $m$ can be calculated by
-$\frac{\partial Var(\hat{f}(x)_{RR}|f(x)=f^{*})}{\partial m}=0,$
-which results in
-$m=1+\frac{\Delta_{MSE}}{f^{*}e^{\epsilon}+1-f^{*}},$ (46)
-where
-$\Delta_{MSE}=e^{\epsilon/2}\sqrt{[(1-f^{*})e^{\epsilon}+f^{*}][f^{*}e^{\epsilon}+(1-f^{*})]}.$
-The corresponding variance is
-$Var^{*}(\hat{f}(x)_{RR})=\frac{2(\Delta_{MSE}+e^{\epsilon})}{n(e^{\epsilon}-1)^{2}}$
-Notice that Eq.
-(46) is valid only when $e^{\epsilon/2}+1\leq m\leq e^{\epsilon}+1$, which translates to $f^{*}\leq\frac{1}{2}$.
-When $f^{*}>\frac{1}{2}$, we have $m<1+e^{\epsilon/2}$, which has been proven to be worse than $m=1+e^{\epsilon/2}$.
-Therefore, we have
-$m=\begin{cases}1+\Delta_{MSE}&\text{ if }f^{*}\leq\frac{1}{2}\\ 1+e^{\epsilon/2}&\text{ if }\frac{1}{2}<f^{*}\leq 1,\end{cases}$
-which corresponds to
-$Var^{*}(\hat{f}(x)_{RR})=\begin{cases}\frac{2(\Delta_{MSE}+e^{\epsilon})}{n(e^{\epsilon}-1)^{2}}&\text{ if }f^{*}\leq\frac{1}{2}\\ \frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}&\text{ if }\frac{1}{2}<f^{*}\leq 1\end{cases}.$
-When it comes to symmetry and asymmetry RAPPOR, since their $Var(a)[a]$ and $Var(a)[b]$ are independent of $m$, the corresponding $Var(\hat{f}(x))$ decreases when $m$ increases.
-Thus, when $m$ is large enough, we have
+Appendix G Proof regarding Preferring RR to RAPPOR in CMS
+The variance of symmetry and asymmetry RAPPOR is listed below:
+$\forall a,b:Var(sRP(a)[b])=\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}.$ (42)
+$Var(aRP(a)[b])=\\ \frac{1}{(e^{\epsilon}-1)^{2}}\begin{cases}(e^{\epsilon}+1)^{2}&\text{ if }a=b\\ 4e^{\epsilon}&\text{ if }a\neq b,\end{cases}$ (43)
+Since their $Var(a)[a]$ and $Var(a)[b]$ are independent of $m$, the corresponding $Var(\hat{f}(x))$ decreases when $m$ increases. Thus, when $m$ is large enough, we have
 $Var(\hat{f}(x)_{RP})\rightarrow\frac{1}{n}[Var(R|\neq)(1-f(x))+\\ Var(R|=)f(x)].$ (44)
+Following Appendix B and given unbiased CMS, the worst-case MSE of symmetry and asymmetry RAPPOR are identical to their variance, which are $\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}$ and $\frac{f^{*}(e^{\epsilon}-1)^{2}+4^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$, respectively.
 For asymmetric RAPPOR, we realize $Var(\hat{f}(x)_{aRP})>Var^{*}(\hat{f}(x)_{RR})$ when $f^{*}>0$, so the CMS with asymmetric RAPPOR cannot serve as the optimized CMS.
@@
 Thus, the CMS with symmetric RAPPOR is also ruled out.
-As a result, Theorem 15 is proved.
-Corollary 3.4 is proved automatically when substituting $f^{*}=1$ into Theorem 15.
-Now, we will explore the optimized parameters for $l_{2}$ loss. By considering only the pairwise independence of $\mathcal{H}$, we integrate Eq. (6) into Eq. (11):
-$l_{2}(\hat{f})=\sum_{x\in[d]}\frac{m}{n(m-1)^{2}}\bigl{[}\bigl{(}Var(R|=)+(m-1)Var(R|\neq)\\ +\frac{m-1}{m}\bigr{)}(1-f(x))+mVar(R|=)f(x)\bigr{]}\\ =\frac{m}{n(m-1)^{2}}\bigl{[}\bigl{(}Var(R|=)+(m-1)Var(R|\neq)\\ +\frac{m-1}{m}\bigr{)}(d-1)+mVar(R|=)\bigr{]}$ (48)
-We will first use RR to perturb the hashed values. Substituting Eq. (12) into Eq. (48), we obtain
-$l_{2}(\hat{f}_{RR})=\frac{(d+e^{\epsilon}-1)m^{2}}{n(m-1)(e^{\epsilon}-1)^{2}}+\\ \frac{(d-1)(2m+e^{\epsilon}-1)}{n(m-1)(e^{\epsilon}-1)}.$ (49)
-By studying its derivative with respect to $m$, we can find its minimum when
-$m=1+\frac{\Delta_{l}}{e^{\epsilon}+d-1},$
-where
-$\Delta_{l}=\sqrt{(e^{\epsilon}+d-1)(de^{\epsilon}-e^{\epsilon}+1)}e^{\epsilon/2}.$
-In practice, we choose the closest integer to $1+\frac{\Delta_{l}}{e^{\epsilon}+d-1}$ as $m$. The corresponding minimum value is
-$\min_{m}l_{2}(\hat{f}_{RR})=\frac{2(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}}.$ (50)
-If the size of candidates $d$ is large enough, we have
-$l_{2}(\hat{f}_{RR})\rightarrow\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}},$ (51)
-and the corresponding $m\rightarrow 1+e^{\epsilon}$.
-Similar to the worst-case MSE, we could prove that the CMS with symmetric RAPPOR will always have larger $l_{2}$ loss than the CMS+RR.
+As a result, The choice of RR is preferred.
+Similar to Appendix B, we can use the variance to derive $l_{2}$ loss, which are $\frac{de^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}$ and $\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$ for symmetric and asymmetric RAPPOR, respectively.
+we observe that CMS with symmetric RAPPOR will always have a larger $l_{2}$ loss than the CMS+RR.
 For asymmetric RAPPOR, its $l_{2}$ loss is the same as that of CMS+RR only when $d$ is large enough.
@@
 $m=\Omega(\frac{(e^{\epsilon}-1)^{2}}{(1+\tau)e^{\epsilon}}).$ (45)
-On the other hand, CMS+RR only requires $\log_{2}(1+e^{\epsilon})$ bits to output the perturbed result.
-Thus, CMS+RR is always preferred when optimizing $l_{2}$ loss due to its low communication cost.
-Now, we have proven the $l_{2}$-loss part in Theorem 3.7 and Corollary 3.5.
-For $l_{1}$ loss, given the Cauchy-Schwarz inequality, we have $l_{1}\leq\sqrt{d\>l_{2}}$. Combining this with Eq. (50) leads to
-$l_{1}(\hat{f})\leq\sqrt{\frac{2d(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}}}.$ (53)
-Define $\lceil l_{1}\rceil$ to be the upper bound of $l_{1}$. If the size of candidates $d$ is large enough, we have
-$\lceil l_{1}(\hat{f})\rceil\rightarrow\frac{2de^{\epsilon/2}}{\sqrt{n}e^{\epsilon}-1}.$ (54)
-Similar to $l_{2}$ loss, the corresponding $m\rightarrow 1+e^{\epsilon}$ (in practice, we choose the closest integer). Thus, the $l_{1}$-loss part in Theorem 3.7 and Corollary 3.5 are proven.
-Appendix E Proof regarding Imperfect Hashing
+On the other hand, CMS+RR only requires $\log_{2}(1+e^{\epsilon})$ bits to output the perturbed result. Thus, CMS+RR is always preferred when optimizing $l_{2}$ loss due to its low communication cost.
+Appendix H Proof regarding Imperfect Hashing
 Let’s start by proving Theorem 3.9.
-Observe $\frac{\partial Var(g(x\mid\mathcal{H}^{\prime}))}{\partial m^{\prime}}$, and note that it is linear to $f(x)$.
+Observe $\frac{\partial Var(g(x\mid\mathcal{H}_{api}))}{\partial m^{\prime}}$, and note that it is linear to $f(x)$.
 Considering the derivative at $f(x)=1$ and $f(x)=0$ yields:
-$-\max\{\frac{2m^{\prime}Var(R|=)}{n(m^{\prime}-1)^{3}},\\ \frac{(m^{\prime}-1)(Var(R|\neq)+1)+(m^{\prime}+1)Var(R|=)}{n(m^{\prime}-1)^{3}}\}\\ \leq\frac{\partial Var(g(x|\mathcal{H}^{\prime}))}{\partial m^{\prime}}<0.$ (55)
+$-\max\{\frac{2m^{\prime}Var(R|=)}{n(m^{\prime}-1)^{3}},\\ \frac{(m^{\prime}-1)(Var(R|\neq)+1)+(m^{\prime}+1)Var(R|=)}{n(m^{\prime}-1)^{3}}\}\\ \leq\frac{\partial Var(g(x|\mathcal{H}_{api}))}{\partial m^{\prime}}<0.$ (46)
 Observing that $Var(R|=)\geq Var(R|\neq)$ is satisfied by the LDP mechanism such as randomized response and RAPPOR, we have
-$-\frac{2m^{\prime}Var(R|=)+m^{\prime}-1}{n(m^{\prime}-1)^{3}}\leq\frac{\partial Var(g(x|\mathcal{H}^{\prime}))}{\partial m^{\prime}}<0.$ (56)
+$-\frac{2m^{\prime}Var(R|=)+m^{\prime}-1}{n(m^{\prime}-1)^{3}}\leq\frac{\partial Var(g(x|\mathcal{H}_{api}))}{\partial m^{\prime}}<0.$ (47)
 At the same time, we have $m^{\prime}=m-\frac{m}{(2q+1)^{2}}$.
 Denote $Var(\hat{f}(x))$ as the variance of the CMS using perfect hashing.
-We want $Var(g(x|\mathcal{H}^{\prime}))$ to approach $Var(\hat{f}(x))$, which is formulated as $Var(g(x|\mathcal{H}^{\prime}))\leq Var(\hat{f}(x))(1+\tau)$, where $\tau$ is a small number like 0.01.
+We want $Var(g(x|\mathcal{H}_{api}))$ to approach $Var(\hat{f}(x))$, which is formulated as $Var(g(x|\mathcal{H}_{api}))\leq Var(\hat{f}(x))(1+\tau)$, where $\tau$ is a small number like 0.01.
 This can be translated to
@@
 thus proving Theorem 3.9.
-When $m=1+e^{\epsilon/2}$, $Var(\hat{f}(x))$ becomes $\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$, and the left-hand side of Eq.
-(57) decreases with $\epsilon$.
+When $m=1+e^{\epsilon/2}$, $Var(\hat{f}(x))$ becomes $\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$, and $\frac{Var(R|=)}{Var(\hat{f}(x))}$ decreases with $\epsilon$.
 Thus, we can substitute $\epsilon=0$ into Eq.
 (48), and we have $2q+1>\sqrt{\frac{1}{\tau}}$.
-Corollary 3.6 is proved.
-If $\forall x:Var(g(x))\leq(1+\tau)Var(\hat{f}(x))$, then $l_{2}(g)\leq(1+\tau)l_{2}(\hat{f}(x))$.
+The MSE part of Corollary 3.5 is proved.
+If $\forall x:Var(g(x))\leq(1+\tau)Var(\hat{f}(x))$, then $l_{2}(g)\leq(1+\tau)l_{2}(\hat{f})$.
 When $m=1+e^{\epsilon}$, $Var(\hat{f}(x))$ increases with $f(x)$, so we have $Var(\hat{f}(x)|f(x)=0)\leq Var(\hat{f}(x))$.
@@
 (49) decreasing with $\epsilon$.
-Thus, this inequality is still satisfied when $\epsilon=0$, which derives $2q+1>\sqrt{\frac{1}{\tau}}$.
-Thus, Corollary 3.7 is proved.
-Now we will prove Theorem 3.9. First, by calculating the derivative of $Pr(h^{\prime}(x_{1})=h^{\prime}(x_{2}))$ with respect to $r$ to find the maximum, we have
-$Pr(h^{\prime}(x_{1})=h^{\prime}(x_{2}))\leq\frac{1}{m}+\frac{1}{4mq(q+1)},$ (59)
-which is equivalent to $\bar{c}(x,x^{\prime})\leq\frac{1}{m}+\frac{1}{4mq(q+1)}$.
-Next, consider the $MSE(\hat{f}(x))$ in Eq.
-(7).
-We aim for $MSE(\hat{f}(x))$ to approach $Var(\hat{f}(x))$, where $\hat{f}(x)$ denotes the estimation with perfect pairwise independent hashing.
-Regarding the bias term of $MSE(\hat{f}(x))$, we have
-$E[\hat{f}(x|\mathcal{H}^{\prime})]-f(x)=\frac{m}{m-1}(\bar{c}(x,x^{\prime})-\frac{1}{m})\sum_{x^{\prime}\in[d]\backslash x}f(x^{\prime})\\ \leq\frac{m}{m-1}\frac{1}{4mq(q+1)}=\frac{1}{4(m-1)q(q+1)}.$ (60)
-Another term in $MSE(\hat{f}(x))$ is $Var(\hat{f}(x))$, which is linear to $\bar{c}(x,x^{\prime})$ and $\bar{c}(x,x^{\prime})^{2}$, and all the parameters of these two terms are positive. Therefore,
-$\frac{Var[\hat{f}(x|\mathcal{H}^{\prime})]}{Var[\hat{f}(x)]}\leq\max\{\frac{\bar{c}(x,x^{\prime})}{1/m},\frac{\bar{c}(x,x^{\prime})^{2}}{1/m^{2}}\}\\ =\frac{(2q+1)^{4}}{16q^{2}(q+1)^{2}}.$ (61)
-To satisfy
-$MSE(f(x)|\mathcal{H}^{\prime})-MSE(\hat{f}(x))\leq 2\tau Var[\hat{f}(x)],$
-it is equivalent to requiring
-$(E[\hat{f}(x|\mathcal{H}^{\prime})]-f(x))^{2}<\tau Var[\hat{f}(x)]$
-$\frac{Var[\hat{f}(x|\mathcal{H}^{\prime})]}{Var[\hat{f}(x)]}<1+\tau$
-Substituting Eqs. (60) and (61) into the above equations proves Theorem 3.9.
-Appendix F Proof regarding Correcting the Original CMS
-F.1 Proof of Theorem 3.10
+Since this inequality is still satisfied when $\epsilon=0$, it derives $2q+1>\sqrt{\frac{1}{\tau}}$.
+Thus, The $l_{2}$ part of Corollary 3.5 is proved.
+Appendix I Proof of Theorem 3.10
 Consider each $c_{j}(x,x^{\prime})$ as an independent Bernoulli random variable with probability $\frac{1}{m}$ of being one.
@@
 Thus, Theorem 3.10 is proved.
-F.2 Proof of Theorem 22
-Given Markov’s inequality, Eq. (22) is equivalent to requiring
-$\frac{\mathop{Var}_{\forall c_{j}(x,x^{\prime})}(E[\hat{f}(x)])}{\tau Var(\hat{f}(x))}\leq\delta.$ (64)
-Theorem 3.10 has demonstrated that
-$\mathop{Var}_{\forall c_{j}(x,x^{\prime})}(E[\hat{f}(x)])\leq\frac{1}{(m-1)k}$ (65)
-The original CMS uses RAPPOR or Hadamard Encoding (HE) to perturb the hash values in Step 2, so their variance can be calculated using Eq. (5), which can be scaled down as
-$Var(\hat{f}(x))>\frac{1}{n}\min\{Var(R|=),Var(R|\neq)\}\\ =\frac{Var(R|\neq)}{n},$ (66)
-given that $Var(R|=)>Var(R|\neq)$ in both RAPPOR and HE. Combining Eq. (65) and (66), we have
-$\frac{\mathop{Var}_{\forall c_{j}(x,x^{\prime})}(E[\hat{f}(x)])}{Var(\hat{f}(x))}<\frac{n}{(m-1)kVar(R|\neq)}.$
-Substituting the above inequality into Eq. (64), the requirement of Eq. (22) is converted as
-$\frac{n}{(m-1)kVar(R|\neq)}<\delta,$ (67)
-which is the same as the condition of Theorem 22.
-Appendix G Decoding Hadamard Encoding
+Appendix J Decoding Hadamard Encoding
 Define $H$ to be a scaled Walsh Hadamard matrix, where $H[i,j]=(-1)^{i\cdot j}$ with $\cdot$ representing bitwise multiplication.
@@
 (2).
-Appendix H Data Source of Section 5.1
-We start with the accuracy.
-Since HE is equivalent to unbiased CMS+RR with $m=2$, we use Eq.
-(45) to compute its $\widehat{MSE}$, $l_{1}$ loss, and $l_{2}$ losses.
-The computed $\widehat{MSE}$ is the same as in [11].
-The $l_{1}$ and $l_{2}$ losses of RHR were taken from [12], but it did not calculate the $\widehat{MSE}$.
-Section 4.3 shows that RHR is a skewed version of unbiased CMS, and the $\widehat{MSE}$ of RHR could be worse than that from the CMS+RR with $m=2$ (i.e., HE), so the $\widehat{MSE}$ of RHR is $\Omega\left(\left(\frac{e^{\epsilon}+1}{e^{\epsilon}-1}\right)^{2}\right)$.
-OLH is equivalent to the unbiased CMS with $m=1+e^{\epsilon}$, so we use Eq.
-(45) and follow Section 3.4 to calculate its $l_{1}$ loss, $l_{2}$ loss, and $\widehat{MSE}$.
-Section 4.5 shows that CMS+HE is equivalent to the unbiased CMS+RR with $m=2$ (i.e., HE), so its $l_{1}$ loss, $l_{2}$ loss, and $\widehat{MSE}$ are the same as HE.
-The $l_{1}$ and $l_{2}$ losses of SS are presented in [10].
-Its $\widehat{MSE}$ is derived using Eq.
-(2) in [9] with
-$p=\frac{ke^{\epsilon}}{ke^{\epsilon}+d-k}$
-$q=\frac{ke^{\epsilon}}{ke^{\epsilon}+d-k}\frac{k-1}{d-1}+\frac{d-k}{ke^{\epsilon}+d-k}\frac{k}{d-1}$
-$k=d/(e^{\epsilon}+1)$
-$f=1$
-where $k$ represents the size of a subset in SS. $f=1$ because the $Var(\hat{f}(x))$ increases with $f(x)$ for SS.
-The $\widehat{MSE}$ of a-RP and s-RP is calculated using Eq. (2) in [9] as well. Both cases where $f(x)=0$ and $f(x)=1$ are calculated, and the worst-case variance is the maximum between them.
-Next, we present the data source regarding communication cost.
-We first compute the cost of OCMS, which requires the server and a client to sync on the assigned hash function and the perturbed response $z^{(i)}$.
-Section 3.5 shows that a hash function is comprised of two random numbers, each ranging from 1 to $\Theta(d)$, so $2\log_{2}d$ bits are sufficient to represent one hash function.
-Then, we need another $\log_{2}m$ bits to sync the perturbed response $z^{(i)}$.
-Given that OCMS has either $m=1+e^{\epsilon/2}$ or $m=1+e^{\epsilon}$ depending on the optimization goal, the bits to transfer $z^{(i)}$ are close to $\epsilon/2$ and $\epsilon$, respectively.
-Thus, the overall communication cost of MSE-OCMS and $l$-OCMS is $22\log_{2}d+\epsilon/$ 2 and $2\log_{2}d+\epsilon$, respectively.
-The costs of HE and RHR are taken from their respective literature [11] and [12].
-It is notable that we do not use the communication cost from their public-randomness scheme because they fail to consider the cost of syncing the public randomness, as pointed out in Section 5.3.
-For OLH, its hash function maps $[d]$ to $[1+e^{\epsilon}]$, and each mapping is mutually independent, so it requires $d\log_{2}(1+e^{\epsilon})\approx d\epsilon$ bits to represent all the mappings.
-The original CMS [7] did not consider pairwise independence, so its hash function mapping from $[d]$ to $[m]$ requires $d\log_{2}m$ bits to communicate.
-SS randomly selects $\frac{d}{e^{\epsilon}+1}$ values from $[d]$, and this subset selection requires $\frac{d}{e^{\epsilon}+1}$ bits to represent.
-RAPPOR generates a one-hot vector with size $d$, which requires $d$ bits to transfer.
-Appendix I Upper Bound of the worst-case MSE estimator
+Appendix K Upper Bound of the worst-case MSE estimator
 The upper bound of $\overline{MSE}(\hat{f})$ is set to be $[1+\frac{2}{t}\bigl{(}\sqrt{t\log(20|\mathbf{x}|)}+\log(20|\mathbf{x}|)\bigr{)}]V$ based on the following theorem:
-Theorem I.1.
+Theorem K.1.
 If $\forall x\in\mathbf{x}:E[(\hat{f}(x)-f(x))^{2}]\leq V$, where $V$ is constant, we have
@@
 where $t$ is the experiment rounds.
-Proof: Similar to Appendix C, we define $Y^{(i)}=\frac{m}{m-1}\hat{y}^{(i)}[h_{j}^{(i)}(x)]-\frac{1}{m-1}$, and $\frac{Y^{(i)}-A}{B-A}$ is a Bernoulli random variable (see Appendix C for the definition of $A$ and $B$).
+Proof: Similar to Appendix F, we define $Y^{(i)}=\frac{m}{m-1}\hat{y}^{(i)}[h_{j}^{(i)}(x)]-\frac{1}{m-1}$, and $\frac{Y^{(i)}-A}{B-A}$ is a Bernoulli random variable (see Appendix F for the definition of $A$ and $B$).
 All the $X^{(i)}$ can be placed into two groups based on whether $X^{(i)}=x$.